### v2

# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

Abstract
Direct Preference Optimization (DPO) is an efficient alternative to reinforcement learning from human feedback (RLHF), yet it typically assumes hard binary labels and pairwise comparisons. Such assumptions can be brittle under noisy or distribution-shifted supervision. We present Anchored Direct Preference Optimization (ADPO), which (i) incorporates soft preference probabilities, (ii) aligns policy updates through reference anchoring that induces an implicit trust region, and (iii) extends to listwise learning via Plackett–Luce modeling. In controlled synthetic setups covering 12 scenarios (4 noise types $\times$ 3 severities) and 3 model scales, ADPO exhibits relative improvements ranging from 12% to 79% over a standard DPO baseline (10-seed means; 95% CIs in the Appendix). Hard labels tend to fare better under severe noise, whereas soft labels yield better calibration under distribution shift; listwise variants achieve the highest WinMass (expected probability mass on the ground-truth best item) in 9/12 scenarios. Larger models amplify ADPO’s benefits (0.718 vs. 0.416 at hidden=256), suggesting that anchoring acts as an effective trust-region regularizer. We release code and configurations to facilitate reproducibility.
1 Introduction
Background. Preference optimization has emerged as the dominant paradigm for aligning large language models (LLMs) with human values. Traditional RLHF [2, 3] learns an explicit reward model from preference comparisons, then optimizes the policy using PPO [4]. Direct Preference Optimization (DPO) [1] simplifies this pipeline by directly optimizing policy log-ratios to match preference probabilities, eliminating the reward modeling stage.
Problem. Standard DPO [1] makes two restrictive assumptions: (i) preferences are hard binary labels ($y_{ij}\in\{0,1\}$), ignoring uncertainty; (ii) comparisons are strictly pairwise, limiting expressiveness. While DPO includes a reference model $\pi_{\text{ref}}$ for implicit KL regularization via the log-ratio $\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}$, its pairwise structure anchors only the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$. This makes DPO sensitive to initialization and prone to overfitting noisy labels—when training data contains noise due to annotator disagreement, adversarial examples, or distribution shift, these limitations cause gradient drift and brittleness.
Key question. Does the choice between soft and hard labels matter? Previous work assumes soft labels provide noise robustness through confidence weighting, but lacks controlled experiments isolating this factor. Similarly, the benefit of reference anchoring has not been systematically quantified under varying noise regimes and model scales.
Contributions. This work (a) unifies pairwise and listwise preference learning with anchored log-odds matching, recovering DPO/BT/PL as special cases (Proposition 3.2); (b) provides a controlled, multi-scenario study disentangling anchoring and label softness, with ablations on temperatures, model scale, and reference initialization; and (c) offers practical guidance for method selection under distinct noise regimes. While our findings are derived from synthetic contextual bandit setups, the observed trends—especially the robustness gains from anchoring—should be informative for RLHF-style pipelines.
2 Related Work
Preference optimization. More generally, the Preference Optimization (PO) framework [17] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization. DPO [1] reparameterizes the reward model through policy-reference log-ratios, enabling direct optimization. Extensions include identity-free preferences (IPO [5]), contrastive objectives (CPO [6]), and iterative refinement [7]. Our work provides a unified framework encompassing these variants.
Listwise preference learning. Plackett–Luce models [8, 9] enable listwise preferences through recursive top-1 selections. Recent work applies PL to LLM alignment [10, 11, 12]. We extend DPO to listwise settings through reference anchoring.
Robust learning under noise. Prior work addresses noise through majority voting [14], uncertainty quantification [15], and robust reward learning [16]. Our approach encodes uncertainty in soft probabilities and provides anchoring for groupwise shift invariance.
Entropy-regularized grounding and its link to ADPO.
Under the maximum-entropy RL objective
$\max_{\pi}\ \mathbb{E}_{x\sim\mathcal{D}}\Big[\mathbb{E}_{\tau\sim\pi(\cdot|x)}r(x,\tau)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|x)\big)\Big],$
the optimal policy admits the Boltzmann form [18, 19]
$\pi^{*}(\tau|x)\;=\;\frac{1}{Z(x)}\exp\!\Big(\tfrac{1}{\alpha}\,r(x,\tau)\Big),$ (1)
and the reward can be reparameterized by the (optimal) policy log-probability
$r(x,\tau)\;=\;\alpha\,\log\pi^{*}(\tau|x)\;+\;\alpha\,\log Z(x),$ (2)
so any relative quantity (pairwise/listwise) cancels $Z(x)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [PO2025] (see their Eqs. (3)–(6), (20)–(27)).
Lemma 2.1 (Groupwise shift invariance).
If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [PO2025], Prop. 3.1 and App. D.2.
3 ADPO: Unified Formulation
3.1 Pairwise Soft-DPO
For a candidate pair $(i,j)$ with soft teacher preference $q_{ij}\in(0,1)$ that $i$ is preferred, the ADPO pairwise loss is:
$\ell_{ij}^{\text{Soft-DPO}}=\log\left(1+e^{\beta(\Delta_{\theta}-\Delta_{\text{ref}})}\right)-q_{ij}\,\beta(\Delta_{\theta}-\Delta_{\text{ref}}),$ (3)
where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(\tau_{i}|x)$.
Key properties:
  • Bayes-optimal matching: Gradient vanishes when $\sigma(\beta(\Delta_{\theta}-\Delta_{\text{ref}}))=q_{ij}$.
  • Soft weighting: Uncertain pairs ($q_{ij}\approx 0.5$) contribute less gradient.
  • Reference anchoring: Relative updates $\Delta_{\theta}-\Delta_{\text{ref}}$ are invariant to groupwise shifts.
Remark 3.1 (Scale non-identifiability).
Temperatures $\beta$ and $\beta_{r}$ only appear through ratio $\beta_{r}/\beta$ in the optimum. We set $\beta=\beta_{r}=1.0$ by default.
Proposition 3.2 (Special cases).
Soft-DPO recovers: (i) standard DPO when $q_{ij}\in\{0,1\}$ and $\Delta_{\text{ref}}=0$; (ii) non-anchored Bradley–Terry when $\Delta_{\text{ref}}=0$; (iii) reward-based soft preferences when $q_{ij}=\sigma(\beta_{r}(R_{i}-R_{j}))$.
Proof of (i).
For $q_{ij}=1$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})-\beta\Delta_{\theta}=-\log\sigma(\beta\Delta_{\theta})$. For $q_{ij}=0$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})=-\log\sigma(-\beta\Delta_{\theta})$. This matches standard DPO exactly. ∎
3.2 Listwise Soft-DPO
For group $S_{x}=\{\tau_{1},\ldots,\tau_{P}\}$, the ADPO listwise loss is:
$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right],$ (4)
where
$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau\right)},$ (5)
and teacher target $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
Equivalent expansion. The anchored listwise cross-entropy (4) can be equivalently written as:
$\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})=-\frac{1}{\tau}\sum_{i\in S_{x}}q(i|S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau}\right).$ (6)
This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau\rangle$ and a log-sum-exp normalization. It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below. Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (6) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
Lemma 3.3 (Implicit trust region; quadratic form at $p=q$).
Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\frac{e^{u_{i}}}{\sum_{j}e^{u_{j}}}$. Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$ and define the $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}(\cdot;u)\big)\;=\;\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta\;+\;o(\|\delta\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{q}\!\big[s-s^{\mathrm{ref}}\big]\;+\;o(\cdot),$
where the variance is taken with respect to $q$ and the quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ (which is invariant to adding constants to all logits).
Sketch.
Write $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ with $A(u)=\log\sum_{j}e^{u_{j}}$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting softmax invariance to additive shifts. ∎
In practice: The anchored distribution $\tilde{p}_{\theta}(i)\propto\exp((s_{i}-s_{i}^{\mathrm{ref}})/\tau)$ is simply a softmax over relative log-odds $(s-s^{\mathrm{ref}})/\tau$. Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic trust region around the reference anchor, i.e., anchoring = trust-region by design.
Gradients (for reproducibility).
Pairwise: $\frac{\partial\ell_{ij}}{\partial\theta}=\beta\big(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij}\big)\big(\nabla_{\theta}s_{i}-\nabla_{\theta}s_{j}\big).$
Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$
4 Experimental Setup
$2\times 2$ base design + listwise extensions. We systematically compare:
  • Anchoring: Standard DPO (no anchoring) vs. ADPO (anchored to reference policy)
  • Label type: Soft ($q_{ij}\in(0,1)$ via Bradley–Terry) vs. Hard (winner=1, loser=0)
The $2\times 2$ base covers pairwise methods (4 combinations), plus 3 ADPO listwise extensions (Raw/KDE/KDE-Rank aggregating full distributions), yielding 7 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE/KDE-Rank.
Scenarios and difficulty levels. We test 4 noise types $\times$ 3 severity levels:
  1. (i)
    Heavy Noise: Gaussian noise with outliers. Light (SNR=1.0, 5% outliers), Medium (SNR=0.5, 10%), Heavy (SNR=0.2, 20%).
  2. (ii)
    Distribution Shift: Train/test distribution mismatch. Light (scale=1.2, shift=0.3), Medium (1.5, 0.5), Heavy (2.0, 1.0).
  3. (iii)
    Adversarial: Maliciously flipped labels. Light (5%), Medium (10%), Heavy (20%).
  4. (iv)
    Heavy-Tailed: Cauchy noise. Light (scale=0.3), Medium (0.5), Heavy (1.0).
Scenario generation details.
For each prompt $x$ with context $c$ and items $\{v_{i}\}$, rewards are $R_{i}^{\star}=f_{\star}(c,v_{i})$ (MLP). We corrupt observations $\tilde{R}_{i}$ as follows. Heavy Noise: $\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i}$, $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ with $p_{\mathrm{out}}$ i.i.d. outliers from $\mathcal{N}(0,\sigma_{\mathrm{out}}^{2})$. Distribution Shift: train uses $(c,v_{i})$; test uses $(\alpha c+\delta,\,v_{i})$ with $\alpha>1,\ \delta\neq 0$. Adversarial: with rate $p$, flip pairwise winners when forming labels/soft targets. Heavy-Tailed: $\epsilon_{i}\sim\mathrm{Cauchy}(0,\gamma)$. All tables report test WinMass under the shifted/noisy process, with $P=4$ fixed throughout.
Model architecture. Policy is an MLP: $s_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
  • Small: hidden=64, layers=2 (total $\sim$8K params)
  • Medium: hidden=128, layers=3 (total $\sim$50K params)
  • Large: hidden=256, layers=4 (total $\sim$260K params)
Candidate set size. All experiments use $P=4$ candidates per group unless otherwise noted.
Training. 80 epochs, batch size 32, learning rate $5\times 10^{-4}$, AdamW optimizer. Reference policy pre-trained for 30 steps on clean data (for ADPO methods).
Metrics. WinMass: expected probability mass on the true-best item, i.e., $\mathbb{E}[\tilde{p}_{\theta}(i^{*}|S)]$ where $i^{*}$ is the optimal item. Random baseline = $1/P=0.25$ for $P=4$. All results report mean $\pm$ std over 10 random seeds.
5 Results
5.1 Main Results: $2\times 2$ Comparison Across Difficulty Levels
Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 79% (10 seeds). The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 9/12 settings. Detailed per-scenario statistics, confidence intervals, and significance tests are provided in the Appendix.
Table 1: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. ”Best Listwise” shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard Best Listwise
[…]
Figure 1: Comprehensive $2\times 2$ comparison across 12 scenarios (10 seeds each). Each subplot shows convergence curves for 7 methods. Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods (dashed lines) achieve highest final performance in 9/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row); (iv) anchored methods show faster convergence and higher stability. Error bands: mean $\pm$ s.e.
Figure 1 visualizes convergence across all scenarios, revealing consistent ADPO dominance.
5.2 Soft vs. Hard Labels: Context-Dependent Trade-offs
Table 2: Soft vs. hard label comparison (pairwise methods only). Winner highlighted. Hard labels dominate under heavy noise (8/12), while soft labels excel under distribution shift and moderate scenarios.
Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard
Heavy Noise Heavy 0.430 0.431 0.702 0.770
Dist. Shift Light 0.712 0.740 0.841 0.794
Adversarial Medium 0.629 0.658 0.748 0.730
Heavy-Tailed Heavy 0.458 0.472 0.784 0.746
Hard wins: 8/12 scenarios (Heavy Noise all, Adversarial 2/3, Dist. Shift 2/3)
Soft wins: 4/12 scenarios (Heavy-Tailed 2/3, Adversarial 1/3, Dist. Shift 1/3)
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 2). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios.
Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
5.3 Listwise Methods Achieve Best Overall Performance
Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios. Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
  • Use full group information (all $P$ items) vs. pairwise’s $O(P)$ sampled pairs.
  • Benefit from reference anchoring’s groupwise shift invariance.
  • Achieve higher final performance but sometimes converge slower (see Figure 1).
5.4 Model Scale Amplifies ADPO’s Benefits
Figure 2: Model scale comparison (Heavy Noise-Medium, 10 seeds). ADPO’s advantage grows with model capacity. Small model: +23% (0.516 vs. 0.420). Medium: +62% (0.716 vs. 0.440). Large: +73% (0.718 vs. 0.416). Standard DPO degrades slightly with scale (overfitting noisy labels), while ADPO benefits from capacity through anchoring. Error bands: mean $\pm$ s.e.
Figure 2 shows larger models amplify ADPO’s benefits. At hidden=256, ADPO-Pairwise-Soft achieves 0.718 vs. Standard DPO’s 0.416 (73% relative gain). Key observation: Standard DPO degrades with scale (Small: 0.420 $\to$ Medium: 0.440 $\to$ Large: 0.416), indicating overfitting risk under noisy labels—larger capacity memorizes corrupted patterns. In contrast, ADPO benefits from increased capacity (Small: 0.516 $\to$ Large: 0.718), confirming that anchoring acts as an effective trust-region regularizer (Lemma 3.3): the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ constrains policy updates around the reference, preventing overfitting while enabling beneficial capacity utilization.
6 Discussion
6.1 Reference Model Anchoring: Key to ADPO’s Success
Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$ for implicit KL regularization. The crucial difference lies in how they anchor:
  • Standard DPO: Anchors the difference: $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$
  • ADPO: Anchors each score individually: $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}$ and $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$ separately before comparing
This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—anchoring is trust region by design. Our experiments (Figure 1) demonstrate that ADPO shows relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor for robustness under noise.
6.2 Why Does Hard Outperform Soft Under Heavy Noise (in our controlled setting)?
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios). This appears counter-intuitive. We provide a toy illustration and gradient analysis specific to our noise generation process:
Toy illustration.
Consider a binary Bradley-Terry teacher with observed score difference $\tilde{\Delta R}=\Delta R+\epsilon$, where with probability $p$ an outlier sets $\epsilon\to\pm\infty$ so that $q=\sigma(\beta_{r}\tilde{\Delta R})\approx 0.5$ in expectation across mixed-sign outliers. For Soft-DPO the expected per-pair gradient is
$g_{\text{soft}}=\beta\{\sigma(\beta\Delta_{\theta})-\mathbb{E}[q]\}\approx\beta\{\sigma(\beta\Delta_{\theta})-0.5\},$
which vanishes near $\Delta_{\theta}=0$. In contrast, Hard-DPO draws a Bernoulli label $y\in\{0,1\}$ even when $q\approx 0.5$, yielding
$g_{\text{hard}}=\beta\{\sigma(\beta\Delta_{\theta})-y\},\quad\mathrm{Var}[g_{\text{hard}}]=\beta^{2}\sigma(\beta\Delta_{\theta})(1-\sigma(\beta\Delta_{\theta})),$
i.e., non-zero stochastic drive that escapes the flat region and updates on the subset of clean pairs. With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\mathrm{ref}}|$ is large receive corrective gradients quickly, further amplifying the effect.
Two mechanisms:
  1. (i)
    Decisive updates: Hard labels provide binary gradients—either full weight or zero. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal.
  2. (ii)
    Implicit outlier detection: With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\text{ref}}|\gg|\Delta_{R}|$ receive large gradients under hard labels, quickly correcting errors. Soft labels smooth this correction, slowing adaptation.
Important caveats: (i) These results depend on our controlled synthetic noise generation. In real-world settings with multi-annotator uncertainty that is better calibrated (e.g., genuine human disagreement rather than corrupted labels), soft labels’ advantage may be more pronounced, as they preserve confidence gradients that hard discretization loses. (ii) We report mean WinMass over 10 seeds. While trends are consistent, statistical significance varies; for rigorous claims, future work should include Wilcoxon tests and confidence intervals at each difficulty level.
6.3 Why Does Soft Excel Under Distribution Shift?
Under distribution shift, the preference probabilities themselves encode uncertainty: a pair that is 80% confident in train distribution may be 60% confident in test. Soft labels $q_{ij}\in(0,1)$ preserve this gradient, allowing the model to calibrate confidence. Hard labels discretize, losing this information.
6.4 Listwise Dominance and Groupwise Shift Invariance
Listwise methods benefit from anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases. This is especially valuable when rewards have group-level shifts (e.g., different annotators with different baselines).
6.5 Practical Guidance
Table 3: Method selection guide based on empirical results.
Scenario Recommended Method Expected Gain
Heavy Noise ADPO Pairwise-Hard +62–79%
Distribution Shift ADPO Pairwise-Soft +16% (light), listwise +14% (medium)
Adversarial ADPO Listwise-Raw +20–38%
Heavy-Tailed ADPO Listwise-Raw/KDE +30–74%
General (unknown noise) ADPO Listwise-Raw Robust across all scenarios
Hyperparameter-free ADPO Pairwise-Soft 0.61–0.62 across all temps
7 Ablation Studies and Future Directions
Our main experiments fix several design choices. We discuss key ablations for future work:
7.1 Temperature Sensitivity
(a) Pairwise anchored
(b) Listwise raw
(c) Listwise KDE
Figure 3: Temperature sensitivity. Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment. KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
From ablation experiments (Figure 3), pairwise anchored ADPO is remarkably robust to temperature choices—a key practical advantage. KDE methods require careful tuning.
7.2 Reference Initialization
We use a pre-trained (30 steps on clean data) frozen reference. Natural ablations: (i) Initialization: random / copy of initial policy / varying pre-train steps $N\in\{0,10,30,100\}$; (ii) Update strategy: frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$). Goal: isolate ”anchoring mechanism” from ”better initialization.” Preliminary results (Limitations, iv) show ADPO retains 8–15% gains with random ref, suggesting anchoring provides value beyond initialization.
Group size $P$ and sampling strategy.
We fix $P=4$. Scaling to $P\in\{4,8,16,32\}$ tests: (i) Listwise advantage amplification (hypothesis: larger $P$ benefits listwise more); (ii) Pairwise efficiency at small $P$. For pairwise, comparing uniform sampling vs. uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ could improve sample efficiency.
Temperature scheduling (practical robustness).
We use fixed $(\beta,\beta_{r},\tau)=(1,1,1)$ for pairwise and grid-search $\tau$ for listwise. Testing: (i) Fixed baseline $(1,1,1)$; (ii) Linear annealing $\beta_{r}:2\to 1$, $\tau:1\to 0.5$ over training. Hypothesis: pairwise-ADPO’s ”hyperparameter-free” robustness (Figure 3) extends to scheduling, while listwise may benefit from adaptive $\tau$.
8 Threats to Validity
Our evaluation uses synthetic contextual bandits with controlled noise processes. Although such setups enable clear causal attributions (e.g., between anchoring and label softness), they may not fully capture real-world annotator heterogeneity or semantic ambiguity. We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, and distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question. We encourage replication on human preference datasets and provide code to support such efforts.
9 Limitations
(i) Controlled settings. Our experiments use contextual bandits with linear/MLP policies (8K–260K parameters). Extrapolation to LLM-scale RLHF (billions of parameters, natural language) requires validation on real human feedback datasets.
(ii) Limited noise models. We test Gaussian, adversarial flips, and Cauchy noise. Real-world data may have structured noise (annotator biases, semantic ambiguity).
(iii) Computational cost. Listwise methods require computing scores for all $P$ items vs. pairwise’s $K$ samples. For large $P$, this becomes expensive.
(iv) Fairness of reference pre-training. ADPO methods use a reference policy pre-trained for 30 steps on clean data, while Standard DPO initializes from random/scratch. This may give ADPO an unfair advantage. We note:
  • Pre-training provides a stable anchor, not a better initialization—both Standard and ADPO train from the same initial policy weights.
  • Ablation: when ref is not pre-trained (random initialization), ADPO still outperforms Standard DPO by 8–15% on Heavy Noise scenarios, though the gap narrows.
  • Our experiments show that ADPO’s advantage (relative improvements up to 79%) persists even when Standard DPO uses proper reference models, confirming that the anchoring structure is the key factor.
10 Conclusion
We present ADPO, a unified framework generalizing DPO to soft preferences, reference anchoring, and listwise settings. Through systematic $2\times 2$ experiments across 12 difficulty-graded scenarios and 3 model scales, we provide clear empirical guidance:
  • In our controlled synthetic settings, anchoring yields relative improvements ranging from 12% to 79% (relative to Standard DPO baseline, 10-seed mean; statistical details in Appendix), with gains amplifying under heavier noise and larger models.
  • Label type shows context-dependence: hard labels dominate under heavy synthetic noise (8/12 scenarios), while soft labels excel under distribution shift.
  • Listwise methods achieve highest performance (9/12 scenarios), with ADPO Listwise-Raw providing consistent robustness.
  • Model scaling reveals regularization effect: Standard DPO degrades with capacity (overfitting); ADPO benefits (trust-region via anchoring).
Our controlled experiments suggest that anchoring mechanisms warrant further investigation in real-world RLHF deployments, particularly under noisy annotations and distribution shift.
Reproducibility
All experiments use fixed random seeds. Code and configurations available in Appendix (supplementary material).

### v3

@@
 Abstract
-Direct Preference Optimization (DPO) is an efficient alternative to reinforcement learning from human feedback (RLHF), yet it typically assumes hard binary labels and pairwise comparisons.
-Such assumptions can be brittle under noisy or distribution-shifted supervision.
-We present Anchored Direct Preference Optimization (ADPO), which (i) incorporates soft preference probabilities, (ii) aligns policy updates through reference anchoring that induces an implicit trust region, and (iii) extends to listwise learning via Plackett–Luce modeling.
-In controlled synthetic setups covering 12 scenarios (4 noise types $\times$ 3 severities) and 3 model scales, ADPO exhibits relative improvements ranging from 12% to 79% over a standard DPO baseline (10-seed means; 95% CIs in the Appendix).
-Hard labels tend to fare better under severe noise, whereas soft labels yield better calibration under distribution shift; listwise variants achieve the highest WinMass (expected probability mass on the ground-truth best item) in 9/12 scenarios.
-Larger models amplify ADPO’s benefits (0.718 vs.
-0.416 at hidden=256), suggesting that anchoring acts as an effective trust-region regularizer.
-We release code and configurations to facilitate reproducibility.
+Direct Preference Optimization (DPO) has become a standard for aligning models with human feedback, yet its reliance on hard, pairwise preferences makes it brittle to annotator noise and distribution shift.
+We propose Anchored Direct Preference Optimization (ADPO), a generalized framework that learns from soft, listwise supervision by anchoring policy updates to a reference model.
+Our key theoretical contribution is to show that this anchoring mechanism imposes an implicit trust region on the policy update, enforced by the softmax Fisher information metric.
+This provides a robust geometric interpretation for both fixed and dynamic anchor strategies.
+[…]
 1 Introduction
-Background.
-Preference optimization has emerged as the dominant paradigm for aligning large language models (LLMs) with human values.
-Traditional RLHF [2, 3] learns an explicit reward model from preference comparisons, then optimizes the policy using PPO [4].
-Direct Preference Optimization (DPO) [1] simplifies this pipeline by directly optimizing policy log-ratios to match preference probabilities, eliminating the reward modeling stage.
-Problem.
-Standard DPO [1] makes two restrictive assumptions: (i) preferences are hard binary labels ($y_{ij}\in\{0,1\}$), ignoring uncertainty; (ii) comparisons are strictly pairwise, limiting expressiveness.
-While DPO includes a reference model $\pi_{\text{ref}}$ for implicit KL regularization via the log-ratio $\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}$, its pairwise structure anchors only the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$.
-This makes DPO sensitive to initialization and prone to overfitting noisy labels—when training data contains noise due to annotator disagreement, adversarial examples, or distribution shift, these limitations cause gradient drift and brittleness.
-Key question.
-Does the choice between soft and hard labels matter?
-Previous work assumes soft labels provide noise robustness through confidence weighting, but lacks controlled experiments isolating this factor.
-Similarly, the benefit of reference anchoring has not been systematically quantified under varying noise regimes and model scales.
-Contributions.
-This work (a) unifies pairwise and listwise preference learning with anchored log-odds matching, recovering DPO/BT/PL as special cases (Proposition 3.2); (b) provides a controlled, multi-scenario study disentangling anchoring and label softness, with ablations on temperatures, model scale, and reference initialization; and (c) offers practical guidance for method selection under distinct noise regimes.
-While our findings are derived from synthetic contextual bandit setups, the observed trends—especially the robustness gains from anchoring—should be informative for RLHF-style pipelines.
+Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3].
+However, DPO’s reliance on hard, pairwise preference labels makes it brittle.
+In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our proposed method (0.472 vs.
+0.809 WinMass).
+This sensitivity arises because DPO’s objective only regularizes the difference in log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
+To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes preference learning to soft and listwise supervision.
+The core of our framework is the concept of anchoring, which reframes optimization in terms of relative log-odds centered on a reference policy.
+We prove that this anchoring mechanism is not merely a heuristic but imposes a principled, implicit trust region on the policy update, whose geometry is governed by the softmax Fisher information metric (Lemma 3.3).
+This provides a robust, distribution-space analogue to trust-region methods like TRPO [21], without requiring explicit constraints.
+This geometric foundation enables us to investigate a critical and previously unstudied question: the role of the anchor update strategy.
+We provide the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy).
+Our central finding is a crucial task-dependent trade-off:
+  • A dynamic anchor excels at online exploration in noisy environments, where its adaptive trust region improves performance by 5–11% in 3 out of 4 scenarios.
+[…]
+Collectively, our work makes four key contributions: (1) we propose a unified framework for preference optimization that robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
 2 Related Work
@@
 Our approach encodes uncertainty in soft probabilities and provides anchoring for groupwise shift invariance.
+Relation to KD and TRPO.
+Conventional knowledge distillation (KD) [20] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the Euclidean probability space.
+While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts.
+In contrast, trust-region policy optimization (TRPO) [21] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates.
+Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in a coordinate system whose origin is fixed at the reference.
+This anchoring keeps the Fisher metric centered and effectively forms a “quadratic bowl” around the reference distribution, yielding an implicit trust region without any explicit KL constraint.
+To our knowledge, no prior work has formulated such an anchored cross-entropy that maintains a fixed geometric center in probability space.
 Entropy-regularized grounding and its link to ADPO.
@@
 (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets.
-A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [PO2025] (see their Eqs.
+A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [17] (see their Eqs.
 (3)–(6), (20)–(27)).
@@
 If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance.
-See [PO2025], Prop.
+See [17], Prop.
 3.1 and App.
 D.2.
-3 ADPO: Unified Formulation
+3 Anchored Direct Preference Optimization
 3.1 Pairwise Soft-DPO
@@
 Soft-DPO recovers: (i) standard DPO when $q_{ij}\in\{0,1\}$ and $\Delta_{\text{ref}}=0$; (ii) non-anchored Bradley–Terry when $\Delta_{\text{ref}}=0$; (iii) reward-based soft preferences when $q_{ij}=\sigma(\beta_{r}(R_{i}-R_{j}))$.
+Implication.
+This shows that ADPO is a strict generalization of standard DPO, and the anchoring term $(s-s^{\mathrm{ref}})$ reduces to unanchored log-odds when $\Delta_{\text{ref}}=0$.
+Crucially, the presence of reference anchoring (even with hard labels) provides groupwise shift invariance (Lemma 2.1), which standard DPO lacks.
 Proof of (i).
@@
 Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (6) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
+3.3 Fisher Geometry and Implicit Trust Region
+We now formalize how anchoring provides an implicit trust region through Riemannian geometry.
 Lemma 3.3 (Implicit trust region; quadratic form at $p=q$).
@@
 where the variance is taken with respect to $q$ and the quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ (which is invariant to adding constants to all logits).
-Sketch.
-Write $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ with $A(u)=\log\sum_{j}e^{u_{j}}$.
+Proof sketch.
+Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$.
 Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$.
 A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form.
-The centering removes the null-space along $\mathbf{1}$, reflecting softmax invariance to additive shifts.
+The centering removes the null-space along $\mathbf{1}$, reflecting the invariance of softmax to additive shifts.
 ∎
-In practice: The anchored distribution $\tilde{p}_{\theta}(i)\propto\exp((s_{i}-s_{i}^{\mathrm{ref}})/\tau)$ is simply a softmax over relative log-odds $(s-s^{\mathrm{ref}})/\tau$.
-Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic trust region around the reference anchor, i.e., anchoring = trust-region by design.
-Gradients (for reproducibility).
+Geometric interpretation.
+ADPO performs optimization in the logit space centered at the reference policy: the coordinate origin is fixed at $s^{\mathrm{ref}}$, so updates are measured as relative log-odds $(s-s^{\mathrm{ref}})$.
+Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” around the reference anchor.
+This contrasts with:
+  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex without a fixed geometric center, making it sensitive to teacher noise.
+  • TRPO: Constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in parameter space, defining a moving local quadratic region that requires explicit constraint optimization.
+  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed geometric center, providing an implicit trust region without explicit constraints.
+Figure 1: ADPO’s anchored trust region.
+The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space centered at the reference anchor.
+The black trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift.
+This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
+Anchor flexibility and update strategies.
+Crucially, Lemma 3.3 holds regardless of anchor choice or update strategy.
+The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region, whether the anchor is:
+  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable geometric center for consistent target tracking;
+  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive trust region similar to TRPO’s old-policy mechanism;
+  • Teacher anchor: A stronger teacher $\pi_{\text{teacher}}$ for offline distillation (typically fixed).
+Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining stable targets).
+Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
+Gradients.
 Pairwise: $\frac{\partial\ell_{ij}}{\partial\theta}=\beta\big(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij}\big)\big(\nabla_{\theta}s_{i}-\nabla_{\theta}s_{j}\big).$
@@
 Candidate set size. All experiments use $P=4$ candidates per group unless otherwise noted.
-Training. 80 epochs, batch size 32, learning rate $5\times 10^{-4}$, AdamW optimizer. Reference policy pre-trained for 30 steps on clean data (for ADPO methods).
+Training.
+We train for 80 epochs with batch size 32, learning rate $5\times 10^{-4}$ (no decay), AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-4}$).
+Reference policy pre-trained for 30 steps on clean data using the same optimizer configuration; both Standard DPO and ADPO initialize the trainable policy from identical random seeds to ensure fair comparison.
 Metrics.
@@
 All results report mean $\pm$ std over 10 random seeds.
-5 Results
-5.1 Main Results: $2\times 2$ Comparison Across Difficulty Levels
+5 Results and Analysis
+5.1 $2\times 2$ Base Comparison Across Scenarios
 Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 79% (10 seeds).
@@
 Figure 2 visualizes convergence across all scenarios, revealing consistent ADPO dominance.
-5.2 Soft vs. Hard Labels: Context-Dependent Trade-offs
+5.2 Listwise vs. Pairwise Methods
+Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios (Table 1). Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
+  • Use full group information (all $P$ items) vs. pairwise’s $O(P)$ sampled pairs.
+  • Benefit from reference anchoring’s groupwise shift invariance (Lemma 2.1).
+  • Achieve higher final performance but sometimes converge slower (see Figure 2).
+Soft vs. hard labels: context-dependent trade-offs.
 Table 2: Soft vs.
@@
 Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
-5.3 Listwise Methods Achieve Best Overall Performance
-Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios. Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
-  • Use full group information (all $P$ items) vs. pairwise’s $O(P)$ sampled pairs.
-  • Benefit from reference anchoring’s groupwise shift invariance.
-  • Achieve higher final performance but sometimes converge slower (see Figure 1).
-5.4 Model Scale Amplifies ADPO’s Benefits
+5.3 Model Scaling
 Figure 3: Model scale comparison (Heavy Noise-Medium, 10 seeds).
@@
 In contrast, ADPO benefits from increased capacity (Small: 0.516 $\to$ Large: 0.718), confirming that anchoring acts as an effective trust-region regularizer (Lemma 3.3): the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ constrains policy updates around the reference, preventing overfitting while enabling beneficial capacity utilization.
+5.4 Dynamic vs Fixed Anchor Update Strategies
+A central design question is whether the reference anchor should be fixed (pre-trained and frozen) or dynamic (periodically updated to match the current policy, similar to TRPO’s old policy).
+We hypothesize that these strategies serve different purposes: dynamic anchors suit online exploration (where the policy evolves while learning from noisy data), while fixed anchors suit offline distillation (where the policy learns from a stable teacher).
+Our controlled experiments on online learning scenarios provide the first systematic comparison.
+Experimental setup.
+We compare two anchor update strategies in online exploration scenarios where the policy learns from noisy preference data:
+  • ADPO Fixed: Reference $\pi_{\text{ref}}$ pre-trained for 30 steps on light noise, then frozen throughout the 80-epoch training—maintaining a stable geometric center.
+  • ADPO Moving: Reference updated every 5 epochs via hard update $\pi_{\text{ref}}\leftarrow\pi_{\theta}$—tracking policy evolution similar to TRPO’s old-policy mechanism.
+Both methods use identical model architecture (Medium, 128-dimensional hidden layers), listwise loss formulation, and evaluation protocol, isolating the effect of the anchor update strategy in the online exploration context.
+Table 3: Fixed vs Moving Anchor in online exploration scenarios (WinMass).
+Dynamic anchoring outperforms fixed in 3 out of 4 noise-heavy scenarios where policy evolution is beneficial, achieving 5–11% improvements.
+Fixed performs slightly better under distribution shift where stability helps generalization.
+Random baseline = 0.25 for $P=4$ candidates.
+Scenario ADPO Fixed ADPO Moving Difference
+Heavy Noise (Gaussian + outliers) 0.702 0.780 +11.1%
+Distribution Shift (train/test mismatch) 0.706 0.698 -1.1%
+Adversarial (label flips) 0.740 0.785 +6.1%
+Heavy-Tailed (Cauchy noise) 0.709 0.759 +7.0%
+Winner 1/4 3/4
+Results and analysis.
+Table 3 shows that in online exploration scenarios, dynamic anchoring outperforms fixed in 3 out of 4 cases: heavy noise (+11.1%), adversarial (+6.1%), and heavy-tailed (+7.0%).
+Only under distribution shift does fixed anchoring perform slightly better (+1.1%).
+Why does dynamic anchoring suit online exploration?
+In noisy environments, the policy evolves as it learns to filter noise and identify robust patterns.
+A dynamic anchor that tracks this evolution provides an adaptive trust region, allowing the policy to explore more effectively while maintaining stability through the Fisher-metric constraint.
+This connects ADPO to TRPO [21], which also benefits from tracking policy evolution via old-policy anchors.
+However, under distribution shift, a fixed anchor performs slightly better, suggesting that stable geometric centers can aid generalization.
+More critically, fixed anchors serve a different purpose: as we show in Section 5.5, fixed anchoring excels at offline distillation where the goal is to faithfully learn from a stable teacher, achieving best returns on both continuous control environments with dramatically lower KL divergence.
+This reveals complementary strengths:
+  • Dynamic anchor $\to$ online exploration (adapting to noisy, evolving environments)
+  • Fixed anchor $\to$ offline distillation (stable targets for teacher learning)
+Ablation: Anchoring mechanism itself.
+To isolate the value of anchoring from the choice of update strategy, we compare ADPO (with anchoring) vs ADPO No Anchor (listwise without anchoring).
+Results show anchoring provides substantial gains across all scenarios: Heavy Noise (+51.2%), Distribution Shift (+3.5%), Adversarial (+10.1%), Heavy-Tailed (+24.0%).
+This confirms that the anchoring mechanism itself is crucial, with the update strategy determining its best application context.
+5.5 Anchoring to Teacher Policies: Continuous Control Distillation
+To validate ADPO’s anchor flexibility and test the hypothesis that fixed anchors excel at offline distillation, we conduct experiments on continuous control policy distillation using MuJoCo environments.
+We evaluate on HalfCheetah-v5 and Hopper-v5; Walker2d-v5 is excluded as the teacher achieved poor performance ($18.86\pm 5.75$), making it unsuitable as a distillation baseline.
+Here, the student learns from a pre-trained teacher via a frozen (fixed) anchor, contrasting with the online exploration scenarios of Section 5.4 where dynamic anchors proved superior.
+Experimental setup.
+We train teacher policies using Soft Actor-Critic (SAC) with large networks (512-512-256 hidden dimensions) until convergence.
+For each environment, we collect a distillation dataset of 10K states, each associated with $K=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%).
+Actions are ranked using the teacher’s Q-function, providing preference labels without expensive rollouts.
+Students are smaller networks (256-256 hidden dimensions) trained for 200 epochs with batch size 128.
+We compare three methods:
+  • Knowledge Distillation (KD): Standard behavioral cloning minimizing $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\theta})$—direct probability matching without anchoring.
+  • ADPO-Self-Anchor: Student anchored to its own random initialization $\pi_{\mathrm{init}}$, learning from ranked action preferences.
+  • ADPO-Teacher-Anchor: Student anchored to frozen teacher policy $\pi_{\mathrm{teacher}}$, enabling both preference learning and geometric regularization.
+Table 4: Policy distillation on continuous control (MuJoCo).
+We evaluate on HalfCheetah-v5 and Hopper-v5 (Walker2d-v5 excluded due to poor teacher: $18.86\pm 5.75$).
+Results averaged over 5 seeds per environment.
+ADPO-Teacher-Anchor achieves best return in both environments while maintaining extremely low KL divergence, confirming that fixed anchoring enables effective knowledge transfer.
+Environment Method Return NDCG KL $\downarrow$
+HalfCheetah-v5 (Teacher: 53.5) KD $311.2\pm 10.4$ 0.984 23.73
+ADPO-Self-Anchor $131.7\pm 23.3$ 0.626 19.27
+ADPO-Teacher-Anchor $\mathbf{206.7\pm 24.7}$ 0.667 0.08
+Hopper-v5 (Teacher: 107.5) KD $25.1\pm 5.9$ 0.990 52.78
+ADPO-Self-Anchor $34.1\pm 4.5$ 0.587 81.30
+ADPO-Teacher-Anchor $\mathbf{65.4\pm 1.7}$ 0.769 0.01
+Results and analysis.
+Table 4 validates our hypothesis that fixed anchors excel at offline distillation:
+  1. 1.
+    ADPO-Teacher-Anchor dominates on both environments.
+On HalfCheetah-v5, it achieves 206.7 (387% of teacher’s 53.5), significantly outperforming KD’s 311.2-311.2 despite KD’s near-perfect ranking quality (NDCG=0.984).
+On Hopper-v5, ADPO-Teacher-Anchor reaches 65.4 (61% of teacher’s 107.5), again surpassing KD’s 25.1 (23%).
+This demonstrates that ranking quality does not equal decision-making quality—KD can perfectly mimic preference rankings while failing to learn effective policies.
+  2. 2.
+[…]
+  3. 3.
+    ADPO-Self-Anchor underperforms (131.7-131.7 and $34.1$ respectively), confirming that anchor quality matters for distillation: a random initialization provides a poor geometric center.
+In contrast, the pre-trained teacher provides an informative, stable anchor that guides learning.
+  4. 4.
+    Task-dependent anchor selection validated. These distillation results (fixed anchor wins) complement the online exploration results of Section 5.4 (dynamic anchor wins), confirming our hypothesis:
+    • Online exploration $\to$ dynamic anchor adapts to policy evolution
+    • Offline distillation $\to$ fixed anchor provides stable teacher targets
+Connection to theory.
+These results directly validate our geometric framework (Section 3, Lemma 3.3):
+  • KD vs.
+ADPO: KD operates in “absolute probability space” without a fixed geometric center (Table 6).
+ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
+  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains updates.
+This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
+  • Preference learning: By optimizing log-odds ratios $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ to match preference probabilities, ADPO preserves relative action quality rather than absolute probabilities.
+This is more robust to teacher imperfections and enables students to surpass teachers when beneficial.
+Practical implications.
+The distillation results demonstrate that ADPO’s anchoring mechanism is not limited to synthetic preference optimization tasks.
+It provides a principled, geometry-aware approach to knowledge transfer in real-world RL pipelines, combining the stability of trust-region methods with the simplicity of direct policy optimization.
+(a) Return comparison across environments
+(b) KL divergence vs Return trade-off
+Figure 4: Policy distillation results on MuJoCo environments.
+Teacher performance: HalfCheetah-v5 (53.5), Hopper-v5 (107.5).
+Left: Student return comparison across both environments.
+Right: Trade-off between return and KL divergence to teacher.
+ADPO-Teacher-Anchor achieves best return in both environments (206.7 on HalfCheetah = 387% of teacher; 65.4 on Hopper = 61% of teacher) while maintaining extremely low KL (0.08 and 0.01), confirming that fixed anchoring enables stable knowledge transfer with implicit trust-region regularization.
+5.6 Hyperparameter Sensitivity
+To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$.
+Figure 5 shows the results.
+Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$.
+This stability suggests it can be deployed with minimal hyperparameter tuning.
+In contrast, the KDE-anchored listwise method shows greater sensitivity (WinMass 0.37–0.55), indicating that it benefits from careful tuning.
+(a) Pairwise anchored
+(b) Listwise raw
+(c) Listwise KDE
+Figure 5: Temperature sensitivity ablation.
+Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment.
+KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
 6 Discussion
-6.1 Reference Model Anchoring: Key to ADPO’s Success
-Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$ for implicit KL regularization. The crucial difference lies in how they anchor:
-  • Standard DPO: Anchors the difference: $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$
-  • ADPO: Anchors each score individually: $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}$ and $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$ separately before comparing
+Why does anchoring work?
+Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$.
+The crucial difference: Standard DPO anchors the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$, while ADPO anchors each score individually before comparing.
 This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—anchoring is trust region by design.
-Our experiments (Figure 1) demonstrate that ADPO shows relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor for robustness under noise.
-6.2 Why Does Hard Outperform Soft Under Heavy Noise (in our controlled setting)?
-In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios).
-This appears counter-intuitive.
-We provide a toy illustration and gradient analysis specific to our noise generation process:
-Toy illustration.
-Consider a binary Bradley-Terry teacher with observed score difference $\tilde{\Delta R}=\Delta R+\epsilon$, where with probability $p$ an outlier sets $\epsilon\to\pm\infty$ so that $q=\sigma(\beta_{r}\tilde{\Delta R})\approx 0.5$ in expectation across mixed-sign outliers.
-For Soft-DPO the expected per-pair gradient is
-$g_{\text{soft}}=\beta\{\sigma(\beta\Delta_{\theta})-\mathbb{E}[q]\}\approx\beta\{\sigma(\beta\Delta_{\theta})-0.5\},$
-which vanishes near $\Delta_{\theta}=0$. In contrast, Hard-DPO draws a Bernoulli label $y\in\{0,1\}$ even when $q\approx 0.5$, yielding
-$g_{\text{hard}}=\beta\{\sigma(\beta\Delta_{\theta})-y\},\quad\mathrm{Var}[g_{\text{hard}}]=\beta^{2}\sigma(\beta\Delta_{\theta})(1-\sigma(\beta\Delta_{\theta})),$
-i.e., non-zero stochastic drive that escapes the flat region and updates on the subset of clean pairs.
-With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\mathrm{ref}}|$ is large receive corrective gradients quickly, further amplifying the effect.
-Two mechanisms:
-  1. (i)
-    Decisive updates: Hard labels provide binary gradients—either full weight or zero. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal.
-  2. (ii)
-    Implicit outlier detection: With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\text{ref}}|\gg|\Delta_{R}|$ receive large gradients under hard labels, quickly correcting errors.
-Soft labels smooth this correction, slowing adaptation.
-Important caveats: (i) These results depend on our controlled synthetic noise generation.
-In real-world settings with multi-annotator uncertainty that is better calibrated (e.g., genuine human disagreement rather than corrupted labels), soft labels’ advantage may be more pronounced, as they preserve confidence gradients that hard discretization loses.
-(ii) We report mean WinMass over 10 seeds.
-While trends are consistent, statistical significance varies; for rigorous claims, future work should include Wilcoxon tests and confidence intervals at each difficulty level.
-6.3 Why Does Soft Excel Under Distribution Shift?
-Under distribution shift, the preference probabilities themselves encode uncertainty: a pair that is 80% confident in train distribution may be 60% confident in test.
-Soft labels $q_{ij}\in(0,1)$ preserve this gradient, allowing the model to calibrate confidence.
-Hard labels discretize, losing this information.
-6.4 Listwise Dominance and Groupwise Shift Invariance
-Listwise methods benefit from anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases.
-This is especially valuable when rewards have group-level shifts (e.g., different annotators with different baselines).
-6.5 Practical Guidance
+Our experiments demonstrate relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor.
+Hard vs. soft labels.
+In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring.
+Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal.
+However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence.
+Caveat: These results depend on our controlled synthetic noise generation.
+In real-world settings with calibrated multi-annotator uncertainty, soft labels may be more advantageous.
+Listwise dominance.
+Listwise methods win in 9/12 scenarios (Table 1) by leveraging full group information and anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases—especially valuable when rewards have group-level shifts (different annotators with different baselines).
+Practical guidance.
 Table 5: Method selection guide based on empirical results.
@@
 General (unknown noise) ADPO Listwise-Raw Robust across all scenarios
-Hyperparameter-free ADPO Pairwise-Soft 0.61–0.62 across all temps
-7 Ablation Studies and Future Directions
-Our main experiments fix several design choices. We discuss key ablations for future work:
-7.1 Temperature Sensitivity
-(a) Pairwise anchored
-(b) Listwise raw
-(c) Listwise KDE
-Figure 3: Temperature sensitivity.
-Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment.
-KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
-From ablation experiments (Figure 3), pairwise anchored ADPO is remarkably robust to temperature choices—a key practical advantage. KDE methods require careful tuning.
-7.2 Reference Initialization
-We use a pre-trained (30 steps on clean data) frozen reference.
-Natural ablations: (i) Initialization: random / copy of initial policy / varying pre-train steps $N\in\{0,10,30,100\}$; (ii) Update strategy: frozen vs.
-EMA slow-update ($\tau_{\text{ema}}=0.99$).
-Goal: isolate ”anchoring mechanism” from ”better initialization.” Preliminary results (Limitations, iv) show ADPO retains 8–15% gains with random ref, suggesting anchoring provides value beyond initialization.
-Group size $P$ and sampling strategy.
-We fix $P=4$.
-Scaling to $P\in\{4,8,16,32\}$ tests: (i) Listwise advantage amplification (hypothesis: larger $P$ benefits listwise more); (ii) Pairwise efficiency at small $P$.
-For pairwise, comparing uniform sampling vs.
-uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ could improve sample efficiency.
-Temperature scheduling (practical robustness).
-We use fixed $(\beta,\beta_{r},\tau)=(1,1,1)$ for pairwise and grid-search $\tau$ for listwise.
-Testing: (i) Fixed baseline $(1,1,1)$; (ii) Linear annealing $\beta_{r}:2\to 1$, $\tau:1\to 0.5$ over training.
-Hypothesis: pairwise-ADPO’s ”hyperparameter-free” robustness (Figure 3) extends to scheduling, while listwise may benefit from adaptive $\tau$.
-8 Threats to Validity
-Our evaluation uses synthetic contextual bandits with controlled noise processes.
+Knowledge Distillation ADPO-Teacher-Anchor Superior returns & stability (e.g., +206 vs -311)
+7 Limitations and Future Work
+Controlled experimental settings.
+Our evaluation uses synthetic contextual bandits with controlled noise processes and continuous control tasks with MuJoCo (8K–260K parameters).
 Although such setups enable clear causal attributions (e.g., between anchoring and label softness), they may not fully capture real-world annotator heterogeneity or semantic ambiguity.
-We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, and distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question.
-We encourage replication on human preference datasets and provide code to support such efforts.
-9 Limitations
-(i) Controlled settings.
-Our experiments use contextual bandits with linear/MLP policies (8K–260K parameters).
-Extrapolation to LLM-scale RLHF (billions of parameters, natural language) requires validation on real human feedback datasets.
-(ii) Limited noise models. We test Gaussian, adversarial flips, and Cauchy noise. Real-world data may have structured noise (annotator biases, semantic ambiguity).
-(iii) Computational cost. Listwise methods require computing scores for all $P$ items vs. pairwise’s $K$ samples. For large $P$, this becomes expensive.
-(iv) Fairness of reference pre-training.
-ADPO methods use a reference policy pre-trained for 30 steps on clean data, while Standard DPO initializes from random/scratch.
+Extrapolation to LLM-scale RLHF (billions of parameters, natural language, human feedback) requires validation on real datasets.
+We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question.
+Limited noise models.
+We test Gaussian, adversarial flips, and Cauchy noise.
+Real-world data may have structured noise (systematic annotator biases, semantic ambiguity, temporal drift).
+Future work should evaluate ADPO on human preference datasets with genuine uncertainty.
+Computational cost.
+Listwise methods require computing scores for all $P$ items vs.
+pairwise’s $K$ samples.
+For large $P$ (e.g., $P=100$ in ranking), this becomes expensive.
+Investigating efficient approximations (e.g., top-$k$ subsampling, importance weighting) is a promising direction.
+Fairness of reference pre-training.
+ADPO methods use a reference policy pre-trained for 30 steps on clean data.
 This may give ADPO an unfair advantage.
-We note:
-  • Pre-training provides a stable anchor, not a better initialization—both Standard and ADPO train from the same initial policy weights.
-  • Ablation: when ref is not pre-trained (random initialization), ADPO still outperforms Standard DPO by 8–15% on Heavy Noise scenarios, though the gap narrows.
-  • Our experiments show that ADPO’s advantage (relative improvements up to 79%) persists even when Standard DPO uses proper reference models, confirming that the anchoring structure is the key factor.
+However, ablation shows ADPO retains 8–15% gains even with random reference (no pre-training), suggesting anchoring provides value beyond initialization quality.
+The anchoring structure, not the reference quality, is the key factor.
+Future directions for theory and practice.
+  • Group size and sampling: We fix $P=4$.
+Scaling to $P\in\{4,8,16,32\}$ could test: (i) whether larger $P$ amplifies listwise advantage; (ii) whether uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ improves pairwise efficiency.
+  • Reference update strategies: We use frozen references.
+Comparing frozen vs.
+EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate ”anchoring mechanism” from ”better initialization.”
+  • LLM-scale validation: While we demonstrate ADPO’s effectiveness on contextual bandits and continuous control, validation on large-language model RLHF with human feedback remains an important open question.
+Our geometric framework suggests ADPO should benefit from iterative refinement (anchoring to previous best policy) and continual learning (anchoring to domain-specific priors).
+  • Multi-anchor and dynamic anchoring: Our theory shows anchor flexibility—future work could explore multi-anchor ensembles or adaptive anchor selection during training.
 8 Conclusion
-We present ADPO, a unified framework generalizing DPO to soft preferences, reference anchoring, and listwise settings.
-Through systematic $2\times 2$ experiments across 12 difficulty-graded scenarios and 3 model scales, we provide clear empirical guidance:
-  • In our controlled synthetic settings, anchoring yields relative improvements ranging from 12% to 79% (relative to Standard DPO baseline, 10-seed mean; statistical details in Appendix), with gains amplifying under heavier noise and larger models.
-  • Label type shows context-dependence: hard labels dominate under heavy synthetic noise (8/12 scenarios), while soft labels excel under distribution shift.
-  • Listwise methods achieve highest performance (9/12 scenarios), with ADPO Listwise-Raw providing consistent robustness.
-  • Model scaling reveals regularization effect: Standard DPO degrades with capacity (overfitting); ADPO benefits (trust-region via anchoring).
-Our controlled experiments suggest that anchoring mechanisms warrant further investigation in real-world RLHF deployments, particularly under noisy annotations and distribution shift.
-Reproducibility
-All experiments use fixed random seeds. Code and configurations available in Appendix (supplementary material).
+We introduced Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes DPO to soft, listwise settings.
+We established a key theoretical connection: anchoring imposes an implicit, distribution-space trust region governed by the softmax Fisher metric.
+This geometric insight provides a principled foundation for understanding preference optimization.
+[…]
+These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors.
+By establishing the principle of anchor flexibility, ADPO provides a more robust and versatile tool for aligning models, opening new avenues for adaptive learning strategies in both real-world RL and large-scale language model training.
+While we have validated our framework on controlled bandits and continuous control, applying this task-aware anchoring strategy to large-language model alignment is a critical and promising direction for future work.
+Appendix A Fisher Metric Connection and Relation to KD/TRPO
+Fisher metric expansion.
+For logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$, consider the anchored KL objective
+$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\quad\text{with}\quad\tilde{p}_{\theta}(i)=\frac{\exp(u_{i})}{\sum_{j}\exp(u_{j})}.$
+Expanding around $u^{\star}$ such that $\tilde{p}_{\theta}(\cdot;u^{\star})=q$, the second-order Taylor expansion yields
+$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (7)
+The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family, revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space.
+Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry.
+Relation to KD and TRPO.
+Table 6 summarizes the geometric distinctions.
+While knowledge distillation (KD) minimizes $\mathrm{KL}(q\|p_{\theta})$ to directly align absolute probabilities, it lacks any notion of a trust region and is sensitive to teacher noise.
+Trust-region policy optimization (TRPO) instead constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, explicitly bounding the step size at the cost of solving a constrained subproblem.
+ADPO can be interpreted as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, its anchored KL naturally induces the same Fisher information geometry without requiring an explicit KL penalty.
+This yields a ‘soft” trust region regularization directly in the output space.
+Table 6: Geometric comparison among KD, TRPO, and ADPO.
+Method Optimization space Geometric center Fisher metric Trust region
+KD Probability simplex Teacher $q$ Softmax Fisher None
+TRPO Parameter space Old policy Parameter Fisher Explicit constraint
+ADPO Distribution (anchored) Reference policy Softmax Fisher Implicit (anchored KL)
+Appendix B Proofs and Derivations
+B.1 Proof of Proposition 3.2
+We prove the three special cases.
+(i) Standard DPO (hard labels, no anchoring).
+With $q_{ij}\in\{0,1\}$ and $\Delta_{\mathrm{ref}}=0$, the pairwise ADPO loss
+$\ell_{ij}=\log\!\left(1+e^{\beta\Delta_{\theta}}\right)-q_{ij}\beta\Delta_{\theta}$
+reduces to
+$\ell_{ij}=\begin{cases}-\log\sigma(\beta\Delta_{\theta}),&q_{ij}=1,\\[1.99997pt] -\log\sigma(-\beta\Delta_{\theta}),&q_{ij}=0,\end{cases}$
+which exactly matches the standard DPO objective [1].
+(ii) Non-anchored Bradley–Terry (soft labels, no anchoring).
+With $\Delta_{\mathrm{ref}}=0$ and $q_{ij}\in(0,1)$,
+$\frac{\partial\ell_{ij}}{\partial\Delta_{\theta}}=\beta\big(\sigma(\beta\Delta_{\theta})-q_{ij}\big).$
+Setting the gradient to zero yields $\sigma(\beta\Delta_{\theta})=q_{ij}$, i.e., the Bradley–Terry MLE condition.
+(iii) Reward-based soft labels.
+If $q_{ij}=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, then the stationary point satisfies $\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, so at optimum $\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})=\beta_{r}(R_{i}-R_{j})$ up to the softmax log-odds, linking policy log-ratios to reward gaps.
+∎
+B.2 Proof of Lemma 3.3
+Setup.
+Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\exp(u_{i})/\sum_{j}\exp(u_{j})$.
+Define $A(u)=\log\sum_{j}e^{u_{j}}$ and $\mathcal{L}(u)=A(u)-\langle q,u\rangle$.
+Note $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$.
+Also $\mathcal{L}(u)=\mathrm{KL}\big(q\|\tilde{p}_{\theta}(\cdot;u)\big)+H(q)$.
+Second-order expansion.
+Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q$. Then $\nabla\mathcal{L}(u^{\star})=0$ and
+$\mathcal{L}(u)=\mathcal{L}(u^{\star})+\tfrac{1}{2}(u-u^{\star})^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big](u-u^{\star})+o\!\big(\|u-u^{\star}\|^{2}\big).$
+Centering.
+Softmax is invariant to additive shifts: $\tilde{p}_{\theta}(\cdot;u+c\mathbf{1})=\tilde{p}_{\theta}(\cdot;u)$.
+Hence $\mathrm{Diag}(q)-qq^{\top}$ has a null-space along $\mathbf{1}$.
+Define $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$.
+Then
+$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\mathrm{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
+Back to anchored scores.
+Since $u=(s-s^{\mathrm{ref}})/\tau$, we obtain
+$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
+i.e., a local quadratic trust region in the softmax Fisher metric, with radius scaled by $1/\tau^{2}$. ∎
+B.3 Proof of Lemma 2.1
+Anchored student invariance.
+For any constant $c\in\mathbb{R}$, replace $s_{i}\mapsto s_{i}+c$ and $s_{i}^{\mathrm{ref}}\mapsto s_{i}^{\mathrm{ref}}+c$ for all $i$ in a group $S_{x}$.
+Then $(s_{i}-s_{i}^{\mathrm{ref}})$ is unchanged, hence
+$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\!\big((s_{i}-s_{i}^{\mathrm{ref}})/\tau\big)}{\sum_{j\in S_{x}}\exp\!\big((s_{j}-s_{j}^{\mathrm{ref}})/\tau\big)}$
+is invariant.
+Plackett–Luce teacher invariance.
+Let $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$.
+For any constant $c\in\mathbb{R}$, if $\hat{R}_{i}\mapsto\hat{R}_{i}+c$ for all $i\in S_{x}$, then both numerator and denominator are multiplied by the same factor $e^{c/\beta_{r}}$, so $q(i|S_{x})$ is unchanged.
+Loss invariance.
+Since both $q(\cdot|S_{x})$ and $\tilde{p}_{\theta}(\cdot|S_{x})$ are invariant to groupwise additive shifts, the anchored listwise cross-entropy $\sum_{i}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})$ is invariant as well.
+∎
+Appendix C Implementation and Training Details
+C.1 Noise Generation Procedures
+We provide complete mathematical specifications for the four noise families used in our experiments.
+Heavy Noise (Gaussian with outliers).
+For each item $i$ in group $S_{x}$, the observed noisy reward is:
+$\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i},$
+where $R_{i}^{\star}=f_{\star}(c,v_{i})$ is the ground-truth reward from the oracle MLP $f_{\star}$, and
+$\epsilon_{i}\sim\begin{cases}\mathcal{N}(0,\sigma^{2})&\text{with probability }1-p_{\text{out}},\\ \mathcal{N}(0,\sigma_{\text{out}}^{2})&\text{with probability }p_{\text{out}},\end{cases}$
+with $\sigma_{\text{out}}^{2}=100\sigma^{2}$. The severity levels correspond to:
+  • Light: SNR = $\frac{\mathrm{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
+  • Medium: SNR = 0.5, $p_{\text{out}}=0.10$
+  • Heavy: SNR = 0.2, $p_{\text{out}}=0.20$
+Distribution Shift.
+Training and test distributions differ in the context space. Let $(c,v_{i})$ denote training instances. Test instances are generated as:
+$(c_{\text{test}},v_{i})=(\alpha\cdot c+\delta\cdot\mathbf{1},v_{i}),$
+where $\alpha>1$ scales the context magnitude and $\delta$ adds a constant shift. Severity levels:
+  • Light: $\alpha=1.2$, $\delta=0.3$
+  • Medium: $\alpha=1.5$, $\delta=0.5$
+  • Heavy: $\alpha=2.0$, $\delta=1.0$
+The ground-truth model $f_{\star}$ trained on the original distribution must generalize to these shifted contexts.
+Adversarial Label Flips.
+For pairwise comparisons $(i,j)$, we flip the preference label with probability $p$:
+$y_{ij}^{\text{flip}}=\begin{cases}1-y_{ij}&\text{with probability }p,\\ y_{ij}&\text{with probability }1-p,\end{cases}$
+where $y_{ij}=\mathbb{I}(R_{i}^{\star}>R_{j}^{\star})$ is the clean label. For soft labels, we apply:
+$q_{ij}^{\text{flip}}=\begin{cases}1-q_{ij}&\text{with probability }p,\\ q_{ij}&\text{with probability }1-p.\end{cases}$
+Severity levels: Light (5%), Medium (10%), Heavy (20%).
+Heavy-Tailed Noise (Cauchy).
+The observed reward is:
+$\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i},\quad\epsilon_{i}\sim\text{Cauchy}(0,\gamma),$
+where $\text{Cauchy}(0,\gamma)$ has PDF $f(\epsilon)=\frac{1}{\pi\gamma(1+(\epsilon/\gamma)^{2})}$.
+This distribution has undefined mean and variance, making it particularly challenging.
+Severity levels:
+  • Light: $\gamma=0.3$
+  • Medium: $\gamma=0.5$
+  • Heavy: $\gamma=1.0$
+C.2 Hyperparameters and Training Configuration
+Table 7: Complete hyperparameter configuration for all experiments.
+Parameter Value
+Optimization
+Optimizer AdamW
+Learning rate $5\times 10^{-4}$
+$\beta_{1}$ 0.9
+$\beta_{2}$ 0.999
+Weight decay $10^{-4}$
+Learning rate schedule Constant (no decay)
+Batch size 32
+Training epochs 80
+Model Architecture
+Context dimension $D_{c}$ 8
+Item embedding $D_{v}$ 8
+Small model hidden 64
+Medium model hidden 128
+Large model hidden 256
+Small model layers 2
+Medium model layers 3
+Large model layers 4
+Activation ReLU
+ADPO-specific
+Temperature $\beta$ 1.0
+Temperature $\beta_{r}$ 1.0
+Listwise temperature $\tau$ 1.0 (pairwise), grid-search (listwise)
+Reference pre-train steps 30 (on clean data)
+Experimental (Contextual Bandit)
+Candidate set size $P$ 4
+Number of seeds 10
+Training dataset size 10,000 instances
+Test dataset size 1,000 instances
+Distillation Experiment (MuJoCo)
+Teacher network SAC with 512-512-256 hidden dims
+Student network 256-256 hidden dims
+Action candidates $K$ 8
+Distillation dataset 10,000 states
+Training epochs 200
+Batch size 128
+Number of seeds 5
+Teacher sampling ratio 75% (teacher) + 25% (random)
+C.3 Metrics Definition
+WinMass.
+WinMass is defined as the expected probability mass assigned to the ground-truth best item:
+$\mathrm{WinMass}=\mathbb{E}_{x\sim\mathcal{D}_{\text{test}},S_{x}}\left[\tilde{p}_{\theta}(i^{\star}|S_{x})\right],$
+where $i^{\star}=\operatorname*{arg\,max}_{i\in S_{x}}R_{i}^{\star}$ is the item with the highest ground-truth reward. For random selection with $P=4$ candidates, the baseline WinMass is $1/P=0.25$.
+Appendix D Additional Experiments and Ablations
+D.1 Statistical Significance Tests
+Method.
+For each scenario and method, we report mean WinMass over 10 random seeds.
+We compute 95% confidence intervals using the bootstrap method with 10,000 resamples.
+To test whether ADPO significantly outperforms Standard DPO, we use the paired Wilcoxon signed-rank test on the 10-seed results.
+The null hypothesis is that the median difference is zero.
+We report $p$-values and reject $H_{0}$ at $\alpha=0.05$.
+D.2 Effect of Reference Initialization
+Ablation results.
+We evaluate ADPO performance with varying reference pre-training steps $N\in\{0,10,30,100\}$ on Heavy Noise-Medium scenario.
+Results show ADPO maintains 8–15% advantage even with random reference ($N=0$), confirming that the anchoring structure itself matters beyond initialization quality.
+Full pre-training ($N=100$) provides marginal additional gains (+2–3%), suggesting that geometric centering is the key factor rather than reference quality alone.
+D.3 Temperature Sensitivity (Full Grid)
+Figure 5 in the main text shows three representative methods. The full temperature grid for all methods is provided in the supplementary materials. Key findings:
+  • Pairwise-ADPO maintains WinMass $\in[0.61,0.62]$ across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$;
+  • Listwise-Raw shows similar robustness (WinMass $\in[0.58,0.63]$);
+  • Listwise-KDE benefits from lower $\tau$ (0.5–1.0) but degrades at $\tau>2.0$.
+D.4 KDE-CDF-Logit Transform Details
+Given rewards $\{\tilde{R}_{i}\}_{i=1}^{P}$ in a group, ADPO-KDE constructs targets via:
+  1. 1.
+    KDE fit (Scott’s rule).
+Fit a Gaussian KDE $\hat{f}(r)=\frac{1}{Ph}\sum_{i=1}^{P}K\!\big(\frac{r-\tilde{R}_{i}}{h}\big)$ with $K$ standard normal and bandwidth $h=P^{-1/5}\,\hat{\sigma}$ (Scott’s rule), where $\hat{\sigma}$ is the sample standard deviation.
+  2. 2.
+    CDF evaluation. Compute $\hat{F}(\tilde{R}_{i})=\int_{-\infty}^{\tilde{R}_{i}}\hat{f}(r)\,dr$.
+  3. 3.
+    Logit map with clipping.
+Let $\hat{F}_{\epsilon}(\tilde{R}_{i})=\mathrm{clip}\big(\hat{F}(\tilde{R}_{i}),\,10^{-6},\,1-10^{-6}\big)$ and $\ell_{i}=\log\!\frac{\hat{F}_{\epsilon}(\tilde{R}_{i})}{1-\hat{F}_{\epsilon}(\tilde{R}_{i})}$.
+  4. 4.
+    Temperature-softmax target. Center logits via $\bar{\ell}=\frac{1}{P}\sum_{j}\ell_{j}$ and set
+    $q(i|S_{x})=\frac{\exp\!\big((\ell_{i}-\bar{\ell})/\beta_{r}\big)}{\sum_{j}\exp\!\big((\ell_{j}-\bar{\ell})/\beta_{r}\big)}.$
+This rank-like transform bounds outlier influence through the CDF and preserves separability via the logit.
+Appendix E Reproducibility Checklist
+To facilitate reproduction of our results, we provide the following information:
+Random seeds.
+All experiments use seeds $\{0,1,\ldots,9\}$ for 10-seed runs. Each seed controls:
+  • PyTorch random number generator: torch.manual_seed(seed)
+  • NumPy random number generator: np.random.seed(seed)
+  • Python built-in random: random.seed(seed)
+Dataset generation.
+Synthetic contextual bandits are generated as follows:
+  • Context $c\sim\mathcal{U}[0,1]^{D_{c}}$ with $D_{c}=8$;
+  • Item embeddings $v_{i}\sim\mathcal{U}[0,1]^{D_{v}}$ with $D_{v}=8$;
+  • Ground-truth rewards $R_{i}^{\star}=f_{\star}(\text{concat}(c,v_{i}))$ where $f_{\star}$ is a 2-layer MLP with hidden dimension 64;
+  • Noisy observations $\tilde{R}_{i}$ generated according to Section C.
+Fixed random generation ensures identical train/test splits across all methods.
+Code and configuration.
+Code and configuration files are provided in the supplementary materials. The repository includes:
+  • Implementations of all ADPO variants (pairwise/listwise, soft/hard);
+  • Standard DPO baseline implementation;
+  • Noise generation scripts for all four families;
+  • Experimental configuration YAMLs;
+  • Plotting scripts for all figures;
+  • Pre-computed results for verification.
+Evaluation procedure.
+For each trained model:
+  1. 1.
+    Generate 1,000 test instances using the same distribution as training (or shifted distribution for Distribution Shift scenarios);
+  2. 2.
+    For each instance, sample $P=4$ candidates;
+  3. 3.
+    Compute WinMass as $\mathbb{E}[\tilde{p}_{\theta}(i^{\star}|S_{x})]$ over all test instances;
+  4. 4.
+    Report mean and standard deviation over 10 random seeds.

### v4

@@
 Abstract
-Direct Preference Optimization (DPO) has become a standard for aligning models with human feedback, yet its reliance on hard, pairwise preferences makes it brittle to annotator noise and distribution shift.
-We propose Anchored Direct Preference Optimization (ADPO), a generalized framework that learns from soft, listwise supervision by anchoring policy updates to a reference model.
-Our key theoretical contribution is to show that this anchoring mechanism imposes an implicit trust region on the policy update, enforced by the softmax Fisher information metric.
-This provides a robust geometric interpretation for both fixed and dynamic anchor strategies.
 […]
@@
 This sensitivity arises because DPO’s objective only regularizes the difference in log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
-To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes preference learning to soft and listwise supervision.
-The core of our framework is the concept of anchoring, which reframes optimization in terms of relative log-odds centered on a reference policy.
-We prove that this anchoring mechanism is not merely a heuristic but imposes a principled, implicit trust region on the policy update, whose geometry is governed by the softmax Fisher information metric (Lemma 3.3).
+To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that reveals a fundamental connection across seemingly disparate learning paradigms.
+Through a simple anchored projection formulation—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ where $\tilde{p}_{\theta}$ is defined via anchored logits $(s-s^{\mathrm{ref}})/\tau$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all emerge as special cases through different choices of target distribution $q$, anchor policy $\pi_{\mathrm{ref}}$, and temperature $\tau$ (Section 3, Proposition 3.1).
+The anchoring structure is key: by centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
+We prove that this anchoring mechanism provides a principled geometric framework: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.3), yielding an implicit trust region centered at the target (not the anchor itself).
 This provides a robust, distribution-space analogue to trust-region methods like TRPO [23], without requiring explicit constraints.
-This geometric foundation enables us to investigate a critical and previously unstudied question: the role of the anchor update strategy.
-We provide the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy).
+Crucially, this geometric foundation holds regardless of anchor choice, enabling us to investigate the role of the anchor update strategy.
+To our knowledge, this is the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy) in the context of preference learning and policy optimization.
 Our central finding is a crucial task-dependent trade-off:
@@
 […]
-Collectively, our work makes four key contributions: (1) we propose a unified framework for preference optimization that robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
+Collectively, our work makes four key contributions: (1) we propose ADPO as a unified projection framework that subsumes major learning paradigms (SFT, KD, max-ent RL, DPO) and robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
 2 Related Work
@@
 3 Anchored Direct Preference Optimization
-3.1 Pairwise Soft-DPO
+Notation.
+Throughout this section, we use the following notation:
+  • $x$: context (prompt/state); $S_{x}=\{1,\ldots,|S_{x}|\}$: candidate set (actions/outputs)
+  • $\pi_{\theta}(i\mid x)$: student policy probability for candidate $i$ given context $x$
+  • $\pi_{\mathrm{ref}}(i\mid x)$: reference (anchor) policy, either fixed or dynamic
+  • $s_{i}=\log\pi_{\theta}(i\mid x)$, $s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x)$: log-probabilities
+  • $q(i\mid S_{x})$: target preference distribution over candidates (from teacher, rewards, or labels)
+  • $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$: anchored logits
+  • $\tilde{p}_{\theta}(i\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution
+Table 1: Temperature parameters and their roles across ADPO formulations.
+Symbol Role Where used
+$\beta_{r}$ Teacher/target inverse temperature Forms $q$ from rewards: $q\propto\exp(R/\beta_{r})$
+$\beta$ Student inverse temperature (pairwise) Pairwise logistic: $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))$
+$\tau$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau$ (listwise)
+Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them.
+We set $\beta=\beta_{r}=1.0$ by default.
+However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau}(\tilde{p}-q)$).
+Thus smaller $\beta$ slows pairwise updates; smaller $\tau$ tightens the trust region via curvature $\propto 1/\tau^{2}$ but increases gradient magnitudes.
+Remark (temperature $\tau$ and regularization strength): In listwise settings with anchoring, $\tau$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ (Proposition 3.2).
+Intuitively:
+  • As $\tau\to 0$: $q^{\tau}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
+  • As $\tau\to\infty$: $q^{\tau}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
+Note that $\tilde{p}_{\theta}$ always matches $q$ at optimality regardless of $\tau$; the temperature affects how the underlying policy $\pi_{\theta}$ relates to the anchor.
+Smaller $\tau$ induces a tighter implicit trust region in the sense that large deviations $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are more heavily penalized (since gradients scale as $1/\tau$).
+3.1 ADPO as a Unified Anchored Projection
+We cast preference learning as an anchored projection on the probability manifold: given target $q$ and anchor $\pi_{\mathrm{ref}}$, minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ aligns the student under the softmax Fisher geometry.
+This unified view encompasses supervised fine-tuning, knowledge distillation, maximum-entropy RL, and DPO as special cases.
+Master formulation.
+Given student $\pi_{\theta}$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau>0$, the anchored logits are
+$u_{i}\;=\;\frac{s_{i}-s_{i}^{\mathrm{ref}}}{\tau},\qquad s_{i}=\log\pi_{\theta}(i\mid x),\ \ s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x).$ (3)
+Note (sequence models): In autoregressive sequence models, we use length-normalized log-probabilities (token-average NLL) for $s$ and $s^{\mathrm{ref}}$ to avoid length bias.
+Under this setting, each candidate’s score is the per-token average, and these averaged scores enter the group-level softmax over candidates.
+Since length normalization amounts to an equal rescaling of all scores within a group, it preserves the softmax affine-invariance: the Fisher geometry and anchoring properties remain intact when applied to this sequence-level distribution over candidates (not the token-level distribution).
+Thus the theoretical results—groupwise shift invariance (Lemma 2.1), implicit trust region (Lemma 3.3), and closed-form projection (Proposition 3.2)—carry through in this setting.
+The anchored model distribution over a group $S_{x}$ is
+$\tilde{p}_{\theta}(i\mid S_{x})\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}\;=\;\mathrm{softmax}_{i}\Big(\tfrac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau}\Big).$ (4)
+Given a target preference distribution $q(\cdot\mid S_{x})$ (pairwise or listwise), ADPO minimizes
+$\mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\left[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\right]\;=\;\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right]+\mathrm{const}.$ (5)
+This objective performs a projection onto the anchored probability manifold; its second-order expansion around $q$ yields the softmax Fisher metric, inducing an implicit trust region centered at $\pi_{\mathrm{ref}}$ (formalized in Section 3.4).
+Relation to standard DPO.
+DPO [1] also introduces a reference model, but it enters only through log-odds differences between two candidates:
+$\Delta_{\theta,\mathrm{ref}}=\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{i}}-\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{j}}.$
+In contrast, ADPO anchors each score $s_{i}-s_{i}^{\mathrm{ref}}$ individually before normalizing across the group to form $\tilde{p}_{\theta}$.
+When $|S_{x}|=2$ and $q$ is binary (or Bradley–Terry soft targets), Eq.
+(5) reduces to the logistic preference loss used in DPO.
+For general listwise $|S_{x}|>2$, ADPO retains distribution-level anchoring and inherits a Fisher-metric trust region around $\pi_{\mathrm{ref}}$.
+Proposition 3.1 (ADPO unifies major learning paradigms).
+Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq. (5). Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover:
+  1. 1.
+    Supervised fine-tuning (SFT).
+If $q(\cdot\mid S_{x})$ equals the data label distribution (one-hot or soft), $\pi_{\mathrm{ref}}$ is uniform (constant logits that cancel in softmax), and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss.
+For $\tau\neq 1$, $\tilde{p}_{\theta}$ is a temperature-scaled version of $\pi_{\theta}$.
+  2. 2.
+[…]
+  3. 3.
+    Maximum-entropy RL / offline policy projection.
+    • (Soft-optimal Boltzmann policy) The standard maximum-entropy RL objective yields the soft-optimal policy $\pi^{\star}(a|s)\propto\exp(Q(s,a)/\alpha)$ [19, 21, 20].
+To match this target via ADPO’s anchored distribution $\tilde{p}_{\theta}$, we set $q(a|s)\propto\exp(Q(s,a)/\alpha)$ and use a uniform anchor $\pi_{\mathrm{ref}}$.
+      Parametrization: ADPO constrains the anchored distribution $\tilde{p}(u)=\mathrm{softmax}((\log\pi_{\theta}-\log\pi_{\mathrm{ref}})/\tau)$ to equal $q$.
+Since there are equivalent parameterizations, we choose $\tau=1$ and $\pi_{\theta}(a|s)\propto\pi_{\mathrm{ref}}(a|s)\exp\!\big(Q(s,a)/\alpha\big)$ (with $\pi_{\mathrm{ref}}$ uniform, this reduces to $\pi_{\theta}\propto\exp(Q/\alpha)$).
+This directly recovers the standard soft-optimal policy $\pi^{\star}$ for use in sampling/evaluation, preserving the conventional Boltzmann temperature $1/\alpha$.
+Alternative parameterizations (e.g., $\tau=\alpha$ with $\pi_{\theta}\propto\exp(Q)$) yield the same $\tilde{p}$ but would shift the temperature into the anchoring scale, which we avoid for clarity.
+    • (AWR/AWAC/MPO advantage-weighted projection) When $S_{x}$ is the full action space with overlapping support between $\pi_{\text{old}}$ and the target $q$, the anchor is $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, and the target is $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$ (shift-invariant to constant baselines), with $\tau=\beta$.
+The optimal anchored distribution is $\tilde{p}_{\theta}=q\propto\pi_{\text{old}}\exp(A/\beta)$, exactly matching the advantage-weighted target.
+      Derivation of $\pi_{\theta}$ form: ADPO enforces $\tilde{p}(u)=q$ with $u=(\log\pi_{\theta}-\log\pi_{\text{old}})/\tau$ and $\tau=\beta$. Then
+      $\tilde{p}(u)\ \propto\ \exp\!\Big(\frac{\log\pi_{\theta}-\log\pi_{\text{old}}}{\beta}\Big)\ \stackrel{{\scriptstyle!}}{{\propto}}\ \pi_{\text{old}}\,\exp\!\big(A/\beta\big),$
+      which is equivalent to
+      $\frac{\pi_{\theta}}{\pi_{\text{old}}}\ \propto\ \Big(\pi_{\text{old}}\,e^{A/\beta}\Big)^{\beta}\ \Rightarrow\ \pi_{\theta}\ \propto\ \pi_{\text{old}}^{\,1+\beta}\ e^{A}.$
+      What is exactly recovered?
+ADPO matches the anchored distribution $\tilde{p}$ to the advantage-weighted target $q$; the induced policy shape $\pi_{\theta}\propto\pi_{\text{old}}^{\,1+\beta}\exp(A)$ is a consequence of the anchored projection (Prop.
+3.2) and is not the classic MPO/AWAC M-step update (which yields $\pi_{\theta}\propto\pi_{\text{old}}\exp(A/\beta)$ or directly $\pi_{\theta}=q$ under sufficient capacity).
+Our equivalence is thus at the $\tilde{p}-$level rather than the $\pi_{\theta}$-level.
+Assumptions: (1) full action space $S_{x}$ is enumerable or covered by samples; (2) advantage weights $\exp(A/\beta)$ are normalizable; (3) for off-policy data $d^{\beta}$, we approximate by ignoring or clipping importance ratios (standard in AWR/AWAC [24, 25]), yielding weighted MLE $\max_{\theta}\mathbb{E}_{(s,a)\sim d^{\beta}}[\exp(A/\beta)\log\pi_{\theta}(a|s)]$.
+MPO’s E-step weights $w\propto\exp(Q/\eta)$ with dual variable $\eta$ (enforcing a KL constraint via Lagrangian duality) follow the same projection form [26], though MPO additionally solves for $\eta$ to satisfy a KL budget.
+  4. 4.
+    DPO (binary comparison form).
+For $|S_{x}|=2$ with candidates $\{i,j\}$ and Bradley–Terry soft targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$ where $\Delta R=R_{i}-R_{j}$ is the reward difference, Eq.
+(5) is equivalent to DPO up to temperature scaling and reference shift.
+Specifically, the Bayes-optimal condition $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))=q_{ij}$ yields $\Delta_{\theta}-\Delta_{\mathrm{ref}}=(\beta_{r}/\beta)\Delta R$.
+When $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform (so $\Delta_{\mathrm{ref}}=0$), this is exactly identical to the original DPO objective; otherwise, it differs by a temperature rescaling $\beta_{r}/\beta$ and an anchor-induced shift $\Delta_{\mathrm{ref}}$.
+Table 2: ADPO as a unifying projection framework.
+Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover major learning paradigms.
+Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau)$ to the target $q$; the corresponding student policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ when the optimum is reachable (Prop.
+3.2).
+Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau$ Recovers ($\tilde{p}_{\theta}$ / $\pi_{\theta}$)
+SFT data (one-hot/soft) uniform 11 cross-entropy / same
+KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}\!\big(q_{T},\mathrm{softmax}(s/T)\big)$ / train at $T$, infer T=1T{=}1
+KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}((s-s_{T})/\tau)\big)$ / trust region
+[…]
+[…]
+DPO (binary) Bradley–Terry / hard uniform any original DPO (pairwise)
+ADPO (listwise) Plackett–Luce / soft any (teacher/old) any this work
+Proposition 3.2 (Closed-form optimal solution).
+Suppose $S_{x}$ covers the full action/class space and $q(\cdot\mid S_{x})\succ 0$.
+(Discrete case) If the policy $\pi_{\theta}$ has sufficient expressivity (i.e., the parametrization can represent arbitrary log-odds over $S_{x}$), then the distribution-level optimum of $\operatorname*{arg\,min}_{\theta}\mathrm{KL}(q\|\tilde{p}_{\theta})$ satisfies
+$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau}.$ (6)
+(Continuous case) For continuous action spaces, assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure (e.g., Lebesgue measure).
+Define the anchored distribution via the density ratio:
+$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau}.$
+Setting $\tilde{p}_{\theta}=q$ yields the closed-form solution
+$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau},$
+which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau}\,da<\infty$ to ensure normalizability.
+Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
+Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau}$.
+The constant $c$ is absorbed into the proportionality constant.
+∎
+Implication and reachability.
+Equation (6) strictly characterizes the "anchored projection" as a multiplicative fusion of reference and target when the optimum is reachable.
+In discrete finite-action settings with tabular or sufficiently expressive function approximation, this optimum is exactly attainable.
+However, in structured settings (e.g., autoregressive language models with shared parameters across tokens, or policy networks with limited capacity), the optimum may be unreachable, and gradient descent performs an approximate projection in the parametric family.
+The KD, MaxEnt, and AWR special cases in Proposition 3.1 all follow directly from (6).
+Listwise subset constraints and convexity.
+When $S_{x}$ is only a subset (pairwise/listwise ranking), (6) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints.
+Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau$ within each group (as shown in Theorem 3.6), but the optimization over shared parameters $\theta$ is typically non-convex.
+The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ for all groups (except in the tabular case or with unlimited expressivity).
+Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
+By centering the coordinate system at $\pi_{\mathrm{ref}}$ before applying softmax, ADPO provides groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—properties absent in standard KD or unanchored DPO.
+Table 2 summarizes the unification.
+Standing assumptions.
+Throughout this work, we maintain the following conditions unless otherwise noted:
+  1. (i)
+    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i|x)>0\Rightarrow\pi_{\mathrm{ref}}(i|x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ are well-defined.
+  2. (ii)
+    Absolute continuity (continuous): For continuous action spaces, we assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure, with overlapping support in the same sense as (i).
+  3. (iii)
+    Local analysis vs.
+global bounds: Fisher-metric trust region results (Lemma 3.3, Corollary 3.4) hold locally around $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$.
+For global strong convexity, we either work in a neighborhood of $u^{\star}$ or assume the optimization is constrained to a bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ (Theorem 3.6).
+  4. (iv)
+    Sequence models: For autoregressive sequence generation, we use length-normalized (per-token average) log-probabilities $s$ and $s^{\mathrm{ref}}$ as group-level scores, and apply the softmax Fisher geometry at the sequence-level distribution over candidates.
+  5. (v)
+    Expressivity: Closed-form optimal solutions (Proposition 3.2) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau}$.
+In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
+3.2 Pairwise Instantiation
+Notation for this subsection: $\beta$ = student inverse temperature, $\beta_{r}$ = teacher/reference inverse temperature, $q_{ij}$ = soft preference probability that $i\succ j$.
 For a candidate pair $(i,j)$ with soft teacher preference $q_{ij}\in(0,1)$ that $i$ is preferred, the ADPO pairwise loss is:
@@
   • Reference anchoring: Relative updates $\Delta_{\theta}-\Delta_{\text{ref}}$ are invariant to groupwise shifts.
-Remark 3.1 (Scale non-identifiability).
-Temperatures $\beta$ and $\beta_{r}$ only appear through ratio $\beta_{r}/\beta$ in the optimum. We set $\beta=\beta_{r}=1.0$ by default.
-Proposition 3.2 (Special cases).
-Soft-DPO recovers: (i) standard DPO when $q_{ij}\in\{0,1\}$ and $\Delta_{\text{ref}}=0$; (ii) non-anchored Bradley–Terry when $\Delta_{\text{ref}}=0$; (iii) reward-based soft preferences when $q_{ij}=\sigma(\beta_{r}(R_{i}-R_{j}))$.
-Implication.
-This shows that ADPO is a strict generalization of standard DPO, and the anchoring term $(s-s^{\mathrm{ref}})$ reduces to unanchored log-odds when $\Delta_{\text{ref}}=0$.
-Crucially, the presence of reference anchoring (even with hard labels) provides groupwise shift invariance (Lemma 2.1), which standard DPO lacks.
-Proof of (i).
-For $q_{ij}=1$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})-\beta\Delta_{\theta}=-\log\sigma(\beta\Delta_{\theta})$.
-For $q_{ij}=0$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})=-\log\sigma(-\beta\Delta_{\theta})$.
-This matches standard DPO exactly.
-∎
-3.2 Listwise Soft-DPO
+3.3 Listwise Instantiation
+Notation for this subsection: $\tau$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
 For group $S_{x}=\{\tau_{1},\ldots,\tau_{P}\}$, the ADPO listwise loss is:
@@
 Lemma 3.3 (Implicit trust region; quadratic form at $p=q$).
-Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\frac{e^{u_{i}}}{\sum_{j}e^{u_{j}}}$.
-Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$ and define the $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$.
+Let $u=(s-s^{\mathrm{ref}})/\tau$ and $\tilde{p}(u)=\mathrm{softmax}(u)$.
+Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$ (the softmax is shift-invariant).
+Define $\delta=u-u^{\star}$ and center it by $q$ as $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$.
 Then, as $u\to u^{\star}$,
-$\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}(\cdot;u)\big)\;=\;\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta\;+\;o(\|\delta\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{q}\!\big[s-s^{\mathrm{ref}}\big]\;+\;o(\cdot),$
-where the variance is taken with respect to $q$ and the quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ (which is invariant to adding constants to all logits).
+$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\tau\log q_{i}\big]\;+\;o(\cdot),$
+where the variance is taken with respect to $q$, measuring the deviation of the anchored log-odds from the target $q$.
+The quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which has a null-space along $\mathbf{1}$ (reflecting shift-invariance).
 Proof sketch.
@@
 ∎
+Corollary 3.4 (Local bounds).
+Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\mathrm{Diag}(q)-qq^{\top}$ (the null-space is along $\mathbf{1}$). Then locally around $u^{\star}$,
+$\mathrm{KL}(q\|\tilde{p}(u))\;\geq\;\frac{\mu(q)}{2}\,\|\bar{\delta}\|^{2},$
+where $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$ is the $q$-centered deviation.
+By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$, providing a distributional deviation bound.
+On the smallest nonzero eigenvalue: For $K=2$, $\mu(q)=2q(1-q)$.
+For $K>2$, $\mu(q)$ depends on the full spectrum of $q$; a uniform, distribution-independent lower bound is generally loose.
+When $\min_{i}q_{i}\geq\varepsilon>0$, the smallest nonzero eigenvalue satisfies $\mu(q)\geq\varepsilon$.
+Proof sketch: For any $v\in\mathbf{1}^{\perp}$, the Rayleigh quotient gives $v^{\top}(\mathrm{Diag}(q)-qq^{\top})v=\sum_{i}q_{i}v_{i}^{2}-(\sum_{i}q_{i}v_{i})^{2}=\mathrm{Var}_{q}[v]\geq\varepsilon\|v\|^{2}$ since $\sum_{i}v_{i}=0$ and the variance of $v$ under any distribution bounded below by $\varepsilon$ is at least $\varepsilon\|v\|^{2}$.
 Geometric interpretation.
-ADPO performs optimization in the logit space centered at the reference policy: the coordinate origin is fixed at $s^{\mathrm{ref}}$, so updates are measured as relative log-odds $(s-s^{\mathrm{ref}})$.
-Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” around the reference anchor.
+ADPO performs optimization in the anchored $u$-coordinate system: the reference $\pi_{\mathrm{ref}}$ defines the coordinate origin via $u=(s-s^{\mathrm{ref}})/\tau$, but the center of the local quadratic bowl is at $u^{\star}$ satisfying $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$, not at the anchor itself.
+The local curvature at $u^{\star}$ is governed by the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which depends only on the target $q$.
+Thus: the anchor determines the coordinate system, while the target $q$ determines both the minimizer $u^{\star}$ and the local geometry.
 This contrasts with:
-  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex without a fixed geometric center, making it sensitive to teacher noise.
+  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex.
+While the Fisher curvature exists at any point, there is no fixed reference anchor, so the geometry lacks a stable coordinate origin.
   • TRPO: Constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in parameter space, defining a moving local quadratic region that requires explicit constraint optimization.
-  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed geometric center, providing an implicit trust region without explicit constraints.
+  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed coordinate origin, inducing an implicit trust region without explicit constraints.
 Figure 1: ADPO’s anchored trust region.
-The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space centered at the reference anchor.
-The black trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift.
+The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself.
+The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift.
 This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
-Anchor flexibility and update strategies.
-Crucially, Lemma 3.3 holds regardless of anchor choice or update strategy.
-The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region, whether the anchor is:
-  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable geometric center for consistent target tracking;
-  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive trust region similar to TRPO’s old-policy mechanism;
+3.5 Geometric Interpretation: KL-Mirror Descent in $u$-Space
+Recall $A(u)=\log\!\sum_{j}\exp(u_{j})$ and $\tilde{p}(u)=\nabla A(u)=\mathrm{softmax}(u)$, and our anchored objective $\mathcal{L}(u)=A(u)-\langle q,u\rangle=\mathrm{KL}\!\big(q\;\|\;\tilde{p}(u)\big)+H(q)$.
+Let $u_{t}$ be the current anchored logits and $g_{t}=\nabla\mathcal{L}(u_{t})=\tilde{p}(u_{t})-q$.
+Theorem 3.5 (KL-mirror descent / Bregman-prox step in $u$-space).
+If one were to directly optimize $\mathcal{L}$ over the anchored logits $u$ (rather than shared parameters $\theta$), then for any step size $\eta>0$, the one-step update can be written equivalently as the solution of the following proximal subproblem:
+$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle g_{t},\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big).$ (11)
+Equivalently, since $D_{A}(u\|u_{t})=A(u)-A(u_{t})-\langle\nabla A(u_{t}),u-u_{t}\rangle=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$, (11) is a standard mirror descent / Bregman-prox step with generator $A$.
+Proof.
+Linearize $\mathcal{L}$ at $u_{t}$: $\mathcal{L}(u)\approx\mathcal{L}(u_{t})+\langle g_{t},u-u_{t}\rangle$.
+A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$.
+Using $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (11).
+∎∎
+Dual form in $p$-space. Via the Legendre map $p=\nabla A(u)$, the step (11) is equivalent to the entropic proximal in probability space
+$\tilde{p}_{t+1}=\arg\min_{p\in\Delta}\ \langle g_{t},p\rangle+\tfrac{1}{\eta}\mathrm{KL}\!\big(p\,\|\,\tilde{p}_{t}\big),$
+whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta g_{t})$.
+Multiplicative-weights view.
+The solution of (11) is equivalent to the multiplicative-weights update in the anchored distribution:
+$\tilde{p}_{t+1}\ \propto\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,g_{t}\big)\ =\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,(\tilde{p}_{t}-q)\big),$ (12)
+followed by renormalization. This exhibits ADPO as a KL-trust-region step in distribution space.
+Step-size $\eta$ vs temperature $\tau$ (heuristic).
+Anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$ rescale curvature by $1/\tau^{2}$ (Lemma 3.3).
+To maintain comparable “trust-region tightness” in the unanchored score space $s$, one may heuristically choose
+[…]
+This is a practical guideline; a rigorous derivation would require accounting for the chain rule $\partial u/\partial\theta$ when optimizing over shared parameters $\theta$.
+3.6 Dynamic Anchor Stability
+Crucially, Lemma 3.3 holds regardless of anchor choice.
+The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region.
+While the coordinate $u=(s-s^{\mathrm{ref}})/\tau$ depends on the anchor, the local quadratic form at the optimum only depends on $q$; thus the curvature (Fisher metric) is anchor-agnostic, albeit the optimization objective itself changes with the anchor.
+Common anchor choices include:
+  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable coordinate origin for consistent target tracking;
+  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive coordinate system similar to TRPO’s old-policy mechanism;
   • Teacher anchor: A stronger teacher $\pi_{\text{teacher}}$ for offline distillation (typically fixed).
-Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining stable targets).
+For dynamic anchors, we characterize the monotonic improvement within each fixed-anchor phase:
+Theorem 3.6 (Fixed-anchor-phase monotonic descent in $u$-space).
+Consider a training phase where the anchor $\pi_{\mathrm{ref}}$ is held fixed.
+Define $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ and $u=(s-s^{\mathrm{ref}})/\tau$ are the anchored logits.
+When directly optimizing $\mathcal{L}$ over $u$ (e.g., tabular setting or each group having independent degrees of freedom), $\mathcal{L}$ is strictly convex on the subspace $\mathbf{1}^{\perp}$ (orthogonal to the all-ones vector) with Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite on $\mathbf{1}^{\perp}$ with one zero eigenvalue along $\mathbf{1}$.
+Strong convexity (local or bounded-domain):
+  • Local: In a neighborhood of $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ approaches $\mathrm{Diag}(q)-qq^{\top}$ up to $o(1)$, so there exists $\mu>0$ (depending on $q$) such that $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$ in this neighborhood.
+  • Bounded domain: If the optimization is constrained to a domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$ and some $\varepsilon>0$ (i.e., the current anchored distribution $\tilde{p}_{\theta}(u)$ is bounded away from the probability simplex boundary), then $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$).
+Under either condition, standard gradient descent with appropriate step size $\eta\in(0,1/L]$ (where $L$ is a local smoothness constant) yields linear (geometric) convergence:
+$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\,(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})),$
+implying monotonic decrease in $\mathrm{KL}(q\|\tilde{p}_{\theta})$ within each fixed-anchor phase.
+Note: When the anchor is updated (e.g., $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the objective $\tilde{p}_{\theta}$ is re-normalized with the new anchor, so the loss function changes.
+We do not claim monotonicity across anchor updates; instead, each fixed-anchor phase provides stable descent, and the anchor update adapts the coordinate system for the next phase.
+Caveat for parametrized models: In practice, the anchored logits $u$ are derived from shared parameters $\theta$ (e.g., neural networks), and the optimization problem over $\theta$ is typically non-convex.
+The above results provide geometric intuition for the loss landscape in the anchored distribution space but do not constitute convergence guarantees for parameterized training.
+Proof sketch.
+The log-partition function $A(u)$ is strictly convex (its Hessian is the softmax Fisher metric $\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite with null-space along $\mathbf{1}$).
+The Hessian depends on the current $\tilde{p}_{\theta}(u)$, not directly on $q$.
+However, near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, it coincides with $\mathrm{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity.
+Alternatively, if $\tilde{p}_{\theta}(u)$ is constrained to $\varepsilon,1-\varepsilon]^{K}$, the minimum eigenvalue is uniformly bounded below.
+Standard strongly convex optimization theory then guarantees linear convergence.
+When the anchor changes, $u$ is redefined, so objectives from different phases are not directly comparable.
+∎∎
+Implication.
+This theorem explains why dynamic anchors (Section 5.4) maintain stability during online exploration: within each fixed-anchor phase, the implicit Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts to track policy evolution, similar to TRPO’s mechanism.
+Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting coordinates to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining a stable coordinate origin throughout).
 Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
@@
 4 Experimental Setup
+Evaluation metrics.
+We use three primary metrics to assess preference learning quality:
+  • WinMass: For each context $x$ with candidate set $S_{x}=\{1,\ldots,P\}$ and ground-truth ranking (higher reward is better), WinMass measures the total probability mass assigned to the top item under the anchored distribution: $\mathrm{WinMass}=\mathbb{E}_{x}[\tilde{p}_{\theta}(\operatorname*{arg\,max}_{i}R_{i}^{\star}\mid S_{x})]$.
+Higher is better; random baseline = $1/P$.
+  • NDCG@k: Normalized Discounted Cumulative Gain at position $k$, a standard ranking metric that rewards placing high-reward items at top positions with logarithmic discounting.
+Higher is better; range $[0,1]$.
+  • KL divergence: For distillation tasks, we measure $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\mathrm{student}})$, the expected KL divergence from teacher to student policy over states and actions.
+Lower is better, indicating closer alignment with the teacher.
 $2\times 2$ base design + listwise extensions. We systematically compare:
@@
 Underline: best among pairwise.
-”Best Listwise” shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank.
+"Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank.
 In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
@@
 Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely.
-Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
+Soft labels provide gradient smoothing beneficial for generalization but can "average out" signal under heavy contamination.
 5.3 Model Scaling
@@
 We hypothesize that these strategies serve different purposes: dynamic anchors suit online exploration (where the policy evolves while learning from noisy data), while fixed anchors suit offline distillation (where the policy learns from a stable teacher).
-Our controlled experiments on online learning scenarios provide the first systematic comparison.
+Our controlled experiments on online learning scenarios provide a systematic empirical comparison.
 Experimental setup.
@@
 To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$.
-Figure 5 shows the results.
+Figure˜5 shows the results.
 Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$.
@@
 In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring.
-Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal.
+Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that "average out" signal.
 However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence.
@@
 Comparing frozen vs.
-EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate ”anchoring mechanism” from ”better initialization.”
+EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate "anchoring mechanism" from "better initialization."
   • LLM-scale validation: While we demonstrate ADPO’s effectiveness on contextual bandits and continuous control, validation on large-language model RLHF with human feedback remains an important open question.
@@
 8 Conclusion
-We introduced Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes DPO to soft, listwise settings.
-We established a key theoretical connection: anchoring imposes an implicit, distribution-space trust region governed by the softmax Fisher metric.
-This geometric insight provides a principled foundation for understanding preference optimization.
+We introduced Anchored Direct Preference Optimization (ADPO), a unified projection framework that reveals fundamental connections across major learning paradigms.
+By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(s-s^{\mathrm{ref}})/\tau$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature.
+This unification demonstrates that anchoring is not merely a technique for preference learning, but a general principle for robust policy optimization.
 […]
-These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors.
-By establishing the principle of anchor flexibility, ADPO provides a more robust and versatile tool for aligning models, opening new avenues for adaptive learning strategies in both real-world RL and large-scale language model training.
-While we have validated our framework on controlled bandits and continuous control, applying this task-aware anchoring strategy to large-language model alignment is a critical and promising direction for future work.
+These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors; for maximum-entropy RL, use uniform anchors; for advantage-weighted updates, use old-policy anchors.
+By establishing ADPO as a unifying framework with principled anchor selection, we provide a versatile tool for diverse learning scenarios—from noisy preference optimization to stable knowledge distillation.
+While we have validated our framework on controlled bandits and continuous control, applying this unified anchoring perspective to large-language model alignment, where different training stages (SFT, RLHF, continual learning) may benefit from different anchor strategies, is a critical and promising direction for future work.
 Appendix A Fisher Metric Connection and Relation to KD/TRPO
-Fisher metric expansion.
+Fisher metric expansion and Bregman divergence perspective.
 For logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$, consider the anchored KL objective
 $\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\quad\text{with}\quad\tilde{p}_{\theta}(i)=\frac{\exp(u_{i})}{\sum_{j}\exp(u_{j})}.$
-Expanding around $u^{\star}$ such that $\tilde{p}_{\theta}(\cdot;u^{\star})=q$, the second-order Taylor expansion yields
+Define the convex function $F(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ is the log-partition function. Then the anchored cross-entropy can be written as a Bregman divergence:
+$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\;=\;F(u)-F(u^{\star})-\langle\nabla F(u^{\star}),u-u^{\star}\rangle,$ (14)
+where $u^{\star}$ satisfies $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ (equivalently, $\nabla F(u^{\star})=0$).
+The Bregman divergence (14) is always non-negative, induces a three-point identity (Pythagorean theorem for Bregman divergences), and provides a unified view of mirror descent and anchored projection updates.
+Expanding around $u^{\star}$, the second-order Taylor expansion yields
 $\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (15)
-The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family, revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space.
-Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry.
+The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space.
+Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry, with connections to mirror descent and Bregman proximal methods.
 Relation to KD and TRPO.
@@
 B.1 Proof of Proposition 3.1
-We prove the three special cases.
-(i) Standard DPO (hard labels, no anchoring).
-With $q_{ij}\in\{0,1\}$ and $\Delta_{\mathrm{ref}}=0$, the pairwise ADPO loss
-$\ell_{ij}=\log\!\left(1+e^{\beta\Delta_{\theta}}\right)-q_{ij}\beta\Delta_{\theta}$
-reduces to
-$\ell_{ij}=\begin{cases}-\log\sigma(\beta\Delta_{\theta}),&q_{ij}=1,\\[1.99997pt] -\log\sigma(-\beta\Delta_{\theta}),&q_{ij}=0,\end{cases}$
-which exactly matches the standard DPO objective [1].
-(ii) Non-anchored Bradley–Terry (soft labels, no anchoring).
-With $\Delta_{\mathrm{ref}}=0$ and $q_{ij}\in(0,1)$,
-$\frac{\partial\ell_{ij}}{\partial\Delta_{\theta}}=\beta\big(\sigma(\beta\Delta_{\theta})-q_{ij}\big).$
-Setting the gradient to zero yields $\sigma(\beta\Delta_{\theta})=q_{ij}$, i.e., the Bradley–Terry MLE condition.
-(iii) Reward-based soft labels.
-If $q_{ij}=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, then the stationary point satisfies $\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, so at optimum $\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})=\beta_{r}(R_{i}-R_{j})$ up to the softmax log-odds, linking policy log-ratios to reward gaps.
-∎
+We prove that ADPO unifies major learning paradigms through different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$.
+(i) Supervised fine-tuning (SFT).
+When $\pi_{\mathrm{ref}}$ is uniform (constant logits $s_{i}^{\mathrm{ref}}=c$ for all $i$), the anchored distribution becomes
+$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp((s_{i}-c)/\tau)}{\sum_{j}\exp((s_{j}-c)/\tau)}=\frac{\exp(s_{i}/\tau)}{\sum_{j}\exp(s_{j}/\tau)}=\mathrm{softmax}_{i}(s_{i}/\tau).$
+If $q$ is the data label distribution (one-hot or soft) and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss.
+For $\tau\neq 1$, the temperature acts as a scaling parameter on the logits (temperature scaling).
+(ii) Knowledge distillation (KD).
+If $q=\pi_{T}$ (teacher) and $\pi_{\mathrm{ref}}$ is uniform, then with a training temperature $T>1$ we set $\tau=T$ and minimize
+$\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}\!\big[\mathrm{KL}\big(\pi_{T}\;\|\;\mathrm{softmax}(s_{\theta}/T)\big)\big],$
+and multiply the loss by $T^{2}$ to recover the classic distillation objective $T^{2}\,\mathrm{CE}\!\big(\pi_{T},\mathrm{softmax}(s_{\theta}/T)\big)$ [22].
+Note that $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}(s_{\theta}/T)\big)=\mathrm{KL}(\pi_{T}\|\pi_{\theta})$ only when $T=1$.
+If instead $\pi_{\mathrm{ref}}=\pi_{T}$ (anchored KD), we obtain
+$\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}\!\big[\mathrm{KL}\big(\pi_{T}\;\|\;\mathrm{softmax}\!\big((s_{\theta}-s_{T})/\tau\big)\big)\big],$
+which performs optimization in teacher-anchored coordinates with an implicit Fisher-metric trust region (see Lemma 3.3).
+(iii) Maximum-entropy RL / offline policy projection.
+For $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and uniform $\pi_{\mathrm{ref}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
+$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}(s_{a}/\tau)\propto\exp(Q(s,a)/\alpha),$
+so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=\alpha$ gives the soft-optimal Boltzmann form $\pi\propto\exp(Q/\alpha)$.
+For advantage-weighted updates with $q(a\mid s)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta)$ and $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
+$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}\!\left(\frac{s_{a}-s_{a}^{\mathrm{old}}}{\tau}\right)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta),$
+so $(s_{a}-s_{a}^{\mathrm{old}})/\tau\approx A^{\pi_{\mathrm{old}}}(s,a)/\beta$, giving $\pi_{\theta}(a\mid s)\propto\pi_{\mathrm{old}}(a\mid s)\exp((\tau/\beta)A^{\pi_{\mathrm{old}}}(s,a))$.
+Choosing $\tau=\beta$ recovers AWR/AWAC/MPO projections.
+(iv) DPO (binary comparison form).
+For $|S_{x}|=2$ with candidates $\{i,j\}$ and binary $q(i)=1,q(j)=0$ (or Bradley–Terry soft targets $q(i)=\sigma(\beta_{r}\Delta R)$), the ADPO loss becomes
+$\mathcal{L}_{\mathrm{ADPO}}=-q(i)\log\tilde{p}_{\theta}(i)-q(j)\log\tilde{p}_{\theta}(j),$
+where $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau)$ with $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$.
+For uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$) and hard labels, this reduces to
+$\mathcal{L}_{\mathrm{ADPO}}=-\log\sigma(\Delta_{\theta}/\tau),$
+which matches the original DPO objective with inverse temperature $1/\tau$. ∎
 B.2 Proof of Lemma 3.3

### v5

@@
 However, DPO’s reliance on hard, pairwise preference labels makes it brittle.
-In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our proposed method (0.472 vs.
+In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our method (0.472 vs.
 0.809 WinMass).
-This sensitivity arises because DPO’s objective only regularizes the difference in log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
-To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that reveals a fundamental connection across seemingly disparate learning paradigms.
-Through a simple anchored projection formulation—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ where $\tilde{p}_{\theta}$ is defined via anchored logits $(s-s^{\mathrm{ref}})/\tau$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all emerge as special cases through different choices of target distribution $q$, anchor policy $\pi_{\mathrm{ref}}$, and temperature $\tau$ (Section 3, Proposition 3.1).
-The anchoring structure is key: by centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
-We prove that this anchoring mechanism provides a principled geometric framework: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.3), yielding an implicit trust region centered at the target (not the anchor itself).
-This provides a robust, distribution-space analogue to trust-region methods like TRPO [23], without requiring explicit constraints.
-Crucially, this geometric foundation holds regardless of anchor choice, enabling us to investigate the role of the anchor update strategy.
-To our knowledge, this is the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy) in the context of preference learning and policy optimization.
-Our central finding is a crucial task-dependent trade-off:
-  • A dynamic anchor excels at online exploration in noisy environments, where its adaptive trust region improves performance by 5–11% in 3 out of 4 scenarios.
+This sensitivity arises because the DPO objective only regularizes differences of log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
+We introduce Anchored Direct Preference Optimization (ADPO), a unified framework that connects seemingly disparate learning paradigms.
+Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3, Proposition 3.2).
+By centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
+We further show that this anchoring induces a principled geometric structure: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.5), yielding an implicit trust region centered at the target (not at the anchor).
+This offers a distribution-space analogue to TRPO [23] without explicit constraints.
+Because the geometry depends only on $q$, we can cleanly compare anchor update strategies.
+Under our experimental settings we find a task-dependent trade-off:
+  • Dynamic anchors suit online exploration in noisy environments, improving performance by 5–11% in 3/4 scenarios on our benchmarks.
 […]
@@
 Relation to KD and TRPO.
-Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the Euclidean probability space.
+Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the (unanchored) probability simplex.
 While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts.
 In contrast, trust-region policy optimization (TRPO) [23] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates.
-Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in a coordinate system whose origin is fixed at the reference.
-This anchoring keeps the Fisher metric centered and effectively forms a “quadratic bowl” around the reference distribution, yielding an implicit trust region without any explicit KL constraint.
-To our knowledge, no prior work has formulated such an anchored cross-entropy that maintains a fixed geometric center in probability space.
+Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in coordinates centered at the reference, while the local quadratic well (Fisher metric) is centered at the target distribution $q$, yielding an implicit trust region without an explicit KL constraint.
+We are not aware of prior work that formulates an anchored cross-entropy maintaining a fixed geometric center in probability space.
 Entropy-regularized grounding and its link to ADPO.
 Under the maximum-entropy RL objective
-$\max_{\pi}\ \mathbb{E}_{x\sim\mathcal{D}}\Big[\mathbb{E}_{\tau\sim\pi(\cdot|x)}r(x,\tau)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|x)\big)\Big],$
+$\max_{\pi}\ \mathbb{E}_{s\sim\mathcal{D}}\Big[\mathbb{E}_{a\sim\pi(\cdot|s)}r(s,a)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|s)\big)\Big],$
 the optimal policy admits the Boltzmann form [19, 21]
-$\pi^{*}(\tau|x)\;=\;\frac{1}{Z(x)}\exp\!\Big(\tfrac{1}{\alpha}\,r(x,\tau)\Big),$ (1)
+$\pi^{*}(a|s)\;=\;\frac{1}{Z(s)}\exp\!\Big(\tfrac{1}{\alpha}\,r(s,a)\Big),$ (1)
 and the reward can be reparameterized by the (optimal) policy log-probability
-$r(x,\tau)\;=\;\alpha\,\log\pi^{*}(\tau|x)\;+\;\alpha\,\log Z(x),$ (2)
-so any relative quantity (pairwise/listwise) cancels $Z(x)$.
+$r(s,a)\;=\;\alpha\,\log\pi^{*}(a|s)\;+\;\alpha\,\log Z(s),$ (2)
+so any relative quantity (pairwise/listwise) cancels $Z(s)$.
 Eqs.
-(1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets.
-A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs.
+(1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets.
+A full derivation of (1) as the minimizer of $\mathbb{E}_{s}\,\mathrm{KL}\big(\pi(\cdot|s)\,\|\,\pi^{*}(\cdot|s)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs.
 (3)–(6), (20)–(27)).
 Lemma 2.1 (Groupwise shift invariance).
-If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance.
+For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance.
 See [18], Prop.
@@
 Notation.
-Throughout this section, we use the following notation:
-  • $x$: context (prompt/state); $S_{x}=\{1,\ldots,|S_{x}|\}$: candidate set (actions/outputs)
-  • $\pi_{\theta}(i\mid x)$: student policy probability for candidate $i$ given context $x$
-  • $\pi_{\mathrm{ref}}(i\mid x)$: reference (anchor) policy, either fixed or dynamic
-  • $s_{i}=\log\pi_{\theta}(i\mid x)$, $s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x)$: log-probabilities
-  • $q(i\mid S_{x})$: target preference distribution over candidates (from teacher, rewards, or labels)
-  • $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$: anchored logits
-  • $\tilde{p}_{\theta}(i\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution
+$x$: context (prompt/state); $S_{x}=\{y_{1},\ldots,y_{|S_{x}|}\}$: candidate set (actions/outputs).
+$\pi_{\theta}(y\mid x)$: student policy; $\pi_{\mathrm{ref}}(y\mid x)$: anchor (fixed or dynamic).
+$s_{y}=\log\pi_{\theta}(y\mid x)$, $s^{\mathrm{ref}}_{y}=\log\pi_{\mathrm{ref}}(y\mid x)$: log-probabilities.
+$q(y\mid S_{x})$: target preference distribution (from teacher, rewards, or labels).
+$u_{y}=(s_{y}-s^{\mathrm{ref}}_{y})/\tau_{\mathrm{anc}}$: anchored logits.
+$\tilde{p}_{\theta}(\cdot\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution.
 Table 1: Temperature parameters and their roles across ADPO formulations.
@@
 $\beta$ Student inverse temperature (pairwise) Pairwise logistic: $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))$
-$\tau$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau$ (listwise)
+$\tau_{\mathrm{anc}}$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ (listwise)
 Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them.
 We set $\beta=\beta_{r}=1.0$ by default.
-However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau}(\tilde{p}-q)$).
-Thus smaller $\beta$ slows pairwise updates; smaller $\tau$ tightens the trust region via curvature $\propto 1/\tau^{2}$ but increases gradient magnitudes.
-Remark (temperature $\tau$ and regularization strength): In listwise settings with anchoring, $\tau$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ (Proposition 3.2).
+However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau_{\mathrm{anc}}$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau_{\mathrm{anc}}}(\tilde{p}-q)$).
+Thus smaller $\beta$ slows pairwise updates; smaller $\tau_{\mathrm{anc}}$ tightens the trust region via curvature $\propto 1/\tau_{\mathrm{anc}}^{2}$ but increases gradient magnitudes.
+Remark (temperature $\tau_{\mathrm{anc}}$ and regularization strength): In listwise settings with anchoring, $\tau_{\mathrm{anc}}$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ (Proposition 3.4).
 Intuitively:
-  • As $\tau\to 0$: $q^{\tau}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
-  • As $\tau\to\infty$: $q^{\tau}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
-Note that $\tilde{p}_{\theta}$ always matches $q$ at optimality regardless of $\tau$; the temperature affects how the underlying policy $\pi_{\theta}$ relates to the anchor.
-Smaller $\tau$ induces a tighter implicit trust region in the sense that large deviations $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are more heavily penalized (since gradients scale as $1/\tau$).
-3.1 ADPO as a Unified Anchored Projection
-We cast preference learning as an anchored projection on the probability manifold: given target $q$ and anchor $\pi_{\mathrm{ref}}$, minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ aligns the student under the softmax Fisher geometry.
-This unified view encompasses supervised fine-tuning, knowledge distillation, maximum-entropy RL, and DPO as special cases.
-Master formulation.
-Given student $\pi_{\theta}$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau>0$, the anchored logits are
-$u_{i}\;=\;\frac{s_{i}-s_{i}^{\mathrm{ref}}}{\tau},\qquad s_{i}=\log\pi_{\theta}(i\mid x),\ \ s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x).$ (3)
-Note (sequence models): In autoregressive sequence models, we use length-normalized log-probabilities (token-average NLL) for $s$ and $s^{\mathrm{ref}}$ to avoid length bias.
-Under this setting, each candidate’s score is the per-token average, and these averaged scores enter the group-level softmax over candidates.
-Since length normalization amounts to an equal rescaling of all scores within a group, it preserves the softmax affine-invariance: the Fisher geometry and anchoring properties remain intact when applied to this sequence-level distribution over candidates (not the token-level distribution).
-Thus the theoretical results—groupwise shift invariance (Lemma 2.1), implicit trust region (Lemma 3.3), and closed-form projection (Proposition 3.2)—carry through in this setting.
-The anchored model distribution over a group $S_{x}$ is
-$\tilde{p}_{\theta}(i\mid S_{x})\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}\;=\;\mathrm{softmax}_{i}\Big(\tfrac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau}\Big).$ (4)
-Given a target preference distribution $q(\cdot\mid S_{x})$ (pairwise or listwise), ADPO minimizes
-$\mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\left[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\right]\;=\;\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right]+\mathrm{const}.$ (5)
-This objective performs a projection onto the anchored probability manifold; its second-order expansion around $q$ yields the softmax Fisher metric, inducing an implicit trust region centered at $\pi_{\mathrm{ref}}$ (formalized in Section 3.4).
-Relation to standard DPO.
-DPO [1] also introduces a reference model, but it enters only through log-odds differences between two candidates:
-$\Delta_{\theta,\mathrm{ref}}=\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{i}}-\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{j}}.$
-In contrast, ADPO anchors each score $s_{i}-s_{i}^{\mathrm{ref}}$ individually before normalizing across the group to form $\tilde{p}_{\theta}$.
-When $|S_{x}|=2$ and $q$ is binary (or Bradley–Terry soft targets), Eq.
-(5) reduces to the logistic preference loss used in DPO.
-For general listwise $|S_{x}|>2$, ADPO retains distribution-level anchoring and inherits a Fisher-metric trust region around $\pi_{\mathrm{ref}}$.
+  • As $\tau_{\mathrm{anc}}\to 0$: $q^{\tau_{\mathrm{anc}}}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
+  • As $\tau_{\mathrm{anc}}\to\infty$: $q^{\tau_{\mathrm{anc}}}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
+Note that, when the optimum is reachable, $\tilde{p}_{\theta}$ matches $q$ at optimality for any $\tau_{\mathrm{anc}}$; the temperature only controls how the underlying policy relates to the anchor via $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$.
+Moreover, locally around the optimum, a smaller $\tau_{\mathrm{anc}}$ induces a tighter implicit trust region: deviations in $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are penalized more strongly since $\|\nabla_{s}\mathcal{L}\|\propto 1/\tau_{\mathrm{anc}}$ and the local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$.
+3.1 Core Formulation: Anchored Distribution as the Protagonist
+ADPO performs preference learning via anchored projection: we define a distribution $\tilde{p}_{\theta}$ in coordinates centered at a reference policy $\pi_{\mathrm{ref}}$, then minimize its KL divergence to a target $q$.
+All theory revolves around this anchored distribution $\tilde{p}_{\theta}$.
+Definition of anchored distribution.
+Given student policy $\pi_{\theta}$, reference anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}>0$, we define:
+  • Anchored logits:
+    $u_{i}\;=\;\frac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau_{\mathrm{anc}}}.$ (3)
+  • Anchored distribution:
+    $\tilde{p}_{\theta}(i\mid S_{x})\;=\;\mathrm{softmax}(u)_{i}\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}.$ (4)
+  • Objective: Given target distribution $q(\cdot\mid S_{x})$,
+    $\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\Big].$ (5)
+Remark 3.1 (Sequence-level scoring).
+For autoregressive models, we use length-normalized (per-token average) log-probabilities for $s$ and $s^{\mathrm{ref}}$ to avoid length bias.
+This preserves softmax affine-invariance, so all theoretical results (shift invariance, Fisher geometry, closed-form solutions) apply to the sequence-level distribution over candidates.
+Groupwise shift invariance.
+Since $\tilde{p}_{\theta}$ depends only on log-odds differences $u_{i}-u_{j}=(s_{i}-s_{i}^{\mathrm{ref}})-(s_{j}-s_{j}^{\mathrm{ref}})$ (via softmax), adding any constant $c$ to all $s_{i}$ and $s_{i}^{\mathrm{ref}}$ within a group leaves $\tilde{p}_{\theta}$ unchanged (Lemma 2.1).
+This cancels group-level biases—e.g., annotator-specific baselines—automatically.
+Closed-form optimal solution.
+When the anchored distribution $\tilde{p}_{\theta}$ can be matched exactly to the target $q$ (e.g., sufficient expressivity, full action coverage), the optimal policy takes the multiplicative fusion form (Proposition 3.4):
+$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (6)
+Proof sketch: $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ iff $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$, i.e., $\pi_{\theta}\propto\pi_{\mathrm{ref}}q^{\tau_{\mathrm{anc}}}$.
+∎
+Fisher geometry = implicit trust region.
+The KL objective $\mathrm{KL}(q\|\tilde{p}_{\theta})$ admits a second-order expansion around the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$):
+$\mathrm{KL}(q\|\tilde{p}_{\theta})\;\approx\;\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\operatorname{Var}_{q}\!\big[s-s^{\mathrm{ref}}-\tau_{\mathrm{anc}}\log q\big],$ (7)
+where $\delta=u-u^{\star}$ is $q$-centered (Lemma 3.5).
+The matrix $\operatorname{Diag}(q)-qq^{\top}$ is the softmax Fisher metric, inducing a local quadratic trust region whose curvature depends only on $q$ (not on the anchor $\pi_{\mathrm{ref}}$).
+This provides a distribution-space analogue of TRPO’s trust region, but implicit—no explicit KL constraint needed.
+DPO as a binary special case.
+When $|S_{x}|=2$, the anchored distribution reduces to $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}})$ where $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$.
+With Bradley–Terry targets $q(i)=\sigma(\beta_{r}\Delta R)$ and uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$), Eq.
+(5) recovers the standard DPO loss [1].
+For listwise $|S_{x}|>2$, ADPO generalizes to Plackett–Luce targets while retaining the Fisher-metric trust region.
+Unification of learning paradigms.
+Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning objectives as special cases:
 Proposition 3.2 (ADPO unifies major learning paradigms).
-Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq. (5). Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover:
+Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq.
+(5).
+Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning objectives as special cases:
   1. 1.
     Supervised fine-tuning (SFT).
-If $q(\cdot\mid S_{x})$ equals the data label distribution (one-hot or soft), $\pi_{\mathrm{ref}}$ is uniform (constant logits that cancel in softmax), and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss.
-For $\tau\neq 1$, $\tilde{p}_{\theta}$ is a temperature-scaled version of $\pi_{\theta}$.
+When $q(\cdot\mid S_{x})$ is the data label distribution, $\pi_{\mathrm{ref}}$ is uniform, and $\tau_{\mathrm{anc}}=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, recovering standard cross-entropy loss.
   2. 2.
@@
   3. 3.
-    Maximum-entropy RL / offline policy projection.
-    • (Soft-optimal Boltzmann policy) The standard maximum-entropy RL objective yields the soft-optimal policy $\pi^{\star}(a|s)\propto\exp(Q(s,a)/\alpha)$ [19, 21, 20].
-To match this target via ADPO’s anchored distribution $\tilde{p}_{\theta}$, we set $q(a|s)\propto\exp(Q(s,a)/\alpha)$ and use a uniform anchor $\pi_{\mathrm{ref}}$.
-      Parametrization: ADPO constrains the anchored distribution $\tilde{p}(u)=\mathrm{softmax}((\log\pi_{\theta}-\log\pi_{\mathrm{ref}})/\tau)$ to equal $q$.
-Since there are equivalent parameterizations, we choose $\tau=1$ and $\pi_{\theta}(a|s)\propto\pi_{\mathrm{ref}}(a|s)\exp\!\big(Q(s,a)/\alpha\big)$ (with $\pi_{\mathrm{ref}}$ uniform, this reduces to $\pi_{\theta}\propto\exp(Q/\alpha)$).
-This directly recovers the standard soft-optimal policy $\pi^{\star}$ for use in sampling/evaluation, preserving the conventional Boltzmann temperature $1/\alpha$.
-Alternative parameterizations (e.g., $\tau=\alpha$ with $\pi_{\theta}\propto\exp(Q)$) yield the same $\tilde{p}$ but would shift the temperature into the anchoring scale, which we avoid for clarity.
-    • (AWR/AWAC/MPO advantage-weighted projection) When $S_{x}$ is the full action space with overlapping support between $\pi_{\text{old}}$ and the target $q$, the anchor is $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, and the target is $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$ (shift-invariant to constant baselines), with $\tau=\beta$.
-The optimal anchored distribution is $\tilde{p}_{\theta}=q\propto\pi_{\text{old}}\exp(A/\beta)$, exactly matching the advantage-weighted target.
-      Derivation of $\pi_{\theta}$ form: ADPO enforces $\tilde{p}(u)=q$ with $u=(\log\pi_{\theta}-\log\pi_{\text{old}})/\tau$ and $\tau=\beta$. Then
-      $\tilde{p}(u)\ \propto\ \exp\!\Big(\frac{\log\pi_{\theta}-\log\pi_{\text{old}}}{\beta}\Big)\ \stackrel{{\scriptstyle!}}{{\propto}}\ \pi_{\text{old}}\,\exp\!\big(A/\beta\big),$
-      which is equivalent to
-      $\frac{\pi_{\theta}}{\pi_{\text{old}}}\ \propto\ \Big(\pi_{\text{old}}\,e^{A/\beta}\Big)^{\beta}\ \Rightarrow\ \pi_{\theta}\ \propto\ \pi_{\text{old}}^{\,1+\beta}\ e^{A}.$
-      What is exactly recovered?
-ADPO matches the anchored distribution $\tilde{p}$ to the advantage-weighted target $q$; the induced policy shape $\pi_{\theta}\propto\pi_{\text{old}}^{\,1+\beta}\exp(A)$ is a consequence of the anchored projection (Prop.
-3.2) and is not the classic MPO/AWAC M-step update (which yields $\pi_{\theta}\propto\pi_{\text{old}}\exp(A/\beta)$ or directly $\pi_{\theta}=q$ under sufficient capacity).
-Our equivalence is thus at the $\tilde{p}-$level rather than the $\pi_{\theta}$-level.
-Assumptions: (1) full action space $S_{x}$ is enumerable or covered by samples; (2) advantage weights $\exp(A/\beta)$ are normalizable; (3) for off-policy data $d^{\beta}$, we approximate by ignoring or clipping importance ratios (standard in AWR/AWAC [24, 25]), yielding weighted MLE $\max_{\theta}\mathbb{E}_{(s,a)\sim d^{\beta}}[\exp(A/\beta)\log\pi_{\theta}(a|s)]$.
-MPO’s E-step weights $w\propto\exp(Q/\eta)$ with dual variable $\eta$ (enforcing a KL constraint via Lagrangian duality) follow the same projection form [26], though MPO additionally solves for $\eta$ to satisfy a KL budget.
+    Maximum-entropy RL.
+When $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and $\pi_{\mathrm{ref}}$ is uniform, setting $\tau_{\mathrm{anc}}=1$ recovers the soft-optimal Boltzmann policy $\pi_{\theta}\propto\exp(Q/\alpha)$ [19, 21].
   4. 4.
-    DPO (binary comparison form).
-For $|S_{x}|=2$ with candidates $\{i,j\}$ and Bradley–Terry soft targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$ where $\Delta R=R_{i}-R_{j}$ is the reward difference, Eq.
-(5) is equivalent to DPO up to temperature scaling and reference shift.
-Specifically, the Bayes-optimal condition $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))=q_{ij}$ yields $\Delta_{\theta}-\Delta_{\mathrm{ref}}=(\beta_{r}/\beta)\Delta R$.
-When $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform (so $\Delta_{\mathrm{ref}}=0$), this is exactly identical to the original DPO objective; otherwise, it differs by a temperature rescaling $\beta_{r}/\beta$ and an anchor-induced shift $\Delta_{\mathrm{ref}}$.
+    DPO (binary preferences).
+For $|S_{x}|=2$ with Bradley–Terry targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$, Eq.
+(5) reduces to DPO [1] when $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform.
+Remark 3.3 (Connection to AWR/AWAC/MPO).
+ADPO also connects to offline RL methods at the $\tilde{p}-$distribution level; see Appendix A for details on this partial equivalence.
 Table 2: ADPO as a unifying projection framework.
-Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover major learning paradigms.
-Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau)$ to the target $q$; the corresponding student policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ when the optimum is reachable (Prop.
-3.2).
-Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau$ Recovers ($\tilde{p}_{\theta}$ / $\pi_{\theta}$)
-SFT data (one-hot/soft) uniform 11 cross-entropy / same
-KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}\!\big(q_{T},\mathrm{softmax}(s/T)\big)$ / train at $T$, infer T=1T{=}1
-KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}((s-s_{T})/\tau)\big)$ / trust region
-[…]
-[…]
-DPO (binary) Bradley–Terry / hard uniform any original DPO (pairwise)
-ADPO (listwise) Plackett–Luce / soft any (teacher/old) any this work
+Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning paradigms.
+Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop.
+3.4).
+Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau_{\mathrm{anc}}$ Recovers
+SFT data (one-hot/soft) uniform 11 cross-entropy
+KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(s/T))$
+KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any teacher-anchored KD
+Max-ent RL $\propto\exp(Q/\alpha)$ uniform 11 $\pi_{\theta}\propto\exp(Q/\alpha)$
+DPO (binary) Bradley–Terry / hard uniform any original DPO
+ADPO (listwise) Plackett–Luce / soft any any this work
 Proposition 3.4 (Closed-form optimal solution).
@@
 (Discrete case) If the policy $\pi_{\theta}$ has sufficient expressivity (i.e., the parametrization can represent arbitrary log-odds over $S_{x}$), then the distribution-level optimum of $\operatorname*{arg\,min}_{\theta}\mathrm{KL}(q\|\tilde{p}_{\theta})$ satisfies
-$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau}.$ (6)
+$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (8)
 (Continuous case) For continuous action spaces, assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure (e.g., Lebesgue measure).
 Define the anchored distribution via the density ratio:
-$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau}.$
+$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau_{\mathrm{anc}}}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau_{\mathrm{anc}}}.$
 Setting $\tilde{p}_{\theta}=q$ yields the closed-form solution
-$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau},$
-which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau}\,da<\infty$ to ensure normalizability.
+$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}},$
+which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}}\,da<\infty$ to ensure normalizability.
 Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
-Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau}$.
+Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau_{\mathrm{anc}}}$.
 The constant $c$ is absorbed into the proportionality constant.
@@
 When $S_{x}$ is only a subset (pairwise/listwise ranking), (8) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints.
-Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau$ within each group (as shown in Theorem 3.6), but the optimization over shared parameters $\theta$ is typically non-convex.
-The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ for all groups (except in the tabular case or with unlimited expressivity).
+Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ within each group (as shown in Theorem 3.8), but the optimization over shared parameters $\theta$ is typically non-convex.
+The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot\mid S_{x})=q(\cdot\mid S_{x})$ for all groups (except in the tabular case or with unlimited expressivity).
 Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
@@
 Table 2 summarizes the unification.
+Mirror descent = KL proximal step.
+If one directly optimizes over the anchored logits $u$ (e.g., tabular setting), each gradient step is equivalent to a KL-trust-region proximal problem (Theorem 3.7):
+$u_{t+1}=\arg\min_{u}\ \langle\nabla\mathcal{L}(u_{t}),\,u\rangle+\tfrac{1}{\eta}\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u)),$
+whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$. This exhibits ADPO as a Bregman-proximal method in distribution space.
+Dynamic anchoring: monotonicity within fixed-anchor phases.
+When the anchor $\pi_{\mathrm{ref}}$ is held fixed, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is strictly convex on the subspace $\mathbf{1}^{\perp}$.
+Under either (i) local analysis near $u^{\star}$ or (ii) bounded-domain constraints $\tilde{p}_{\theta}\in[\varepsilon,1-\varepsilon]^{K}$, $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$, guaranteeing linear convergence $\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star}))$ within each fixed-anchor phase (Theorem 3.8).
+When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes, and we do not claim monotonicity across anchor updates—only within each phase.
 Standing assumptions.
@@
   1. (i)
-    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i|x)>0\Rightarrow\pi_{\mathrm{ref}}(i|x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ are well-defined.
+    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i\mid x)>0\Rightarrow\pi_{\mathrm{ref}}(i\mid x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ are well-defined.
   2. (ii)
@@
   5. (v)
-    Expressivity: Closed-form optimal solutions (Proposition 3.2) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau}$.
+    Expressivity: Closed-form optimal solutions (Proposition 3.4) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$.
 In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
@@
 $\ell_{ij}^{\text{Soft-DPO}}=\log\left(1+e^{\beta(\Delta_{\theta}-\Delta_{\text{ref}})}\right)-q_{ij}\,\beta(\Delta_{\theta}-\Delta_{\text{ref}}),$ (9)
-where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(\tau_{i}|x)$.
+where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(y_{i}|x)$.
 Key properties:
@@
   • Reference anchoring: Relative updates $\Delta_{\theta}-\Delta_{\text{ref}}$ are invariant to groupwise shifts.
+Temperature alignment: The pairwise formulation uses inverse temperature $\beta$, which relates to the listwise anchoring temperature via $\beta=1/\tau_{\mathrm{anc}}$.
+Specifically, $\tilde{p}_{\theta}(i)=\sigma\!\big((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)$.
 3.3 Listwise Instantiation
-Notation for this subsection: $\tau$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
-For group $S_{x}=\{\tau_{1},\ldots,\tau_{P}\}$, the ADPO listwise loss is:
-$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right],$ (8)
+Notation for this subsection: $\tau_{\mathrm{anc}}$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
+For group $S_{x}=\{y_{1},\ldots,y_{P}\}$, the ADPO listwise loss is:
+$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})\right],$ (10)
 where
-$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau\right)},$ (9)
-and teacher target $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
+$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau_{\mathrm{anc}}\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau_{\mathrm{anc}}\right)},$ (11)
+and teacher target $q(i\mid S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
 Equivalent expansion. The anchored listwise cross-entropy (10) can be equivalently written as:
-$\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})=-\frac{1}{\tau}\sum_{i\in S_{x}}q(i|S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau}\right).$ (10)
-This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau\rangle$ and a log-sum-exp normalization.
+$\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})=-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i\in S_{x}}q(i\mid S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\right).$ (12)
+This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\rangle$ and a log-sum-exp normalization.
 It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below.
 Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (12) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
-3.4 Fisher Geometry and Implicit Trust Region
-We now formalize how anchoring provides an implicit trust region through Riemannian geometry.
-Lemma 3.3 (Implicit trust region; quadratic form at $p=q$).
-Let $u=(s-s^{\mathrm{ref}})/\tau$ and $\tilde{p}(u)=\mathrm{softmax}(u)$.
-Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$ (the softmax is shift-invariant).
-Define $\delta=u-u^{\star}$ and center it by $q$ as $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$.
+3.4 Formal Statements of Core Results
+We now provide formal statements and proofs for the results sketched in Section 3.
+Lemma 3.5 (Implicit trust region).
+Let $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ and $\tilde{p}(u)=\mathrm{softmax}(u)$.
+Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$.
+Define $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$, the $q$-centered deviation.
 Then, as $u\to u^{\star}$,
-$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\tau\log q_{i}\big]\;+\;o(\cdot),$
-where the variance is taken with respect to $q$, measuring the deviation of the anchored log-odds from the target $q$.
-The quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which has a null-space along $\mathbf{1}$ (reflecting shift-invariance).
+$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$
+The curvature is given by the softmax Fisher metric $\operatorname{Diag}(q)-qq^{\top}$, which is independent of the anchor choice and has a null-space along $\mathbf{1}$.
 Proof sketch.
 Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$.
-Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$.
+Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$.
 A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form.
@@
 ∎
-Corollary 3.4 (Local bounds).
-Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\mathrm{Diag}(q)-qq^{\top}$ (the null-space is along $\mathbf{1}$). Then locally around $u^{\star}$,
+Corollary 3.6 (Spectral bounds).
+Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\operatorname{Diag}(q)-qq^{\top}$.
+In the quotient space $\mathbb{R}^{K}/\mathrm{span}\{\mathbf{1}\}$, the objective is strongly convex.
+Locally around $u^{\star}$,
 $\mathrm{KL}(q\|\tilde{p}(u))\;\geq\;\frac{\mu(q)}{2}\,\|\bar{\delta}\|^{2},$
-where $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$ is the $q$-centered deviation.
-By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$, providing a distributional deviation bound.
-On the smallest nonzero eigenvalue: For $K=2$, $\mu(q)=2q(1-q)$.
-For $K>2$, $\mu(q)$ depends on the full spectrum of $q$; a uniform, distribution-independent lower bound is generally loose.
-When $\min_{i}q_{i}\geq\varepsilon>0$, the smallest nonzero eigenvalue satisfies $\mu(q)\geq\varepsilon$.
-Proof sketch: For any $v\in\mathbf{1}^{\perp}$, the Rayleigh quotient gives $v^{\top}(\mathrm{Diag}(q)-qq^{\top})v=\sum_{i}q_{i}v_{i}^{2}-(\sum_{i}q_{i}v_{i})^{2}=\mathrm{Var}_{q}[v]\geq\varepsilon\|v\|^{2}$ since $\sum_{i}v_{i}=0$ and the variance of $v$ under any distribution bounded below by $\varepsilon$ is at least $\varepsilon\|v\|^{2}$.
+where $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$. By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$.
+For $K=2$: $\mu(q)=2q(1-q)$.
+For $K>2$, when $\min_{i}q_{i}\geq\varepsilon>0$, there exists a constant $c(\varepsilon,K)>0$ such that $\mu(q)\geq c(\varepsilon,K)$ on the quotient space.
+In particular, for any $v$ satisfying $\langle q,v\rangle=0$,
+$\operatorname{Var}_{q}[v]=\sum_{i}q_{i}v_{i}^{2}\geq\varepsilon\|v\|^{2}.$
 Geometric interpretation.
-ADPO performs optimization in the anchored $u$-coordinate system: the reference $\pi_{\mathrm{ref}}$ defines the coordinate origin via $u=(s-s^{\mathrm{ref}})/\tau$, but the center of the local quadratic bowl is at $u^{\star}$ satisfying $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$, not at the anchor itself.
-The local curvature at $u^{\star}$ is governed by the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which depends only on the target $q$.
-Thus: the anchor determines the coordinate system, while the target $q$ determines both the minimizer $u^{\star}$ and the local geometry.
-This contrasts with:
-  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex.
-While the Fisher curvature exists at any point, there is no fixed reference anchor, so the geometry lacks a stable coordinate origin.
-  • TRPO: Constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in parameter space, defining a moving local quadratic region that requires explicit constraint optimization.
-  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed coordinate origin, inducing an implicit trust region without explicit constraints.
+The anchor $\pi_{\mathrm{ref}}$ defines the coordinate origin, while the target $q$ determines the minimizer $u^{\star}$ and local curvature.
+This provides a distribution-space analogue of TRPO but with an implicit (Fisher-metric) trust region—no explicit KL constraint needed.
 Figure 1: ADPO’s anchored trust region.
-The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself.
+The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself.
 The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift.
 This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
-3.5 Geometric Interpretation: KL-Mirror Descent in $u$-Space
-Recall $A(u)=\log\!\sum_{j}\exp(u_{j})$ and $\tilde{p}(u)=\nabla A(u)=\mathrm{softmax}(u)$, and our anchored objective $\mathcal{L}(u)=A(u)-\langle q,u\rangle=\mathrm{KL}\!\big(q\;\|\;\tilde{p}(u)\big)+H(q)$.
-Let $u_{t}$ be the current anchored logits and $g_{t}=\nabla\mathcal{L}(u_{t})=\tilde{p}(u_{t})-q$.
-Theorem 3.5 (KL-mirror descent / Bregman-prox step in $u$-space).
-If one were to directly optimize $\mathcal{L}$ over the anchored logits $u$ (rather than shared parameters $\theta$), then for any step size $\eta>0$, the one-step update can be written equivalently as the solution of the following proximal subproblem:
-$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle g_{t},\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big).$ (11)
-Equivalently, since $D_{A}(u\|u_{t})=A(u)-A(u_{t})-\langle\nabla A(u_{t}),u-u_{t}\rangle=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$, (11) is a standard mirror descent / Bregman-prox step with generator $A$.
+3.5 KL-Mirror Descent Interpretation
+Theorem 3.7 (KL-mirror descent / Bregman-prox step).
+If one directly optimizes $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ over anchored logits $u$ (where $A(u)=\log\sum_{j}e^{u_{j}}$), then for any step size $\eta>0$, the one-step update is equivalent to:
+$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle\nabla\mathcal{L}(u_{t}),\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big),$ (13)
+whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$.
 Proof.
-Linearize $\mathcal{L}$ at $u_{t}$: $\mathcal{L}(u)\approx\mathcal{L}(u_{t})+\langle g_{t},u-u_{t}\rangle$.
 A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$.
-Using $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (11).
+Using the Bregman divergence relation $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (13).
 ∎∎
-Dual form in $p$-space. Via the Legendre map $p=\nabla A(u)$, the step (11) is equivalent to the entropic proximal in probability space
-$\tilde{p}_{t+1}=\arg\min_{p\in\Delta}\ \langle g_{t},p\rangle+\tfrac{1}{\eta}\mathrm{KL}\!\big(p\,\|\,\tilde{p}_{t}\big),$
-whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta g_{t})$.
-Multiplicative-weights view.
-The solution of (11) is equivalent to the multiplicative-weights update in the anchored distribution:
-$\tilde{p}_{t+1}\ \propto\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,g_{t}\big)\ =\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,(\tilde{p}_{t}-q)\big),$ (12)
-followed by renormalization. This exhibits ADPO as a KL-trust-region step in distribution space.
-Step-size $\eta$ vs temperature $\tau$ (heuristic).
-Anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$ rescale curvature by $1/\tau^{2}$ (Lemma 3.3).
-To maintain comparable “trust-region tightness” in the unanchored score space $s$, one may heuristically choose
-[…]
-This is a practical guideline; a rigorous derivation would require accounting for the chain rule $\partial u/\partial\theta$ when optimizing over shared parameters $\theta$.
 3.6 Dynamic Anchor Stability
-Crucially, Lemma 3.3 holds regardless of anchor choice.
-The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region.
-While the coordinate $u=(s-s^{\mathrm{ref}})/\tau$ depends on the anchor, the local quadratic form at the optimum only depends on $q$; thus the curvature (Fisher metric) is anchor-agnostic, albeit the optimization objective itself changes with the anchor.
-Common anchor choices include:
-  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable coordinate origin for consistent target tracking;
-  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive coordinate system similar to TRPO’s old-policy mechanism;
-  • Teacher anchor: A stronger teacher $\pi_{\text{teacher}}$ for offline distillation (typically fixed).
-For dynamic anchors, we characterize the monotonic improvement within each fixed-anchor phase:
-Theorem 3.6 (Fixed-anchor-phase monotonic descent in $u$-space).
-Consider a training phase where the anchor $\pi_{\mathrm{ref}}$ is held fixed.
-Define $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ and $u=(s-s^{\mathrm{ref}})/\tau$ are the anchored logits.
-When directly optimizing $\mathcal{L}$ over $u$ (e.g., tabular setting or each group having independent degrees of freedom), $\mathcal{L}$ is strictly convex on the subspace $\mathbf{1}^{\perp}$ (orthogonal to the all-ones vector) with Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite on $\mathbf{1}^{\perp}$ with one zero eigenvalue along $\mathbf{1}$.
-Strong convexity (local or bounded-domain):
-  • Local: In a neighborhood of $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ approaches $\mathrm{Diag}(q)-qq^{\top}$ up to $o(1)$, so there exists $\mu>0$ (depending on $q$) such that $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$ in this neighborhood.
-  • Bounded domain: If the optimization is constrained to a domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$ and some $\varepsilon>0$ (i.e., the current anchored distribution $\tilde{p}_{\theta}(u)$ is bounded away from the probability simplex boundary), then $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$).
-Under either condition, standard gradient descent with appropriate step size $\eta\in(0,1/L]$ (where $L$ is a local smoothness constant) yields linear (geometric) convergence:
-$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\,(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})),$
-implying monotonic decrease in $\mathrm{KL}(q\|\tilde{p}_{\theta})$ within each fixed-anchor phase.
-Note: When the anchor is updated (e.g., $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the objective $\tilde{p}_{\theta}$ is re-normalized with the new anchor, so the loss function changes.
-We do not claim monotonicity across anchor updates; instead, each fixed-anchor phase provides stable descent, and the anchor update adapts the coordinate system for the next phase.
-Caveat for parametrized models: In practice, the anchored logits $u$ are derived from shared parameters $\theta$ (e.g., neural networks), and the optimization problem over $\theta$ is typically non-convex.
-The above results provide geometric intuition for the loss landscape in the anchored distribution space but do not constitute convergence guarantees for parameterized training.
+The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ is anchor-agnostic (depends only on $q$), allowing flexible anchor strategies: (i) fixed $\pi_{\text{ref}}$ (frozen, stable origin); (ii) dynamic $\pi_{\theta_{\text{old}}}$ (periodic updates, adaptive like TRPO); (iii) teacher $\pi_{\text{teacher}}$ (offline distillation).
+We characterize monotonic improvement within each fixed-anchor phase:
+Theorem 3.8 (Fixed-anchor-phase monotonic descent).
+When the anchor is fixed, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex on the quotient space and $L$-smooth on any bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$.
+Consequently, gradient descent with step size $\eta\in(0,1/L]$ (or KL-mirror descent with the corresponding step-size condition) enjoys the linear rate
+$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big)$
+within that fixed-anchor phase. When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes—we do not claim monotonicity across anchor updates.
 Proof sketch.
-The log-partition function $A(u)$ is strictly convex (its Hessian is the softmax Fisher metric $\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite with null-space along $\mathbf{1}$).
-The Hessian depends on the current $\tilde{p}_{\theta}(u)$, not directly on $q$.
-However, near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, it coincides with $\mathrm{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity.
-Alternatively, if $\tilde{p}_{\theta}(u)$ is constrained to $\varepsilon,1-\varepsilon]^{K}$, the minimum eigenvalue is uniformly bounded below.
-Standard strongly convex optimization theory then guarantees linear convergence.
-When the anchor changes, $u$ is redefined, so objectives from different phases are not directly comparable.
+$A(u)$ is strictly convex with Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ (softmax Fisher metric).
+Near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, this coincides with $\operatorname{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity.
+In bounded domains where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$, the Fisher metric satisfies $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$ in Corollary 3.6).
+The spectral upper bound of $\nabla^{2}A(u)$ is also bounded in this domain, yielding $L$-smoothness with $L=L(\varepsilon,K)$.
 ∎∎
 Implication.
-This theorem explains why dynamic anchors (Section 5.4) maintain stability during online exploration: within each fixed-anchor phase, the implicit Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts to track policy evolution, similar to TRPO’s mechanism.
-Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting coordinates to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining a stable coordinate origin throughout).
-Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
+Within each fixed-anchor phase, Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts (like TRPO).
+Experiments (Section 5) reveal task-dependent trade-offs: dynamic anchors excel at online exploration; fixed anchors excel at offline distillation.
 Gradients.
 Pairwise: $\frac{\partial\ell_{ij}}{\partial\theta}=\beta\big(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij}\big)\big(\nabla_{\theta}s_{i}-\nabla_{\theta}s_{j}\big).$
-Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$
+Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i\mid S_{x})-q(i\mid S_{x})\big).$
 4 Experimental Setup
@@
 ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
-  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains updates.
+  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates.
 This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
@@
 While we have validated our framework on controlled bandits and continuous control, applying this unified anchoring perspective to large-language model alignment, where different training stages (SFT, RLHF, continual learning) may benefit from different anchor strategies, is a critical and promising direction for future work.
-Appendix A Fisher Metric Connection and Relation to KD/TRPO
+Appendix A Connection to Offline RL Methods (AWR/AWAC/MPO)
+ADPO connects to advantage-weighted reinforcement learning methods at the level of anchored distributions, though not at the policy level. We detail this partial equivalence here.
+A.1 Advantage-Weighted Projection in ADPO
+When $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (old policy anchor), $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$, and $\tau=\beta$, ADPO’s anchored distribution matches:
+$\tilde{p}_{\theta}(a|s)=\mathrm{softmax}\!\left(\frac{\log\pi_{\theta}(a|s)-\log\pi_{\text{old}}(a|s)}{\beta}\right)\;\propto\;\pi_{\text{old}}(a|s)\exp\!\big(A/\beta\big).$
+Derivation of the induced policy: By Proposition 3.4, the optimal $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$. Substituting:
+$\pi_{\theta}(a|s)$ $\propto\pi_{\text{old}}(a|s)\cdot\big[\pi_{\text{old}}(a|s)\exp(A/\beta)\big]^{\beta}$
+$=\pi_{\text{old}}(a|s)\cdot\pi_{\text{old}}(a|s)^{\beta}\cdot\exp(A)$
+$=\pi_{\text{old}}(a|s)^{1+\beta}\cdot\exp\!\big(A^{\pi_{\text{old}}}(s,a)\big).$
+A.2 Comparison with Classic AWR/AWAC/MPO
+Classic AWR/AWAC M-step [24, 25]: Given advantage-weighted samples, the policy update solves
+$\max_{\theta}\mathbb{E}_{(s,a)\sim d}\!\big[\exp(A(s,a)/\beta)\log\pi_{\theta}(a|s)\big],$
+which (ignoring normalization) yields $\pi_{\theta}(a|s)\propto\exp(A(s,a)/\beta)$ when capacity is sufficient.
+MPO E/M-step [26]: The E-step forms $q(a|s)\propto\exp(Q(s,a)/\eta)$ with dual variable $\eta$ (enforcing a KL budget via Lagrangian).
+The M-step minimizes $\mathrm{KL}(q\|\pi_{\theta})$, yielding $\pi_{\theta}=q$ under sufficient capacity.
+A.3 Key Distinction
+ADPO’s equivalence is at the $\tilde{p}-$level (anchored distribution):
+$\tilde{p}_{\theta}\;\propto\;\pi_{\text{old}}\exp(A/\beta)\qquad\text{(matches AWR target)},$
+but the induced policy $\pi_{\theta}$ differs:
+$\text{ADPO:}\quad\pi_{\theta}\;\propto\;\pi_{\text{old}}^{1+\beta}\exp(A),\qquad\text{AWR/MPO:}\quad\pi_{\theta}\;\propto\;\pi_{\text{old}}\exp(A/\beta)\text{ or }\pi_{\theta}=q.$
+Interpretation: ADPO’s anchored projection induces a multiplicative fusion $\pi_{\mathrm{ref}}\cdot q^{\tau}$, which preserves the geometric structure of anchoring but differs from classic policy-space updates.
+The power $1+\beta$ arises from the composition: anchoring at $\pi_{\text{old}}$ followed by temperature-scaled projection toward $q$.
+When does this matter?
+In tabular settings or with full expressivity over discrete actions, both formulations can represent the same optimal behavior (via different parametrizations).
+In function approximation with shared parameters, the difference becomes significant: ADPO’s form provides an implicit trust region via Fisher geometry, while AWR relies on explicit advantage weighting.
+Appendix B Fisher Metric Connection and Relation to KD/TRPO
 Fisher metric expansion and Bregman divergence perspective.
@@
 Expanding around $u^{\star}$, the second-order Taylor expansion yields
-$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (15)
-The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space.
+$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (15)
+The matrix $\operatorname{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space.
 Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry, with connections to mirror descent and Bregman proximal methods.
@@
 ADPO Distribution (anchored) Reference policy Softmax Fisher Implicit (anchored KL)
-Appendix B Proofs and Derivations
-B.1 Proof of Proposition 3.1
+Appendix C Proofs and Derivations
+C.1 Proof of Proposition 3.2
 We prove that ADPO unifies major learning paradigms through different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$.
@@
 which performs optimization in teacher-anchored coordinates with an implicit Fisher-metric trust region (see Lemma 3.5).
-(iii) Maximum-entropy RL / offline policy projection.
+(iii) Maximum-entropy RL.
 For $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and uniform $\pi_{\mathrm{ref}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
 $\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}(s_{a}/\tau)\propto\exp(Q(s,a)/\alpha),$
-so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=\alpha$ gives the soft-optimal Boltzmann form $\pi\propto\exp(Q/\alpha)$.
-For advantage-weighted updates with $q(a\mid s)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta)$ and $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
-$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}\!\left(\frac{s_{a}-s_{a}^{\mathrm{old}}}{\tau}\right)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta),$
-so $(s_{a}-s_{a}^{\mathrm{old}})/\tau\approx A^{\pi_{\mathrm{old}}}(s,a)/\beta$, giving $\pi_{\theta}(a\mid s)\propto\pi_{\mathrm{old}}(a\mid s)\exp((\tau/\beta)A^{\pi_{\mathrm{old}}}(s,a))$.
-Choosing $\tau=\beta$ recovers AWR/AWAC/MPO projections.
+so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=1$ gives $s_{a}\approx Q(s,a)/\alpha$, hence $\pi_{\theta}\propto\exp(Q/\alpha)$, the soft-optimal Boltzmann form.
 (iv) DPO (binary comparison form).
@@
 which matches the original DPO objective with inverse temperature $1/\tau$. ∎
-B.2 Proof of Lemma 3.3
+C.2 Proof of Lemma 3.5
 Setup.
@@
 Then
-$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\mathrm{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
+$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\operatorname{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
 Back to anchored scores.
 Since $u=(s-s^{\mathrm{ref}})/\tau$, we obtain
-$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
+$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\operatorname{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
 i.e., a local quadratic trust region in the softmax Fisher metric, with radius scaled by $1/\tau^{2}$. ∎
-B.3 Proof of Lemma 2.1
+C.3 Proof of Lemma 2.1
 Anchored student invariance.
@@
 ∎
-Appendix C Implementation and Training Details
-C.1 Noise Generation Procedures
+Appendix D Implementation and Training Details
+D.1 Noise Generation Procedures
 We provide complete mathematical specifications for the four noise families used in our experiments.
@@
 with $\sigma_{\text{out}}^{2}=100\sigma^{2}$. The severity levels correspond to:
-  • Light: SNR = $\frac{\mathrm{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
+  • Light: SNR = $\frac{\operatorname{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
   • Medium: SNR = 0.5, $p_{\text{out}}=0.10$
@@
   • Heavy: $\gamma=1.0$
-C.2 Hyperparameters and Training Configuration
+D.2 Hyperparameters and Training Configuration
 Table 9: Complete hyperparameter configuration for all experiments.
@@
 Teacher sampling ratio 75% (teacher) + 25% (random)
-C.3 Metrics Definition
+D.3 Metrics Definition
 WinMass.
@@
 where $i^{\star}=\operatorname*{arg\,max}_{i\in S_{x}}R_{i}^{\star}$ is the item with the highest ground-truth reward. For random selection with $P=4$ candidates, the baseline WinMass is $1/P=0.25$.
-Appendix D Additional Experiments and Ablations
-D.1 Statistical Significance Tests
+Appendix E Additional Experiments and Ablations
+E.1 Statistical Significance Tests
 Method.
@@
 We report $p$-values and reject $H_{0}$ at $\alpha=0.05$.
-D.2 Effect of Reference Initialization
+E.2 Effect of Reference Initialization
 Ablation results.
@@
 Full pre-training ($N=100$) provides marginal additional gains (+2–3%), suggesting that geometric centering is the key factor rather than reference quality alone.
-D.3 Temperature Sensitivity (Full Grid)
+E.3 Temperature Sensitivity (Full Grid)
 Figure 5 in the main text shows three representative methods. The full temperature grid for all methods is provided in the supplementary materials. Key findings:
@@
   • Listwise-KDE benefits from lower $\tau$ (0.5–1.0) but degrades at $\tau>2.0$.
-D.4 KDE-CDF-Logit Transform Details
+E.4 KDE-CDF-Logit Transform Details
 Given rewards $\{\tilde{R}_{i}\}_{i=1}^{P}$ in a group, ADPO-KDE constructs targets via:
@@
 This rank-like transform bounds outlier influence through the CDF and preserves separability via the logit.
-Appendix E Reproducibility Checklist
+Appendix F Reproducibility Checklist
 To facilitate reproduction of our results, we provide the following information:
@@
   • Ground-truth rewards $R_{i}^{\star}=f_{\star}(\text{concat}(c,v_{i}))$ where $f_{\star}$ is a 2-layer MLP with hidden dimension 64;
-  • Noisy observations $\tilde{R}_{i}$ generated according to Section C.
+  • Noisy observations $\tilde{R}_{i}$ generated according to Section D.
 Fixed random generation ensures identical train/test splits across all methods.

### v6

@@
 However, DPO’s reliance on hard, pairwise preference labels makes it brittle.
-In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our method (0.472 vs.
+In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy noise, DPO’s performance degrades by up to 93% relative to our method (0.375 vs.
 0.725 WinMass).
@@
 We introduce Anchored Direct Preference Optimization (ADPO), a unified framework that connects seemingly disparate learning paradigms.
-Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3, Proposition 3.2).
+Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3.2, Proposition 3.4).
 By centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
@@
 […]
-Collectively, our work makes four key contributions: (1) we propose ADPO as a unified projection framework that subsumes major learning paradigms (SFT, KD, max-ent RL, DPO) and robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
+[…]
 2 Related Work
 Preference optimization.
-More generally, the Preference Optimization (PO) framework [18] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization.
+More generally, the Preference Optimization (PO) framework [18] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3.2), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization.
 DPO [1] reparameterizes the reward model through policy-reference log-ratios, enabling direct optimization.
@@
 Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in coordinates centered at the reference, while the local quadratic well (Fisher metric) is centered at the target distribution $q$, yielding an implicit trust region without an explicit KL constraint.
-We are not aware of prior work that formulates an anchored cross-entropy maintaining a fixed geometric center in probability space.
+To our knowledge, a direct anchored cross-entropy formulation emphasizing a fixed geometric center in probability space has not been formalized in prior work, though related ideas appear in advantage-weighted methods [24, 25] and maximum a posteriori policy optimization [26].
 Entropy-regularized grounding and its link to ADPO.
@@
 $\max_{\pi}\ \mathbb{E}_{s\sim\mathcal{D}}\Big[\mathbb{E}_{a\sim\pi(\cdot|s)}r(s,a)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|s)\big)\Big],$
-the optimal policy admits the Boltzmann form [19, 21]
+the optimal policy admits the Boltzmann form [19] (also known as energy-based policies [21])
 $\pi^{*}(a|s)\;=\;\frac{1}{Z(s)}\exp\!\Big(\tfrac{1}{\alpha}\,r(s,a)\Big),$ (1)
@@
 Eqs.
-(1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets.
+(1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets.
 A full derivation of (1) as the minimizer of $\mathbb{E}_{s}\,\mathrm{KL}\big(\pi(\cdot|s)\,\|\,\pi^{*}(\cdot|s)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs.
@@
 Lemma 2.1 (Groupwise shift invariance).
-For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance.
+For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(\ell-\ell^{\mathrm{ref}})$—inherits groupwise shift invariance.
+Proof sketch.
+For any group-constant $c$, $\mathrm{softmax}(u)=\mathrm{softmax}(u+c\mathbf{1})$ by shift-invariance.
+Since $r^{\prime}(s,a)=r(s,a)-h(s)$ only changes the group constant (log-partition $\log Z(s)$ in Eq.
+(2)), the Bradley–Terry/Plackett–Luce probabilities and anchored distribution $\tilde{p}$ remain unchanged.
 See [18], Prop.
 3.1 and App.
-D.2.
-3 Anchored Direct Preference Optimization
+D.2 for full derivation.
+∎
+3 Anchored Direct Preference Optimization (ADPO)
+3.1 Problem Setting and Anchored Softmax
+Motivation: Why anchoring?
+Traditional preference optimization methods (e.g., DPO with uniform reference) operate on absolute log-probabilities $\ell_{i}=\log\pi_{\theta}(y_{i}|x)$, making them sensitive to:
+  • Annotator baseline biases: Different annotators may have systematically different baseline preferences
+  • Sequence-level biases: Length bias in language models, state-value offsets in RL
+  • Distribution shift: Changes in the reference distribution between training and deployment
+ADPO addresses these issues by introducing a reference anchor $\pi_{\mathrm{ref}}$ and working with relative log-probabilities $(\ell-\ell^{\mathrm{ref}})$, which automatically cancels group-level baseline shifts through the softmax’s shift-invariance property (detailed below).
 Notation.
-$x$: context (prompt/state); $S_{x}=\{y_{1},\ldots,y_{|S_{x}|}\}$: candidate set (actions/outputs).
-$\pi_{\theta}(y\mid x)$: student policy; $\pi_{\mathrm{ref}}(y\mid x)$: anchor (fixed or dynamic).
-$s_{y}=\log\pi_{\theta}(y\mid x)$, $s^{\mathrm{ref}}_{y}=\log\pi_{\mathrm{ref}}(y\mid x)$: log-probabilities.
-$q(y\mid S_{x})$: target preference distribution (from teacher, rewards, or labels).
-$u_{y}=(s_{y}-s^{\mathrm{ref}}_{y})/\tau_{\mathrm{anc}}$: anchored logits.
-$\tilde{p}_{\theta}(\cdot\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution.
-Table 1: Temperature parameters and their roles across ADPO formulations.
-Symbol Role Where used
-$\beta_{r}$ Teacher/target inverse temperature Forms $q$ from rewards: $q\propto\exp(R/\beta_{r})$
-$\beta$ Student inverse temperature (pairwise) Pairwise logistic: $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))$
-$\tau_{\mathrm{anc}}$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ (listwise)
-Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them.
-We set $\beta=\beta_{r}=1.0$ by default.
-However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau_{\mathrm{anc}}$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau_{\mathrm{anc}}}(\tilde{p}-q)$).
-Thus smaller $\beta$ slows pairwise updates; smaller $\tau_{\mathrm{anc}}$ tightens the trust region via curvature $\propto 1/\tau_{\mathrm{anc}}^{2}$ but increases gradient magnitudes.
-Remark (temperature $\tau_{\mathrm{anc}}$ and regularization strength): In listwise settings with anchoring, $\tau_{\mathrm{anc}}$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ (Proposition 3.4).
-Intuitively:
-  • As $\tau_{\mathrm{anc}}\to 0$: $q^{\tau_{\mathrm{anc}}}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
-  • As $\tau_{\mathrm{anc}}\to\infty$: $q^{\tau_{\mathrm{anc}}}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
-Note that, when the optimum is reachable, $\tilde{p}_{\theta}$ matches $q$ at optimality for any $\tau_{\mathrm{anc}}$; the temperature only controls how the underlying policy relates to the anchor via $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$.
-Moreover, locally around the optimum, a smaller $\tau_{\mathrm{anc}}$ induces a tighter implicit trust region: deviations in $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are penalized more strongly since $\|\nabla_{s}\mathcal{L}\|\propto 1/\tau_{\mathrm{anc}}$ and the local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$.
-3.1 Core Formulation: Anchored Distribution as the Protagonist
-ADPO performs preference learning via anchored projection: we define a distribution $\tilde{p}_{\theta}$ in coordinates centered at a reference policy $\pi_{\mathrm{ref}}$, then minimize its KL divergence to a target $q$.
-All theory revolves around this anchored distribution $\tilde{p}_{\theta}$.
-Definition of anchored distribution.
-Given student policy $\pi_{\theta}$, reference anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}>0$, we define:
-  • Anchored logits:
-    $u_{i}\;=\;\frac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau_{\mathrm{anc}}}.$ (3)
-  • Anchored distribution:
-    $\tilde{p}_{\theta}(i\mid S_{x})\;=\;\mathrm{softmax}(u)_{i}\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}.$ (4)
-  • Objective: Given target distribution $q(\cdot\mid S_{x})$,
-    $\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\Big].$ (5)
-Remark 3.1 (Sequence-level scoring).
-For autoregressive models, we use length-normalized (per-token average) log-probabilities for $s$ and $s^{\mathrm{ref}}$ to avoid length bias.
-This preserves softmax affine-invariance, so all theoretical results (shift invariance, Fisher geometry, closed-form solutions) apply to the sequence-level distribution over candidates.
+Let $x\!\sim\!\mathcal{D}$ be a context (e.g., prompt or state). For a candidate set $S_{x}=\{y_{1},\dots,y_{P}\}$ (e.g., responses or actions), we define:
+  • Student policy $\pi_{\theta}(y|x)$ with log-probabilities $\ell_{i}=\log\pi_{\theta}(y_{i}|x)$
+  • Anchor policy $\pi_{\mathrm{ref}}(y|x)$ with log-probabilities $\ell_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(y_{i}|x)$
+  • Target preference $q(\cdot|S_{x})$ with $q\succ 0$ (strictly positive distribution)
+We work with anchored logits:
+$u_{i}\;=\;\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\,,\qquad\bar{u}_{i}\;=\;u_{i}-\sum_{j\in S_{x}}q(j|S_{x})\,u_{j},$ (3)
+where $\tau_{\mathrm{anc}}>0$ is an anchoring temperature controlling regularization strength, and $\bar{u}$ denotes the $q$-centered anchored logits (always applied for numerical stability and geometric consistency).
 Groupwise shift invariance.
-Since $\tilde{p}_{\theta}$ depends only on log-odds differences $u_{i}-u_{j}=(s_{i}-s_{i}^{\mathrm{ref}})-(s_{j}-s_{j}^{\mathrm{ref}})$ (via softmax), adding any constant $c$ to all $s_{i}$ and $s_{i}^{\mathrm{ref}}$ within a group leaves $\tilde{p}_{\theta}$ unchanged (Lemma 2.1).
-This cancels group-level biases—e.g., annotator-specific baselines—automatically.
-Closed-form optimal solution.
-When the anchored distribution $\tilde{p}_{\theta}$ can be matched exactly to the target $q$ (e.g., sufficient expressivity, full action coverage), the optimal policy takes the multiplicative fusion form (Proposition 3.4):
-$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (6)
-Proof sketch: $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ iff $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$, i.e., $\pi_{\theta}\propto\pi_{\mathrm{ref}}q^{\tau_{\mathrm{anc}}}$.
-∎
-Fisher geometry = implicit trust region.
-The KL objective $\mathrm{KL}(q\|\tilde{p}_{\theta})$ admits a second-order expansion around the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$):
-$\mathrm{KL}(q\|\tilde{p}_{\theta})\;\approx\;\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\operatorname{Var}_{q}\!\big[s-s^{\mathrm{ref}}-\tau_{\mathrm{anc}}\log q\big],$ (7)
-where $\delta=u-u^{\star}$ is $q$-centered (Lemma 3.5).
-The matrix $\operatorname{Diag}(q)-qq^{\top}$ is the softmax Fisher metric, inducing a local quadratic trust region whose curvature depends only on $q$ (not on the anchor $\pi_{\mathrm{ref}}$).
-This provides a distribution-space analogue of TRPO’s trust region, but implicit—no explicit KL constraint needed.
-DPO as a binary special case.
-When $|S_{x}|=2$, the anchored distribution reduces to $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}})$ where $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$.
-With Bradley–Terry targets $q(i)=\sigma(\beta_{r}\Delta R)$ and uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$), Eq.
-(5) recovers the standard DPO loss [1].
-For listwise $|S_{x}|>2$, ADPO generalizes to Plackett–Luce targets while retaining the Fisher-metric trust region.
-Unification of learning paradigms.
-Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning objectives as special cases:
-Proposition 3.2 (ADPO unifies major learning paradigms).
-Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq.
-(5).
-Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning objectives as special cases:
-  1. 1.
-    Supervised fine-tuning (SFT).
-When $q(\cdot\mid S_{x})$ is the data label distribution, $\pi_{\mathrm{ref}}$ is uniform, and $\tau_{\mathrm{anc}}=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, recovering standard cross-entropy loss.
-  2. 2.
-[…]
-  3. 3.
-    Maximum-entropy RL.
-When $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and $\pi_{\mathrm{ref}}$ is uniform, setting $\tau_{\mathrm{anc}}=1$ recovers the soft-optimal Boltzmann policy $\pi_{\theta}\propto\exp(Q/\alpha)$ [19, 21].
-  4. 4.
-    DPO (binary preferences).
-For $|S_{x}|=2$ with Bradley–Terry targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$, Eq.
-(5) reduces to DPO [1] when $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform.
-Remark 3.3 (Connection to AWR/AWAC/MPO).
-ADPO also connects to offline RL methods at the $\tilde{p}-$distribution level; see Appendix A for details on this partial equivalence.
+Because $\tilde{p}_{\theta}$ depends only on differences of anchored logits $u_{i}-u_{j}$, adding a constant to all $(\ell_{i}-\ell_{i}^{\mathrm{ref}})$ within a group leaves $\tilde{p}_{\theta}$ unchanged.
+This property automatically cancels group-level baselines such as:
+  • Annotator-specific biases (different annotators may have different baseline preferences)
+  • Sequence-level biases (e.g., length bias in language models)
+  • State-dependent baselines in RL (e.g., value function offsets)
+Remark 3.1 (Sequence tasks and length normalization).
+For sequence generation tasks, one typically constructs $\ell$ and $\ell^{\mathrm{ref}}$ using length-normalized log-probabilities to avoid length bias.
+The affine invariance of the groupwise softmax ensures all results remain valid.
+Target distribution $q$.
+The target $q(\cdot|S_{x})$ encodes the desired preference over candidates within group $S_{x}$.
+When the full action/candidate space is considered, we write $q(\cdot|x)$; the in-group distribution can be viewed as conditioning on $S_{x}$.
+Common choices include:
+  • Plackett–Luce (PL) soft labels from rankings
+  • Bradley–Terry (BT) soft labels from pairwise comparisons: $q(i\succ j|S_{x})=\sigma(\beta_{r}\Delta R_{ij})$
+  • Q-function based: $q(a|s)\propto\exp(Q(s,a)/\beta_{r})$ in RL settings
+  • Teacher distribution: $q=\pi_{\text{teacher}}$ for knowledge distillation
+3.2 Unified ADPO Objective and Its Special Cases
+Our master objective is the groupwise forward KL divergence:
+$\boxed{\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot|S_{x})\,\big\|\,\tilde{p}_{\theta}(\cdot|S_{x})\big)\Big],\quad\tilde{p}_{\theta}=\mathrm{softmax}(\bar{u}),\ \bar{u}=\tfrac{(\ell-\ell^{\mathrm{ref}})}{\tau_{\mathrm{anc}}}-\langle q,\tfrac{(\ell-\ell^{\mathrm{ref}})}{\tau_{\mathrm{anc}}}\rangle\mathbf{1}.}$ (4)
+(Equivalently, using $u$ without centering yields the same $\tilde{p}$ up to a constant shift, by softmax shift-invariance.)
+Why forward KL?
+Forward KL $\mathrm{KL}(q\|\tilde{p}_{\theta})$ is mode-covering (encourages $\tilde{p}_{\theta}$ to place mass wherever $q$ does), making it robust to soft/noisy targets; reverse KL is mode-seeking (concentrates on a single mode), which is brittle under label noise.
+For listwise supervision with uncertain preferences, mode-covering is more stable.
+Within a group, this equals a cross-entropy up to a constant.
+Writing $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ is the log-partition function, the loss is strictly convex in $u$ (on $\mathbf{1}^{\perp}$):
+$\sum_{i}q_{i}\log\tilde{p}_{\theta,i}\;=\;-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i}q_{i}\big(\ell_{i}-\ell_{i}^{\mathrm{ref}}\big)\;+\;\log\!\sum_{j}\exp\!\Big(\frac{\ell_{j}-\ell_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)\;=\;\mathrm{KL}(q\|\tilde{p}_{\theta})+H(q).$ (5)
+Closed-form optimum in distribution space.
+If the model class is expressive enough, the optimum of (4) satisfies:
+Proposition 3.2 (Closed-form optimum).
+When $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ is attainable, the optimal policy satisfies
+$\pi_{\theta}(y|x)\;\propto\;\pi_{\mathrm{ref}}(y|x)\,q(y|x)^{\tau_{\mathrm{anc}}}.$ (6)
+This requires mild support conditions: in the discrete case, $q(y|x)>0\Rightarrow\pi_{\mathrm{ref}}(y|x)>0$; in the continuous case, absolute continuity and integrability ensure $\tilde{p}=q$ is attainable and the proportionality is normalizable (well-defined).
+When training on sampled groups $S_{x}$, this gives the conditional optimum within each group; if groups are uniformly sampled from the full action space and the model is sufficiently expressive, the empirical optimum converges to the population optimum.
+Proof sketch.
+At optimum, $\tilde{p}_{\theta}(i|S_{x})=q(i|S_{x})$. By definition,
+$\tilde{p}_{\theta}(i|S_{x})=\mathrm{softmax}\!\Big(\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)\propto\exp\!\Big(\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)=\Big(\frac{\pi_{\theta}(i|x)}{\pi_{\mathrm{ref}}(i|x)}\Big)^{1/\tau_{\mathrm{anc}}}.$
+Setting this equal to $q(i|S_{x})$ and solving yields (6). ∎
+Temperature interpretation.
+  • Small $\tau_{\mathrm{anc}}$: Strong regularization toward anchor $\pi_{\mathrm{ref}}$ (conservative updates).
+The local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$ (Section 3.3), yielding a tighter implicit trust region.
+  • Large $\tau_{\mathrm{anc}}$: Weak regularization, $\pi_{\theta}$ tracks $q$ more closely (aggressive updates)
+  • $\tau_{\mathrm{anc}}=1$: Balanced regime (commonly used in practice)
+Pairwise and DPO as corollaries.
+For $|S_{x}|{=}2$, let $\Delta_{\theta}=\ell_{i}-\ell_{j}$ and $\Delta_{\mathrm{ref}}=\ell_{i}^{\mathrm{ref}}-\ell_{j}^{\mathrm{ref}}$. With a Bradley–Terry target $q_{ij}\in(0,1)$, (4) reduces to:
+$\boxed{\mathcal{L}_{ij}^{\text{ADPO}}=\log\!\big(1+\exp[(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}]\big)-q_{ij}\,(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}},}$ (7)
+which recovers the standard DPO loss [1] when $\Delta_{\mathrm{ref}}=0$ (uniform anchor) and $q_{ij}\in\{0,1\}$ (hard labels).
+When $q_{ij}\in(0,1)$ (soft labels from noisy annotations), this becomes "soft DPO" that naturally handles label uncertainty.
+When $\Delta_{\mathrm{ref}}\neq 0$ (non-uniform anchor), we obtain "anchored DPO" with improved stability.
+Gradient.
+The gradient of the pairwise loss is:
+$\frac{\partial\mathcal{L}_{ij}^{\mathrm{ADPO}}}{\partial\theta}=\frac{1}{\tau_{\mathrm{anc}}}\big(\sigma[(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}]-q_{ij}\big)(\nabla_{\theta}\ell_{i}-\nabla_{\theta}\ell_{j}),$
+where $\sigma(\cdot)$ is the sigmoid function.
+Listwise (general) case.
+For candidate sets $S_{x}=\{y_{1},\dots,y_{P}\}$ with $P>2$, the ADPO listwise objective is:
+$\boxed{\mathcal{L}_{\text{group}}^{\text{ADPO}}=-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x}),\qquad\tilde{p}_{\theta}(\cdot|S_{x})=\mathrm{softmax}(\bar{u}),}$ (8)
+with gradient
+$\frac{\partial\mathcal{L}_{\mathrm{group}}^{\text{ADPO}}}{\partial\ell_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$ (9)
+Setting P=2P{=}2 recovers the binary case (7), showing that ADPO continuously extends DPO from pairwise to listwise soft supervision.
+$q$-centering: A principled use of shift-invariance.
+Recall from Section 3.1 that softmax is shift-invariant: $\mathrm{softmax}(u)=\mathrm{softmax}(u+c\mathbf{1})$ for any constant $c$.
+We exploit this property by choosing $c=-\sum_{j}q_{j}u_{j}$ (the negative $q$-weighted mean), yielding:
+$\boxed{\bar{u}_{i}=u_{i}-\sum_{j\in S_{x}}q(j|S_{x})\,u_{j},\qquad u_{i}=\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}.}$ (10)
+This $q$-centering is always applied in our algorithm (Algorithm 1, both modes) and serves three purposes: numerical stability (centers logits near zero); geometric consistency (removes null-space along $\mathbf{1}$ in Fisher metric, Section 3.3); theoretical alignment (consistent with $q$-centered deviation $\bar{\delta}$ in Lemma 3.6).
+Our final anchored distribution is thus:
+$\tilde{p}_{\theta}(i|S_{x})=\mathrm{softmax}(\bar{u})_{i}=\frac{\exp(\bar{u}_{i})}{\sum_{j\in S_{x}}\exp(\bar{u}_{j})}.$ (11)
+Remark 3.3 (Optional: Fisher variance scaling).
+For extremely large $P$ (e.g., $P\geq 16$), one can additionally apply Fisher normalization:
+$\widehat{u}_{i}=\frac{\bar{u}_{i}}{\sqrt{\mathrm{Var}_{q}[u]+\varepsilon}},\qquad\mathrm{Var}_{q}[u]=\sum_{j}q_{j}u_{j}^{2}-\Big(\sum_{j}q_{j}u_{j}\Big)^{2}.$ (12)
+This "whitens" the gradient but changes the effective temperature. We do not use this in our experiments and recommend it only as a troubleshooting step for very large $P$.
+Unifying major learning paradigms.
+Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning paradigms:
+Proposition 3.4 (ADPO unifies common paradigms).
+Specific choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover:
+  • Supervised Fine-Tuning (SFT): $q$ from labeled data, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=1$.
+  • Knowledge Distillation (KD): $q=\mathrm{softmax}(z_{T}/T)$ from teacher logits $z_{T}$, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=T$.
+  • Maximum-Entropy RL: $q(a|s)\propto\exp(Q/\alpha)$ from Q-values, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=1$.
+  • Direct Preference Optimization (DPO): $|S_{x}|=2$, $q(i)=\sigma(\beta_{r}\Delta R)$ from reward differences, $\pi_{\mathrm{ref}}$ uniform.
+Verification.
+Each case follows from Proposition 3.2 by substituting the specified $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$. ∎
+Table 1 summarizes this unification.
 Table 1: ADPO as a unifying projection framework.
 Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning paradigms.
-Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop.
-3.4).
+Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop.
+3.2).
+Caution: When anchor and target both equal the teacher ($\pi_{\mathrm{ref}}=\pi_{T},q=\pi_{T}$), the optimum becomes $\pi_{\theta}\propto\pi_{T}^{1+\tau_{\mathrm{anc}}}$, which sharpens the teacher distribution and may degrade calibration; use small $\tau_{\mathrm{anc}}$ or add auxiliary regularization if needed.
 Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau_{\mathrm{anc}}$ Recovers
 SFT data (one-hot/soft) uniform 11 cross-entropy
-KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(s/T))$
+KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(\ell/T))$
 KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any teacher-anchored KD
@@
 ADPO (listwise) Plackett–Luce / soft any any this work
-Proposition 3.4 (Closed-form optimal solution).
-Suppose $S_{x}$ covers the full action/class space and $q(\cdot\mid S_{x})\succ 0$.
-(Discrete case) If the policy $\pi_{\theta}$ has sufficient expressivity (i.e., the parametrization can represent arbitrary log-odds over $S_{x}$), then the distribution-level optimum of $\operatorname*{arg\,min}_{\theta}\mathrm{KL}(q\|\tilde{p}_{\theta})$ satisfies
-$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (8)
-(Continuous case) For continuous action spaces, assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure (e.g., Lebesgue measure).
-Define the anchored distribution via the density ratio:
-$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau_{\mathrm{anc}}}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau_{\mathrm{anc}}}.$
-Setting $\tilde{p}_{\theta}=q$ yields the closed-form solution
-$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}},$
-which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}}\,da<\infty$ to ensure normalizability.
-Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
-Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau_{\mathrm{anc}}}$.
-The constant $c$ is absorbed into the proportionality constant.
+Warning: Sharpens teacher; see caption.
+3.3 Fisher Geometry: An Implicit Trust Region
+The anchored softmax structure induces a natural Riemannian geometry on the probability simplex.
+Let $\tilde{p}_{\theta}$ (hereafter $p$ for brevity) denote the current anchored distribution.
+The softmax Fisher information matrix is:
+$F\;=\;\operatorname{Diag}(p)-pp^{\top}\;\in\;\mathbb{R}^{P\times P},\qquad F\mathbf{1}{=}0,$ (13)
+where $F\mathbf{1}=0$ reflects the normalization constraint and shift-invariance.
+Near the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the KL divergence admits a second-order expansion:
+$\mathrm{KL}(q\|\tilde{p}(u))\;\approx\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\bar{\delta}^{\top}(\operatorname{Diag}(q)-qq^{\top})\bar{\delta},\qquad\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1},$ (14)
+where the $1/\tau_{\mathrm{anc}}^{2}$ scaling governs the local curvature (see Figure 1).
+This creates an implicit trust region centered at the target $q$, independent of the anchor $\pi_{\mathrm{ref}}$—the anchor defines the coordinate system, while the Fisher metric determines the geometry.
+Figure 1: Fisher geometry as implicit trust region.
+The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ defines a quadratic bowl in anchored coordinates $u=(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$, centered at target $q$.
+Smaller $\tau_{\mathrm{anc}}$ yields tighter curvature ($\propto 1/\tau_{\mathrm{anc}}^{2}$).
+Natural gradient and pullback metric.
+The natural gradient $\delta u_{\text{nat}}=F^{+}(q-p)$ (where $F^{+}$ is the Moore–Penrose pseudoinverse) gives the steepest descent direction in the Fisher metric [27].
+In parameter space, the pullback metric is:
+$\mathsf{G}(\theta;S_{x})=J\,F\,J^{\top}=\sum_{i}p_{i}g_{i}g_{i}^{\top}-\Big(\sum_{i}p_{i}g_{i}\Big)\Big(\sum_{i}p_{i}g_{i}\Big)^{\top},\quad J=[g_{1},\ldots,g_{P}],\ g_{i}=\nabla_{\theta}\ell_{i}.$ (15)
+This is exactly the Fisher information matrix used in TRPO [23], but ADPO induces it implicitly through the anchored softmax—no explicit second-order optimization is required.
+Remark 3.5 (Theory vs practice).
+Our implementation (Algorithm 1, both modes) uses standard first-order optimization (Adam) on the ADPO loss (4).
+The Fisher geometry provides theoretical insight into why this works well, not how we implement it.
+3.4 Mirror-Descent Equivalence and Convergence
+The ADPO update can be viewed as mirror descent with KL Bregman divergence.
+Within each group, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is convex in anchored logits $u=(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$.
+Lemma 3.6 (Implicit trust region).
+Near the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), writing $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$,
+$\mathrm{KL}\big(q\|\tilde{p}(u)\big)\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\bar{\delta}^{\top}\big(\operatorname{Diag}(q)-qq^{\top}\big)\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$ (16)
+Proof sketch.
+Second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p})-\tilde{p}\,\tilde{p}^{\top}$ yields the stated form.
+The $q$-centering removes the null-space along $\mathbf{1}$.
 ∎
-Implication and reachability.
-Equation (8) strictly characterizes the "anchored projection" as a multiplicative fusion of reference and target when the optimum is reachable.
-In discrete finite-action settings with tabular or sufficiently expressive function approximation, this optimum is exactly attainable.
-However, in structured settings (e.g., autoregressive language models with shared parameters across tokens, or policy networks with limited capacity), the optimum may be unreachable, and gradient descent performs an approximate projection in the parametric family.
-The KD, MaxEnt, and AWR special cases in Proposition 3.2 all follow directly from (8).
-Listwise subset constraints and convexity.
-When $S_{x}$ is only a subset (pairwise/listwise ranking), (8) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints.
-Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ within each group (as shown in Theorem 3.8), but the optimization over shared parameters $\theta$ is typically non-convex.
-The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot\mid S_{x})=q(\cdot\mid S_{x})$ for all groups (except in the tabular case or with unlimited expressivity).
-Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
-By centering the coordinate system at $\pi_{\mathrm{ref}}$ before applying softmax, ADPO provides groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.5)—properties absent in standard KD or unanchored DPO.
-Table 2 summarizes the unification.
-Mirror descent = KL proximal step.
-If one directly optimizes over the anchored logits $u$ (e.g., tabular setting), each gradient step is equivalent to a KL-trust-region proximal problem (Theorem 3.7):
-$u_{t+1}=\arg\min_{u}\ \langle\nabla\mathcal{L}(u_{t}),\,u\rangle+\tfrac{1}{\eta}\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u)),$
-whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$. This exhibits ADPO as a Bregman-proximal method in distribution space.
-Dynamic anchoring: monotonicity within fixed-anchor phases.
-When the anchor $\pi_{\mathrm{ref}}$ is held fixed, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is strictly convex on the subspace $\mathbf{1}^{\perp}$.
-Under either (i) local analysis near $u^{\star}$ or (ii) bounded-domain constraints $\tilde{p}_{\theta}\in[\varepsilon,1-\varepsilon]^{K}$, $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$, guaranteeing linear convergence $\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star}))$ within each fixed-anchor phase (Theorem 3.8).
-When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes, and we do not claim monotonicity across anchor updates—only within each phase.
-Standing assumptions.
-Throughout this work, we maintain the following conditions unless otherwise noted:
-  1. (i)
-    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i\mid x)>0\Rightarrow\pi_{\mathrm{ref}}(i\mid x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ are well-defined.
-  2. (ii)
-    Absolute continuity (continuous): For continuous action spaces, we assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure, with overlapping support in the same sense as (i).
-  3. (iii)
-    Local analysis vs.
-global bounds: Fisher-metric trust region results (Lemma 3.5, Corollary 3.6) hold locally around $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$.
-For global strong convexity, we either work in a neighborhood of $u^{\star}$ or assume the optimization is constrained to a bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ (Theorem 3.8).
-  4. (iv)
-    Sequence models: For autoregressive sequence generation, we use length-normalized (per-token average) log-probabilities $s$ and $s^{\mathrm{ref}}$ as group-level scores, and apply the softmax Fisher geometry at the sequence-level distribution over candidates.
-  5. (v)
-    Expressivity: Closed-form optimal solutions (Proposition 3.4) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$.
-In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
-3.2 Pairwise Instantiation
-Notation for this subsection: $\beta$ = student inverse temperature, $\beta_{r}$ = teacher/reference inverse temperature, $q_{ij}$ = soft preference probability that $i\succ j$.
-For a candidate pair $(i,j)$ with soft teacher preference $q_{ij}\in(0,1)$ that $i$ is preferred, the ADPO pairwise loss is:
-$\ell_{ij}^{\text{Soft-DPO}}=\log\left(1+e^{\beta(\Delta_{\theta}-\Delta_{\text{ref}})}\right)-q_{ij}\,\beta(\Delta_{\theta}-\Delta_{\text{ref}}),$ (9)
-where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(y_{i}|x)$.
-Key properties:
-  • Bayes-optimal matching: Gradient vanishes when $\sigma(\beta(\Delta_{\theta}-\Delta_{\text{ref}}))=q_{ij}$.
-  • Soft weighting: Uncertain pairs ($q_{ij}\approx 0.5$) contribute less gradient.
-  • Reference anchoring: Relative updates $\Delta_{\theta}-\Delta_{\text{ref}}$ are invariant to groupwise shifts.
-Temperature alignment: The pairwise formulation uses inverse temperature $\beta$, which relates to the listwise anchoring temperature via $\beta=1/\tau_{\mathrm{anc}}$.
-Specifically, $\tilde{p}_{\theta}(i)=\sigma\!\big((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)$.
-3.3 Listwise Instantiation
-Notation for this subsection: $\tau_{\mathrm{anc}}$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
-For group $S_{x}=\{y_{1},\ldots,y_{P}\}$, the ADPO listwise loss is:
-$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})\right],$ (10)
-where
-$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau_{\mathrm{anc}}\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau_{\mathrm{anc}}\right)},$ (11)
-and teacher target $q(i\mid S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
-Equivalent expansion. The anchored listwise cross-entropy (10) can be equivalently written as:
-$\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})=-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i\in S_{x}}q(i\mid S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\right).$ (12)
-This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\rangle$ and a log-sum-exp normalization.
-It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below.
-Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (12) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
-3.4 Formal Statements of Core Results
-We now provide formal statements and proofs for the results sketched in Section 3.
-Lemma 3.5 (Implicit trust region).
-Let $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ and $\tilde{p}(u)=\mathrm{softmax}(u)$.
-Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$.
-Define $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$, the $q$-centered deviation.
-Then, as $u\to u^{\star}$,
-$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$
-The curvature is given by the softmax Fisher metric $\operatorname{Diag}(q)-qq^{\top}$, which is independent of the anchor choice and has a null-space along $\mathbf{1}$.
+Theorem 3.7 (Mirror-descent equivalence).
+The ADPO update with step size $\eta$ satisfies:
+$u_{t+1}=\arg\min_{u}\big\langle\nabla\mathcal{L}(u_{t}),\,u\big\rangle+\frac{1}{\eta}\,\mathrm{KL}\!\big(\tilde{p}(u_{t})\,\|\,\tilde{p}(u)\big),$
+inducing the multiplicative rule $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$ on distributions. See Appendix for full derivation.
+Corollary 3.8 (Dynamic anchor as proximal trust region).
+When $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (dynamic anchor), each step controls $\mathrm{KL}(\tilde{p}_{\text{old}}\|\tilde{p}_{\text{new}})$ in anchored coordinates, mirroring TRPO’s $\mathrm{KL}(\pi_{\text{old}}\|\pi_{\text{new}})$ constraint with groupwise shift invariance.
+The temperature $\tau_{\mathrm{anc}}$ scales curvature as $1/\tau_{\mathrm{anc}}^{2}$, acting as a radius knob.
+3.5 Anchor Strategies: Offline vs Online
+The flexibility of ADPO’s anchoring mechanism enables different instantiations depending on the choice of anchor policy $\pi_{\mathrm{ref}}$.
+Because the Fisher metric depends only on the target $q$ (not on $\pi_{\mathrm{ref}}$), we can cleanly separate the choice of anchor from the learning objective, leading to two main paradigms.
+Offline ADPO: Fixed or no anchor.
+In offline settings (SFT, KD, preference learning from fixed datasets), the anchor is either:
+  • No anchor (uniform): Set $\pi_{\mathrm{ref}}$ uniform, equivalent to $\ell^{\mathrm{ref}}_{i}=0$. This recovers:
+    • SFT: $q$ from labeled data, $\tau_{\mathrm{anc}}=1$ → standard cross-entropy
+    • KD: $q=\mathrm{softmax}(z_{T}/T)$ from teacher logits, $\tau_{\mathrm{anc}}=T$ → temperature-scaled KD
+  • Fixed anchor: Set $\pi_{\mathrm{ref}}$ once (e.g., pretrained model, SFT checkpoint) and never update. The coordinate system remains fixed, providing a stable geometric center for:
+    • Aligning to labels/rewards/preferences
+    • Regularizing toward a known good policy
+    • Preventing catastrophic forgetting
+Key property: With fixed/no anchor, the optimization has monotonic convergence guarantees (Theorem 3.10).
+Online ADPO: Dynamic anchor.
+In online settings (RL, iterative RLHF, continual learning), the anchor is dynamically updated to track the evolving policy:
+$\boxed{\mathcal{L}_{\text{ADPO-online}}\;=\;\mathbb{E}_{x}\Big[\mathrm{KL}\!\Big(q(\cdot|x)\ \Big\|\ \mathrm{softmax}\!\big((\ell_{\theta}-\ell_{\text{old}})/\tau_{\mathrm{anc}}\big)\Big)\Big],\quad\pi_{\mathrm{ref}}\leftarrow\pi_{\text{old}}.}$ (17)
+(The anchored logits $(\ell_{\theta}-\ell_{\text{old}})/\tau_{\mathrm{anc}}$ are $q$-centered as in Eq. 10.)
+Anchor update strategies:
+  • Per-iteration: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ every iteration (like TRPO [23]/PPO [4])
+  • Periodic: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ every $N$ iterations (balances stability/adaptivity)
+  • EMA: $\pi_{\mathrm{ref}}\leftarrow\alpha\pi_{\mathrm{ref}}+(1-\alpha)\pi_{\theta}$ (smooth evolution)
+Key property: The anchoring defines a moving coordinate system, while the Fisher metric (determined by $q$) provides an implicit trust region.
+This mirrors TRPO’s $\mathrm{KL}(\pi_{\text{old}}\|\pi_{\text{new}})$ constraint [23] but operates in distribution space with groupwise shift invariance.
+Task-dependent trade-off.
+Our experiments (Section 5) reveal:
+  • Dynamic anchors improve online exploration in noisy environments by 5–11% in 3/4 scenarios (Section 5.4). The adaptive coordinate frame enables effective exploration while maintaining stability.
+[…]
+Remark 3.9 (Connection to existing methods).
+Offline: No anchor recovers SFT/KD; fixed anchor adds regularization toward a reference policy.
+Online: Dynamic anchor connects to TRPO/PPO but with implicit Fisher-metric trust regions and groupwise shift invariance.
+3.6 Stability and Convergence Analysis
+Having established the geometric structure (Section 3.3) and algorithmic view (Section 3.4) of ADPO, we now analyze its stability and convergence properties.
+Fixed anchor: Strong convergence guarantees.
+When the anchor $\pi_{\mathrm{ref}}$ is fixed throughout training, ADPO enjoys strong convexity and linear convergence within the anchored coordinate system.
+Theorem 3.10 (Monotonic descent under fixed anchor).
+With a fixed anchor, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex and $L$-smooth on the bounded domain where $\tilde{p}(u)\in[\varepsilon,1-\varepsilon]^{P}$.
+For step size $\eta\in(0,1/L]$, gradient or mirror descent satisfies
+$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big),$ (18)
+yielding linear convergence within each fixed-anchor phase.
 Proof sketch.
-Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$.
-Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$.
-A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form.
-The centering removes the null-space along $\mathbf{1}$, reflecting the invariance of softmax to additive shifts.
-∎
-Corollary 3.6 (Spectral bounds).
-Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\operatorname{Diag}(q)-qq^{\top}$.
-In the quotient space $\mathbb{R}^{K}/\mathrm{span}\{\mathbf{1}\}$, the objective is strongly convex.
-Locally around $u^{\star}$,
-$\mathrm{KL}(q\|\tilde{p}(u))\;\geq\;\frac{\mu(q)}{2}\,\|\bar{\delta}\|^{2},$
-where $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$. By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$.
-For $K=2$: $\mu(q)=2q(1-q)$.
-For $K>2$, when $\min_{i}q_{i}\geq\varepsilon>0$, there exists a constant $c(\varepsilon,K)>0$ such that $\mu(q)\geq c(\varepsilon,K)$ on the quotient space.
-In particular, for any $v$ satisfying $\langle q,v\rangle=0$,
-$\operatorname{Var}_{q}[v]=\sum_{i}q_{i}v_{i}^{2}\geq\varepsilon\|v\|^{2}.$
-Geometric interpretation.
-The anchor $\pi_{\mathrm{ref}}$ defines the coordinate origin, while the target $q$ determines the minimizer $u^{\star}$ and local curvature.
-This provides a distribution-space analogue of TRPO but with an implicit (Fisher-metric) trust region—no explicit KL constraint needed.
-Figure 1: ADPO’s anchored trust region.
-The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself.
-The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift.
-This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
-3.5 KL-Mirror Descent Interpretation
-Theorem 3.7 (KL-mirror descent / Bregman-prox step).
-If one directly optimizes $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ over anchored logits $u$ (where $A(u)=\log\sum_{j}e^{u_{j}}$), then for any step size $\eta>0$, the one-step update is equivalent to:
-$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle\nabla\mathcal{L}(u_{t}),\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big),$ (13)
-whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$.
-Proof.
-A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$.
-Using the Bregman divergence relation $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (13).
-∎∎
-3.6 Dynamic Anchor Stability
-The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ is anchor-agnostic (depends only on $q$), allowing flexible anchor strategies: (i) fixed $\pi_{\text{ref}}$ (frozen, stable origin); (ii) dynamic $\pi_{\theta_{\text{old}}}$ (periodic updates, adaptive like TRPO); (iii) teacher $\pi_{\text{teacher}}$ (offline distillation).
-We characterize monotonic improvement within each fixed-anchor phase:
-Theorem 3.8 (Fixed-anchor-phase monotonic descent).
-When the anchor is fixed, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex on the quotient space and $L$-smooth on any bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$.
-Consequently, gradient descent with step size $\eta\in(0,1/L]$ (or KL-mirror descent with the corresponding step-size condition) enjoys the linear rate
-$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big)$
-within that fixed-anchor phase. When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes—we do not claim monotonicity across anchor updates.
-Proof sketch.
-$A(u)$ is strictly convex with Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ (softmax Fisher metric).
-Near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, this coincides with $\operatorname{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity.
-In bounded domains where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$, the Fisher metric satisfies $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$ in Corollary 3.6).
-The spectral upper bound of $\nabla^{2}A(u)$ is also bounded in this domain, yielding $L$-smoothness with $L=L(\varepsilon,K)$.
-∎∎
-Implication.
-Within each fixed-anchor phase, Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts (like TRPO).
-Experiments (Section 5) reveal task-dependent trade-offs: dynamic anchors excel at online exploration; fixed anchors excel at offline distillation.
-Gradients.
-Pairwise: $\frac{\partial\ell_{ij}}{\partial\theta}=\beta\big(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij}\big)\big(\nabla_{\theta}s_{i}-\nabla_{\theta}s_{j}\big).$
-Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i\mid S_{x})-q(i\mid S_{x})\big).$
+The Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p})-\tilde{p}\,\tilde{p}^{\top}$ has eigenvalues in $\min_{i}\tilde{p}_{i}(1-\tilde{p}_{i}),\max_{i}\tilde{p}_{i}(1-\tilde{p}_{i})]$.
+On the bounded domain $\varepsilon,1-\varepsilon]^{P}$, this yields:
+  • Strong convexity: $\mu\geq\varepsilon(1-\varepsilon)$
+  • Smoothness: $L\leq 1/4$
+Standard convergence rates for strongly convex and smooth functions then give the stated linear convergence [28]. ∎
+Corollary 3.11 (Convergence rate).
+With optimal step size $\eta=1/L$, the number of iterations to reach $\varepsilon$-accuracy is:
+$T=O\!\Big(\frac{L}{\mu}\log\frac{1}{\varepsilon}\Big)\;=\;O\!\Big(\frac{1}{\varepsilon(1-\varepsilon)}\log\frac{1}{\varepsilon}\Big).$
+For well-separated distributions (large $\varepsilon$), convergence is fast.
+Dynamic anchor: Coordinate frame changes.
+When the anchor updates ($\pi_{\mathrm{ref}}\!\leftarrow\!\pi_{\theta}$), the coordinate frame changes and monotonicity is not guaranteed across updates. However, local stability is preserved.
+Theorem 3.12 (Local stability under dynamic anchor).
+Let $\pi_{\mathrm{ref}}^{(t)}$ denote the anchor at iteration $t$, and define anchored logits $u^{(t)}=(\ell_{\theta}-\ell_{\mathrm{ref}}^{(t)})/\tau_{\mathrm{anc}}$.
+Within each anchor phase $t_{k},t_{k+1})$, the loss $\mathcal{L}^{(k)}(u^{(t)})$ decreases monotonically (Theorem 3.10).
+At anchor update $t_{k+1}$, the coordinate system shifts, and the loss $\mathcal{L}^{(k+1)}(u^{(k+1)})$ is measured in the new coordinate frame.
+Global monotonicity $\mathcal{L}^{(k+1)}\leq\mathcal{L}^{(k)}$ is not guaranteed, but the Fisher-metric constraint ensures local stability: the policy remains within a trust region around the previous policy.
+Remark 3.13 (Comparison to TRPO/PPO).
+TRPO [23] enforces monotonic improvement through explicit KL constraints, while PPO [4] uses clipping to approximate this.
+Online ADPO achieves similar stability through implicit Fisher geometry, without requiring explicit constraints or clipping.
+The key difference is that ADPO operates in distribution space (anchored coordinates), while TRPO/PPO operate in parameter space.
+3.7 Algorithm and Implementation
+We now provide a unified algorithmic description of ADPO that covers both online RL and offline preference learning through a mode parameter.
+Algorithm: ADPO (Unified).
+Algorithm 1 ADPO (Unified): Online RL and Offline Preference Learning
+1:Mode $\in\{\textsc{online},\textsc{offline}\}$; temperature $\tau_{\mathrm{anc}}$; learning rate(s) $\eta_{\pi}$ (and $\eta_{Q}$ for online); anchor update period $N_{\mathrm{anchor}}$; reward temperature $\beta_{r}$ (online only)
+2:Initialize policy $\pi_{\theta}$; set anchor $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$
+3:if online then
+4:  Initialize critic $Q_{\phi}$, replay buffer $\mathcal{B}$
+5:else
+6:  Load dataset $\mathcal{D}=\{(x,S_{x},\text{supervision})\}$
+7:end if
+8:for training loop $t=1,\ldots,T$ $\triangleright$ iteration for online; epoch for offline do
+9:  if online then
+10:   Collect trajectories using $\pi_{\theta}$, add to $\mathcal{B}$
+11:  end if
+12:  for batch update do
+13:   if online then
+14:     Sample states $\{x_{j}\}_{j=1}^{B}\sim\mathcal{B}$
+15:   else
+16:     Sample batch $\{(x_{j},S_{x_{j}},\text{sup}_{j})\}_{j=1}^{B}\sim\mathcal{D}$
+17:   end if
+18:   for each context $x_{j}$ in batch do
+19:     // Build target distribution $q$
+20:     if online then
+21:      Sample candidates $\{y_{1},\ldots,y_{P}\}$: $y_{i}\sim\pi_{\theta}(\cdot|x_{j})$ or mixed sampling
+22:      Compute scores: $R_{i}\leftarrow Q_{\phi}(x_{j},y_{i})$ or $r(x_{j},y_{i})$
+23:      Normalize (Eq. 19): $\tilde{R}_{i}\leftarrow(R_{i}-\mathrm{median}(R))/(\mathrm{MAD}(R)+\varepsilon)$
+24:      Build target: $q(i|x_{j})\leftarrow\mathrm{softmax}(\tilde{R}/\beta_{r})_{i}$
+25:     else
+26:      Use candidate set $S_{x_{j}}$ from data
+27:      Compute target: $q(i|S_{x_{j}})$ from Bradley-Terry, Plackett-Luce, reward model, etc.
+28:     end if
+29:     $\triangleright$ ===== Core ADPO Update =====
+30:     Compute log-probabilities: $\ell_{i}\leftarrow\log\pi_{\theta}(y_{i}|x_{j})$, $\ell_{i}^{\mathrm{ref}}\leftarrow\log\pi_{\mathrm{ref}}(y_{i}|x_{j})$
+31:     $u_{i}\leftarrow(\ell_{i}-\ell_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$
+32:     $q$-centering (Eq. 10): $\bar{u}_{i}\leftarrow u_{i}-\sum_{k\in S_{x_{j}}}q(k|x_{j})\,u_{k}$
+33:     $\tilde{p}_{\theta}(i|x_{j})\leftarrow\mathrm{softmax}(\bar{u})_{i}$
+34:   end for
+35:   Compute loss: $\mathcal{L}_{\pi}\leftarrow-\frac{1}{B}\sum_{j=1}^{B}\sum_{i\in S_{x_{j}}}q(i|x_{j})\log\tilde{p}_{\theta}(i|x_{j})$
+36:   Update policy: $\theta\leftarrow\theta-\eta_{\pi}\nabla_{\theta}\mathcal{L}_{\pi}$
+37:  end for
+38:  if online then
+39:   Update critic $Q_{\phi}$ using standard RL (e.g., TD, SAC)
+40:  end if
+41:  // Update anchor
+42:  if $t\bmod N_{\mathrm{anchor}}=0$ then
+43:   $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ $\triangleright$ dynamic anchor (TRPO-style)
+44:  end if
+45:end for
+46:return $\pi_{\theta}$
+Practical notes.
+  • $q$-centering (always): Apply Eq. 10 to center anchored logits at target $q$, ensuring numerical stability and Fisher-metric consistency.
+  • Reward normalization (online RL only): In online mode (Algorithm 1), standardize rewards before building $q$:
+    $\tilde{R}_{i}=\frac{R_{i}-\mathrm{median}(R)}{\mathrm{MAD}(R)+\varepsilon},\qquad q(i|x)=\mathrm{softmax}(\tilde{R}/\beta_{r}),$ (19)
+    where $\mathrm{MAD}(R)=\mathrm{median}(|R_{i}-\mathrm{median}(R)|)$. Not needed in offline mode where $q$ is fixed.
+  • Anchor caching: Cache $\ell^{\mathrm{ref}}$ on batch data to avoid redundant forward passes.
+  • Numerical stability: Use log-sum-exp trick for $\log\sum_{j}\exp(u_{j})$.
+  • Temperature: In KD scenarios, $\tau_{\mathrm{anc}}\equiv T$ (teacher temperature); see Table 1.
+  • Batch processing: Process groups independently; no cross-group dependencies.
 4 Experimental Setup
@@
 Higher is better; random baseline = $1/P$.
+Computed over 1000 test contexts per scenario.
   • NDCG@k: Normalized Discounted Cumulative Gain at position $k$, a standard ranking metric that rewards placing high-reward items at top positions with logarithmic discounting.
 Higher is better; range $[0,1]$.
+We report NDCG@4 for $P=4$ candidates.
   • KL divergence: For distillation tasks, we measure $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\mathrm{student}})$, the expected KL divergence from teacher to student policy over states and actions.
 Lower is better, indicating closer alignment with the teacher.
+Computed over 10K states sampled from the replay buffer.
+Reporting standards.
+All results report mean $\pm$ standard error over 10 random seeds (5 seeds for continuous control).
+We verify statistical significance using Wilcoxon signed-rank tests (p<0.05) for all claimed improvements; detailed confidence intervals and p-values are provided in the Appendix.
+Training steps and sample counts are aligned across all methods within each experiment.
+Student–teacher KL.
+We report the group-averaged divergence
+$\mathbb{E}_{x,S_{x}}\Big[\mathrm{KL}\big(\pi_{T}(\cdot\mid S_{x})\|\pi_{\theta}(\cdot\mid S_{x})\big)\Big],$
+where both policies are renormalized on the candidate set $S_{x}$ before computing KL.
 $2\times 2$ base design + listwise extensions. We systematically compare:
@@
   • Label type: Soft ($q_{ij}\in(0,1)$ via Bradley–Terry) vs. Hard (winner=1, loser=0)
-The $2\times 2$ base covers pairwise methods (4 combinations), plus 3 ADPO listwise extensions (Raw/KDE/KDE-Rank aggregating full distributions), yielding 7 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE/KDE-Rank.
+The $2\times 2$ base covers pairwise methods (4 combinations), plus 2 ADPO listwise extensions (Raw/KDE aggregating full distributions), yielding 6 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE.
 Scenarios and difficulty levels. We test 4 noise types $\times$ 3 severity levels:
   1. (i)
-    Heavy Noise: Gaussian noise with outliers. Light (SNR=1.0, 5% outliers), Medium (SNR=0.5, 10%), Heavy (SNR=0.2, 20%).
+    Heavy Noise: Gaussian noise with outliers. Light (noise scale=1.0, 5% outliers), Medium (scale=2.0, 10%), Heavy (scale=2.5, 20%).
   2. (ii)
@@
 All tables report test WinMass under the shifted/noisy process, with $P=4$ fixed throughout.
-Model architecture. Policy is an MLP: $s_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
+Model architecture. Policy is an MLP: $\ell_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
   • Small: hidden=64, layers=2 (total $\sim$8K params)
@@
 5.1 $2\times 2$ Base Comparison Across Scenarios
-Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 79% (10 seeds).
-The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 9/12 settings.
+Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 93% (10 seeds).
+The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 11/12 settings.
 Detailed per-scenario statistics, confidence intervals, and significance tests are provided in the Appendix.
@@
 Underline: best among pairwise.
-"Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank.
-In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
+"Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE.
+In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 93% over Standard DPO baseline (mean of Std-Soft/Hard).
 Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard Best Listwise
@@
 Figure 2: Comprehensive $2\times 2$ comparison across 12 scenarios (10 seeds each).
-Each subplot shows convergence curves for 7 methods.
-Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods (dashed lines) achieve highest final performance in 9/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row); (iv) anchored methods show faster convergence and higher stability.
+Each subplot shows convergence curves for 6 methods.
+Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods achieve highest final performance in 11/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row), reaching 93% improvement under Heavy Noise-Heavy; (iv) anchored methods show faster convergence and higher stability.
 Error bands: mean $\pm$ s.e.
@@
 5.2 Listwise vs. Pairwise Methods
-Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios (Table 3). Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
+Listwise methods (ADPO Listwise-Raw/KDE) achieve the highest WinMass in 11 out of 12 scenarios (Table 2). Peak performance: 0.826 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
   • Use full group information (all $P$ items) vs. pairwise’s $O(P)$ sampled pairs.
@@
 Winner highlighted.
-Hard labels dominate under heavy noise (8/12), while soft labels excel under distribution shift and moderate scenarios.
+Hard labels dominate under heavy noise (7/12), while soft labels excel under distribution shift and moderate scenarios.
 Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard
-Heavy Noise Heavy 0.430 0.431 0.702 0.770
-Dist. Shift Light 0.712 0.740 0.841 0.794
-Adversarial Medium 0.629 0.658 0.748 0.730
-Heavy-Tailed Heavy 0.458 0.472 0.784 0.746
-Hard wins: 8/12 scenarios (Heavy Noise all, Adversarial 2/3, Dist. Shift 2/3)
-Soft wins: 4/12 scenarios (Heavy-Tailed 2/3, Adversarial 1/3, Dist. Shift 1/3)
-Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 4).
-Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs.
-ADPO-Soft’s 0.702 (+9.7%).
-However, soft labels excel under distribution shift (Light: 0.841 vs.
-0.794, +5.9%) and moderate adversarial scenarios.
-Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely.
-Soft labels provide gradient smoothing beneficial for generalization but can "average out" signal under heavy contamination.
+Heavy Noise Heavy 0.375 0.418 0.725 0.675
+Dist. Shift Light 0.653 0.717 0.752 0.738
+Adversarial Medium 0.636 0.643 0.730 0.744
+Heavy-Tailed Heavy 0.513 0.508 0.722 0.714
+Hard wins: 7/12 scenarios (Heavy Noise 2/3, Adversarial 2/3, Heavy-Tailed 2/3)
+Soft wins: 5/12 scenarios (Dist. Shift 2/3, Heavy Noise 1/3, others)
+The soft vs.
+hard label comparison reveals a nuanced picture (Table 3).
+Hard labels win in 7/12 scenarios, particularly under Heavy Noise (2/3), Adversarial (2/3), and Heavy-Tailed (2/3) settings.
+However, soft labels excel under distribution shift (2/3 scenarios) and achieve the single highest improvement: ADPO-Soft reaches 0.725 vs.
+ADPO-Hard’s 0.675 under Heavy Noise-Heavy (+7.4%).
+Interpretation: The choice depends on noise type—hard labels provide decisive signals when noise is extreme and random, while soft labels preserve gradient information beneficial for systematic distribution shifts and moderate noise.
 5.3 Model Scaling
@@
 We train teacher policies using Soft Actor-Critic (SAC) with large networks (512-512-256 hidden dimensions) until convergence.
-For each environment, we collect a distillation dataset of 10K states, each associated with $K=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%).
+For each environment, we collect a distillation dataset of 10K states, each associated with $P=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%).
 Actions are ranked using the teacher’s Q-function, providing preference labels without expensive rollouts.
@@
   • ADPO-Self-Anchor: Student anchored to its own random initialization $\pi_{\mathrm{init}}$, learning from ranked action preferences.
-  • ADPO-Teacher-Anchor: Student anchored to frozen teacher policy $\pi_{\mathrm{teacher}}$, enabling both preference learning and geometric regularization.
+  • ADPO-Self-Anchor-EMA: Student anchored to an exponential moving average (EMA) of its own policy, providing a slowly moving reference that tightens coupling while remaining stable.
 Table 5: Policy distillation on continuous control (MuJoCo).
-We evaluate on HalfCheetah-v5 and Hopper-v5 (Walker2d-v5 excluded due to poor teacher: $18.86\pm 5.75$).
 Results averaged over 5 seeds per environment.
-ADPO-Teacher-Anchor achieves best return in both environments while maintaining extremely low KL divergence, confirming that fixed anchoring enables effective knowledge transfer.
+ADPO methods achieve superior returns while maintaining dramatically lower KL divergence.
+ADPO-self-anchor uses fixed initial policy as anchor; ADPO-self-anchor-EMA uses exponential moving average.
+Note: Anchor strategy shows task-dependent effectiveness—fixed anchor excels on HalfCheetah, moving anchor on Hopper.
 Environment Method Return NDCG KL $\downarrow$
-HalfCheetah-v5 (Teacher: 53.5) KD $311.2\pm 10.4$ 0.984 23.73
-ADPO-Self-Anchor $131.7\pm 23.3$ 0.626 19.27
-ADPO-Teacher-Anchor $\mathbf{206.7\pm 24.7}$ 0.667 0.08
-Hopper-v5 (Teacher: 107.5) KD $25.1\pm 5.9$ 0.990 52.78
-ADPO-Self-Anchor $34.1\pm 4.5$ 0.587 81.30
-ADPO-Teacher-Anchor $\mathbf{65.4\pm 1.7}$ 0.769 0.01
+HalfCheetah-v5 (Teacher: 8476.5) KD $309.0\pm 98.2$ 0.857 30.50
+ADPO-self-anchor $\mathbf{279.3\pm 53.2}$ 0.765 10.45
+ADPO-self-anchor-EMA $166.8\pm 227.0$ 0.858 0.62
+Hopper-v5 (Teacher: 1169.8) KD $36.9\pm 20.2$ 0.799 16.98
+ADPO-self-anchor $177.3\pm 158.4$ 0.684 7.61
+ADPO-self-anchor-EMA $\mathbf{855.6\pm 269.1}$ 0.755 3.97
 Results and analysis.
-Table 6 validates our hypothesis that fixed anchors excel at offline distillation:
+Table 5 validates our hypothesis that anchoring excels at offline distillation, with task-dependent anchor strategies:
   1. 1.
-    ADPO-Teacher-Anchor dominates on both environments.
-On HalfCheetah-v5, it achieves 206.7 (387% of teacher’s 53.5), significantly outperforming KD’s 311.2-311.2 despite KD’s near-perfect ranking quality (NDCG=0.984).
-On Hopper-v5, ADPO-Teacher-Anchor reaches 65.4 (61% of teacher’s 107.5), again surpassing KD’s 25.1 (23%).
-This demonstrates that ranking quality does not equal decision-making quality—KD can perfectly mimic preference rankings while failing to learn effective policies.
+    ADPO dramatically outperforms KD on both environments.
+On HalfCheetah-v5, ADPO-self-anchor achieves 279.3 vs KD’s 309.0-309.0 (588-point improvement), while on Hopper-v5, ADPO-self-anchor-EMA reaches 855.6 (73% of teacher’s 1169.8) vs KD’s 36.9 (3%).
+This demonstrates that ranking quality does not equal decision-making quality—KD achieves high NDCG (0.857, 0.799) but fails to learn effective policies.
   2. 2.
@@
   3. 3.
-    ADPO-Self-Anchor underperforms (131.7-131.7 and $34.1$ respectively), confirming that anchor quality matters for distillation: a random initialization provides a poor geometric center.
-In contrast, the pre-trained teacher provides an informative, stable anchor that guides learning.
+    Task-dependent anchor strategy confirmed.
+Fixed anchor (ADPO-self-anchor) achieves best return on HalfCheetah (279.3), while moving anchor (ADPO-self-anchor-EMA) excels on Hopper (855.6).
+Both achieve lower KL than KD, with moving anchor providing the tightest coupling (0.62 and 3.97).
+This validates the theoretical prediction from Section 3.5.
   4. 4.
-    Task-dependent anchor selection validated. These distillation results (fixed anchor wins) complement the online exploration results of Section 5.4 (dynamic anchor wins), confirming our hypothesis:
+    Anchor strategy trade-off validated. These distillation results complement the online exploration results of Section 5.4, confirming our hypothesis:
     • Online exploration $\to$ dynamic anchor adapts to policy evolution
-    • Offline distillation $\to$ fixed anchor provides stable teacher targets
+    • Offline distillation $\to$ both fixed and moving anchors excel, with task-dependent preferences
 Connection to theory.
-These results directly validate our geometric framework (Section 3, Lemma 3.5):
-  • KD vs.
-ADPO: KD operates in “absolute probability space” without a fixed geometric center (Table 8).
-ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
-  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates.
+These results directly validate our geometric framework (Section 3.2, Lemma 3.6):
+  • KD vs. ADPO: KD operates in absolute probability space without a fixed geometric center. ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
+  • Anchor flexibility: Whether anchoring to an old policy (online RL) or to a fixed reference (e.g., the initial policy or an EMA of the student in offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates.
 This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
@@
 This is more robust to teacher imperfections and enables students to surpass teachers when beneficial.
-Practical implications.
-The distillation results demonstrate that ADPO’s anchoring mechanism is not limited to synthetic preference optimization tasks.
-It provides a principled, geometry-aware approach to knowledge transfer in real-world RL pipelines, combining the stability of trust-region methods with the simplicity of direct policy optimization.
-(a) Return comparison across environments
-(b) KL divergence vs Return trade-off
-Figure 4: Policy distillation results on MuJoCo environments.
-Teacher performance: HalfCheetah-v5 (53.5), Hopper-v5 (107.5).
-Left: Student return comparison across both environments.
-Right: Trade-off between return and KL divergence to teacher.
-ADPO-Teacher-Anchor achieves best return in both environments (206.7 on HalfCheetah = 387% of teacher; 65.4 on Hopper = 61% of teacher) while maintaining extremely low KL (0.08 and 0.01), confirming that fixed anchoring enables stable knowledge transfer with implicit trust-region regularization.
-5.6 Hyperparameter Sensitivity
-To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$.
-Figure˜5 shows the results.
-Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$.
-This stability suggests it can be deployed with minimal hyperparameter tuning.
-In contrast, the KDE-anchored listwise method shows greater sensitivity (WinMass 0.37–0.55), indicating that it benefits from careful tuning.
-(a) Pairwise anchored
-(b) Listwise raw
-(c) Listwise KDE
-Figure 5: Temperature sensitivity ablation.
-Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment.
-KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
-6 Discussion
-Why does anchoring work?
-Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$.
-The crucial difference: Standard DPO anchors the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$, while ADPO anchors each score individually before comparing.
-This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.5)—anchoring is trust region by design.
-Our experiments demonstrate relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor.
-Hard vs. soft labels.
-In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring.
-Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that "average out" signal.
-However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence.
-Caveat: These results depend on our controlled synthetic noise generation.
-In real-world settings with calibrated multi-annotator uncertainty, soft labels may be more advantageous.
-Listwise dominance.
-Listwise methods win in 9/12 scenarios (Table 3) by leveraging full group information and anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases—especially valuable when rewards have group-level shifts (different annotators with different baselines).
 Practical guidance.
@@
 General (unknown noise) ADPO Listwise-Raw Robust across all scenarios
-Knowledge Distillation ADPO-Teacher-Anchor Superior returns & stability (e.g., +206 vs -311)
+Knowledge Distillation ADPO-Self-Anchor Superior returns & stability (e.g., 855.6 vs 36.9 on Hopper; 279.3 vs -309.0 on HalfCheetah)
 6 Limitations and Future Work
@@
 We introduced Anchored Direct Preference Optimization (ADPO), a unified projection framework that reveals fundamental connections across major learning paradigms.
-By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(s-s^{\mathrm{ref}})/\tau$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature.
+By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature.
 This unification demonstrates that anchoring is not merely a technique for preference learning, but a general principle for robust policy optimization.
@@
 with $\sigma_{\text{out}}^{2}=100\sigma^{2}$. The severity levels correspond to:
-  • Light: SNR = $\frac{\operatorname{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
-  • Medium: SNR = 0.5, $p_{\text{out}}=0.10$
-  • Heavy: SNR = 0.2, $p_{\text{out}}=0.20$
+  • Light: noise scale $\sigma=1.0$, $p_{\text{out}}=0.05$
+  • Medium: noise scale $\sigma=2.0$, $p_{\text{out}}=0.10$
+  • Heavy: noise scale $\sigma=2.5$, $p_{\text{out}}=0.20$
 Distribution Shift.
@@
 E.3 Temperature Sensitivity (Full Grid)
-Figure 5 in the main text shows three representative methods. The full temperature grid for all methods is provided in the supplementary materials. Key findings:
+Temperature sensitivity analysis shows that ADPO maintains robust performance across a wide range of temperature values. Key findings:
   • Pairwise-ADPO maintains WinMass $\in[0.61,0.62]$ across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$;